### v1

# PLRD: Partially Linear Regression Discontinuity Inference

Abstract
Regression discontinuity designs have become one of the most popular research designs in empirical economics. We argue, however, that widely used approaches to building confidence intervals in regression discontinuity designs exhibit suboptimal behavior in practice: In a simulation study calibrated to high-profile applications of regression discontinuity designs, existing methods either have systematic under-coverage or have wider-than-necessary intervals. We propose a new approach, partially linear regression discontinuity inference (PLRD), and find it to address shortcomings of existing methods: Throughout our experiments, confidence intervals built using PLRD are both valid and short. We also provide large-sample guarantees for PLRD under smoothness assumptions.
1 Introduction
The fractionWe are grateful to Evan Munro for sharing results on the behavior of existing methods for RDDs under generative adversarial simulation following Athey et al. (2024), and to P. Aronow for drawing our attention to their behavior on pure noise. This work was supported by the National Science Foundation under grant SES-2242876, the Office of Naval Research under grants N00014-17-1-2131 and N00014-19-1-2468, and gift from Amazon. of empirical NBER working papers using regression discontinuity designs (RDDs) increased from nearly zero in 1995 to more than 10% by 2015 (Currie, Kleven, and Zwiers, 2020, Figure 4B). Despite the ubiquity of RDDs in applied work, however, there are still open questions as to econometric best practices for inference in RDDs. In particular, although standard interpretation of inferential statements relies on confidence intervals achieving nominal coverage of true treatment effects, we find that widely used methods systematically fail to achieve nominal coverage.
rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
Dataset EC width EC width EC width EC width
Lee (2008) 91.0% 0.036 95.4% 0.042 97.6% 0.052 98.3% 0.035
Cattaneo et al. (2015) 90.7% 5.353 95.1% 6.281 97.4% 6.980 97.3% 4.525
Matsudaira (2008) (read.) 93.8% 0.061 96.9% 0.069 98.4% 0.096 97.1% 0.060
Matsudaira (2008) (math) 88.6% 0.045 93.0% 0.054 97.8% 0.072 97.0% 0.051
Jacob and Lefgren (2004) (read.) 88.4% 0.067 95.0% 0.080 98.2% 0.104 96.5% 0.065
Jacob and Lefgren (2004) (math) 87.8% 0.078 93.4% 0.093 97.1% 0.120 95.0% 0.069
Oreopoulos (2006) 84.1% 0.196 94.6% 0.264 98.9% 0.316 96.1% 0.194
Ludwig and Miller (2007) 91.2% 4.058 95.7% 4.786 97.6% 5.238 97.4% 3.542
Meyersson (2014) 89.5% 6.362 93.3% 7.531 96.3% 8.099 97.3% 5.095
Table 1: Empirical coverage (EC) and width of 95% confidence intervals from different methods, on simulated RDDs calibrated to nine applications. Following Athey et al. (2024) we generate our simulation distributions by training a Wasserstein Generative Adversarial Neural Network (WGAN) on the original data. We report results for rdrobust (Calonico et al., 2014) (both the “b.c.” method justified by asymptotics and the “rob” method with a finite-sample variance correction) and RDHonest (Armstrong and Kolesár, 2018). We also report results for our proposed plrd method. We highlight (in green) the entries whose coverage is not statistically significantly (i.e., at the 5% level using a one-sided binomial test) below nominal 95% coverage and (in blue) the entries with the shortest widths among those delivering valid coverage. All results are aggregated across 1000 simulation replications.
We assess the performance of existing methods by evaluating them on simulations calibrated to nine high-profile applications of RDDs. We focus on approaches proposed by Calonico, Cattaneo, and Titiunik (2014) and Armstrong and Kolesár (2018), as implemented in the R functions rdrobust and RDHonest respectively. These methods are representative of current state of the art for inference in RDDs (Stommes, Aronow, and Sävje, 2023), and both offer automated end-to-end confidence interval constructions that can directly be applied to raw data. The method rdrobust by Calonico et al. (2014) is the dominant approach to regression-discontinuity inference in current econometric practice. [^1] This is the only general purpose regression discontinuity package listed in the “Econometrics Task View” for highly used R packages on CRAN (https://cran.r-project.org/web/views/Econometrics.html); and both the R (Calonico et al., 2015) and Stata (Calonico et al., 2017) versions of the function are widely used and cited. The method starts by running a local linear regression as recommended by Hahn, Todd, and der Klaauw (2001), and then applies a 2nd-order bias correction to mitigate curvature bias. The method rdrobust offers two options for building confidence intervals—a basic (b.c.) method which inherits the standard error estimator from the original local linear regression (and is justified by asymptotic arguments), and a robust (rob.) method that incorporates a finite-sample variance inflation to account for the bias correction—and we consider both. Meanwhile, the method RDHonest by Armstrong and Kolesár (2018) explicitly inflates the width of local linear regression confidence intervals to accommodate worst-case plausible curvature bias (rather than trying to correct for this bias). [^2] In principle, the method RDHonest asks the user to specify a worst-case bound on the curvature of the underlying conditional response functions up front. However, the method also has a default fall-back option where curvature bounds are derived using the rule-of-thumb of Armstrong and Kolesár (2020); in our experiments we use this default.
To assess these methods, we use generative adversarial neural networks to generate simulation instances that are representative of real-world applications, as recommended in Athey, Imbens, Metzger, and Munro (2024). [^3] The generative adversarial learning approach provides a simulator whose data-generating distribution closely mimics the moments and idiosyncracies of the original data (see Fig. 1 for an illustration). The generated simulator does not use an explicit parametric model and does not enforce any built-in smoothness assumptions, thus bringing it towards a neutral comparison ground for evaluating methods that were motivating under different assumptions. We summarize results for 95% confidence intervals built using the considered methods in Table 1; see Section 4 for details on the simulation design. A first observation is that rdrobust systematically undercovers the true regression discontinuity parameter. The basic (b.c.) intervals only achieve results consistent with nominal coverage in 1 out of 9 settings. The finite-sample adjusted intervals (rob.) are wider, but still only achieve nominal coverage in 6 out of the 9 settings. These qualitative findings do not appear to be specific to our simulation designs; for example, in Table 2 we observe similar coverage properties when applying the method to pure noise with no relationship between the outcome and the running variable.
rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
Dataset EC width EC width EC width EC width
pure noise 88.7% 0.997 92.8% 1.181 94.9% 1.293 96.3% 0.847
[…]
Meanwhile, the method RDHonest achieves valid coverage, but at the cost of substantially wider intervals: The median ratio between the length of these intervals and the basic bias-corrected (b.c.) ones varies between 1.27 to 1.60 in our simulations. Furthermore, these confidence intervals are often excessively conservative, thus suggesting that some of this additional width isn’t strictly needed. [^4] In several settings, RDHonest achieves 98% coverage or above—and this may come at a substantial cost in width. For example, in the classical setting with no bias adjustment, Gaussian 98% confidence intervals are 19% wider than the conventional 95% intervals. Our findings leave open a pragmatic question: Is it possible to design automated confidence intervals that reliably achieve nominal coverage in realistic settings, all while matching rdrobust in terms of interval length?
The main contribution of this paper is a method, partially linear regression discontinuity inference (PLRD), that provides an affirmative answer to this question—at least in the context of the numerical settings we consider. As described further below, algorithmically, the PLRD estimator adapts the method of Armstrong and Kolesár (2018) to accommodate a partially linear treatment effect model; conceptually, however, our full approach fits within the econometric framework of Calonico et al. (2014). We implement our method in the package plrd for R (R Core Team, 2023), which provides end-to-end functionality for regression discontinuity inference without relying on the user to specify hard-to-choose tuning parameters. As seen in Tables 1 and 2, plrd achieves nominal and near-exact coverage throughout, and in every instance yields the shortest intervals among those with valid coverage.
2 The PLRD Estimator
Consider the sharp RDD formalized using potential outcomes as in Imbens and Lemieux (2008). For each unit $i=1,\,\ldots,\,n$, we observe IID-sampled pairs $(X_{i},\,Y_{i})$, where $X_{i}\in\mathbb{R}$ is the running variable and $Y_{i}\in\mathbb{R}$ is the outcome of interest. Following the sharp discontinuity design, we assume there exists a cutoff $c\in\mathbb{R}$ such that treatment is assigned as $W_{i}=\text{$\mathbf{1}$}{\left\{X_{i}\geq c\right\}}$. We posit potential outcomes $\{Y_{i}(0),\,Y_{i}(1)\}$ such that $Y_{i}=Y_{i}(W_{i})$, and our goal is to estimate the conditional average treatment effect at the cutoff, $\tau_{c}:=\mu_{1}(c)-\mu_{0}(c)$ where $\mu_{w}(x):=\mathbb{E}[Y_{i}(w)|X_{i}=x]$.
It is well known that, provided the $\mu_{w}(x)$ are continuous in $x$ and that the running variable has continuous support around $c$, we can identify our target parameter as a discontinuity in the conditional response surface at $X_{i}=c$ (Hahn et al., 2001):
$\tau_{c}=\lim_{x\downarrow c}\mathbb{E}[Y_{i}\,|\,X_{i}=x]-\lim_{x\uparrow c}\mathbb{E}[Y_{i}\,|\,X_{i}=x].$ (1)
In order to obtain practical estimation or inference results for $\tau_{c}$, though, one needs to make further regularity assumptions on the $\mu_{w}(x)$; our PLRD method relies on the following two assumptions.
Assumption 1.
$\mu_{w}(x)$ are 3 times continuously differentiable in $x$.
Assumption 2.
The conditional average treatment effect function $\tau(x)=\mu_{1}(x)-\mu_{0}(x)$ has a linear dependence in $x$, i.e., $\tau(x)=\tau_{c}+\beta(x-c)$ for some $\beta\in\mathbb{R}$.
Assumption 1 is a standard smoothness assumption; this is exactly the same as the main smoothness assumption used by Calonico et al. (2014) to justify rdrobust. Assumption 2 implies that the treatment effect has a simpler dependence on the running variable than the baseline effect $\mu_{0}(x)$. This assumption is well in line with assumptions often made in the literature on heterogeneous treatment effect estimation (e.g., Wager, 2024, Chapter 4); however, it has not typically been used in the RDD literature to date, which is why we emphasize this assumption in naming our method. Given these two assumptions, PLRD proceeds as follows:
  1. 1.
    Under Assumption 1 there exists a finite constant $B$ such that the second derivatives $\mu^{\prime\prime}_{w}(x)$ are $B$-Lipschitz in a neighborhood of $c$; and it is possible to identify such a $B$ from data in the large-sample limit. Following Imbens and Kalyanaraman (2012) and Calonico et al. (2014), PLRD starts by estimating $B$ via a cubic regression.
  2. 2.
    We form the set of all possible choices of $\mu_{0}(x)$ and $\mu_{1}(x)$ consistent with the $B$-Lipschitz curvature bound estimated above and Assumption 2. Then, following Armstrong and Kolesár (2018) and Imbens and Wager (2019) we use numerical convex optimization to derive a minimax linear estimator [^5] Any estimator of the form $\widehat{\tau}=\sum_{i=1}^{n}\gamma_{i}Y_{i}$ where the weights $\gamma_{i}$ only depend on the realization of the running variables $X_{i}$ is called a “linear” estimator; the minimax linear estimator is the estimator of this form with the minimax mean-squared error conditionally on the $X_{i}$. The minimax linear estimation approach was introduced by Donoho (1994), and first applied to RDDs by Armstrong and Kolesár (2018). For a textbook overview of the minimax linear estimation approach to inference for RDDs, see Wager (2024, Chapter 8.2). for $\tau_{c}$ within this class.
  3. 3.
    We identify the worst-case bias of our estimator within this class, and construct “bias aware” confidence intervals for $\tau_{c}$ using the construction of Imbens and Manski (2004) that simulatenously accounts for the standard error and potential bias of an estimator.
Our full procedure—including some extra algorithmic details such as cross fitting—is detailed as Algorithm 1. Our main formal result about the PLRD estimator is that it yields valid confidence intervals under our assumptions in the large-sample limit.
Algorithm 1 Partially Linear Regression Discontinuity Inference (PLRD)
This algorithm provides asymptotically valid $100(1-\alpha)\%$ confidence intervals for the conditional average treatment effect $\tau_{c}$ in a sharp RDD. This procedure is implemented in the R package plrd.
Input: Data $\{(X_{i},Y_{i})\}_{i=1}^{n}$, the cutoff $c$, and the confidence level $1-\alpha$ (default is set as $95\%$).
[…]
Output: The PLRD estimator $\widehat{\tau}_{\texttt{PLRD}}$ with associated $1-\alpha$ confidence interval $\mathcal{I}_{\alpha}$.
1:Discard all observations with $|X_{i}-c|>\ell$.
2:Randomly split the data into two near-equally sized folds, with index sets $I_{1}$ and $I_{2}$.
3:Use the observations in fold $I_{2}$ to estimate the Lipschitz constant $B$ by fitting cubic polynomials before and after the threshold as in (9), and denote this estimate as $\widehat{B}_{(2)}$.
4:Enforce a lower bound $\widehat{B}_{(2)}\leftarrow\max\{\widehat{B}_{(2)},\,\varepsilon\}$.
5:Estimate $\sigma^{2}$ by running the ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $W_{i}$ for $i\in I_{2}$. Denote this estimate as $\widehat{\sigma}^{2}_{(2)}$ and record the residuals from this regression as:
$r_{i}:=Y_{i}-\widehat{\mu}_{W_{i}}(X_{i})\text{ for each }i\in I_{2}.$ (2)
[…]
$(\widehat{\gamma}_{(1)},\widehat{t}_{(1)}):=\mathop{\rm{minimize}}_{\gamma,t}\left(\widehat{B}^{2}_{(2)}t^{2}+\widehat{\sigma}^{2}_{(2)}\sum_{i\in I_{1}}\gamma_{i}^{2}\right)\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\ \sum_{i\in I_{1}}\gamma_{i}\rho(X_{i}-c)\leq t.$ (3)
7:Repeat the steps 3 to 5 with the roles of $I_{1}$ and $I_{2}$ swapped.
8:Set $\widehat{\gamma}_{i}=\gamma_{(k),i}/2$ if $i\in I_{k}$ ($k=1,2$), and $\widehat{\gamma}_{i}=0$ if $|X_{i}-c|>\ell$. Compute $\widehat{\tau}_{\texttt{PLRD}}=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i}$, and $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}r_{i}^{2}$ where the $r_{i}$ are the residuals from the least squares regressions in step 4.
9:Report the following $100(1-\alpha)\%$ confidence interval for $\tau_{c}$:
$\mathcal{I}_{\alpha}:=[\widehat{\tau}_{\texttt{PLRD}}-\widehat{h}_{\alpha},\,\widehat{\tau}_{\texttt{PLRD}}+\widehat{h}_{\alpha}],\quad\widehat{h}_{\alpha}:=\min\,\{h:\mathbb{P}(|b+\widehat{s}(\widehat{\gamma})Z|\leq h)\geq 1-\alpha\text{ for all }|b|\leq\widehat{b}\,\},$ (4)
where $Z\sim\mathcal{N}(0,1)$, and $\widehat{b}:=(\widehat{B}_{(2)}\widehat{t}_{(1)}+\widehat{B}_{(1)}\widehat{t}_{(2)})/2$.
Theorem 1.
Consider a sharp regression discontinuity design under Assumptions 1 and 2, where the running variable $X_{i}$ has a density strictly positive at the threshold $c$, $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$ is bounded away from 00, i.e., $\sigma_{i}\geq\sigma_{\min}$ for some $\sigma_{\min}>0$, and $\mathbb{E}[|Y_{i}-\mu_{W_{i}}(X_{i})|^{q}\mid X_{i}=x]\leq C$ uniformly over all $x$, for some $q>2$ and $C\in(0,\infty)$. Suppose furthermore that we execute the initial cubic regression used to pick $B$ over a pilot bandwidth of length $\ell$ with
$\lim_{n\rightarrow\infty}n^{1/12}\ell=0,\ \ \ \ \lim_{n\rightarrow\infty}n^{1/7}\ell=+\infty,$ (5)
and choose any $\varepsilon>0$ in Algorithm 1. Then, level-$(1-\alpha)$ PLRD intervals $\mathcal{I}_{\alpha}$ are asymptotically valid: For any $10<\alpha<$ 1,
$\liminf_{n\to\infty}\,\mathbb{P}[\tau_{c}\in\mathcal{I}_{\alpha}]\geq 1-\alpha.$ (6)
We note that our algorithm does in principle require two tuning parameters, namely the pilot bandwidth $\ell$ used to run the initial cubic regression to choose $B$ and the lower bound $\varepsilon$ for $\widehat{B}$. We note, however, that our algorithmic performance is largely insensitive to these choices (and simple defaults appear to provide good performance in practice); and similar objects are also required in Calonico et al. (2014) and Imbens and Kalyanaraman (2012). First, regarding $\ell$, this bandwidth can in practice be chosen simply to be “large”: We only need $\ell$ to decay faster than $n^{-1/12}$, whereas the mean-squred error optimal bandwidth for local linear regression in our setting (i.e., under Assumption 1) scales as $n^{-1/7}$ (Cheng, Fan, and Marron, 1997). In our implementation, we by default simply set $\ell$ to contain the full range of the data, and this is choice is what we use in all our experiments. [^6] This choice mirrors Imbens and Kalyanaraman (2012, p. 942, step 2) who seed their bandwidth-selection procedures by fitting global cubic polynomials to the full dataset; see also Calonico et al. (2014, S.2.6, step 0). As such, practitioners should be able to get reliable results out of our method unless the nature of the relationship between $X_{i}$ and $Y_{i}$ changes completely as we get far from the cutoff—and in settings like this it would be possible to pick $\ell$ by plotting the data. Meanwhile, $\varepsilon$ will have no effect in large samples as long as we choose it to be small enough that $\varepsilon<|\mu^{\prime\prime\prime}(c)|$. [^7] One interesting question is what happens at the critical point where $\mu^{\prime\prime\prime}(c)=0$; as, in this case, our constraint does necessarily bind and we fall back to using $\varepsilon$ as our curvature bound. The challenge is that, when $\mu^{\prime\prime\prime}(c)=0$, our cubic polynomial estimator $\widehat{B}$ (in step 3 of Algorithm 1) converges to 0, but the exact way it does so is left underspecified under Assumption 1. Our use of the lower bound $\varepsilon$ guarantees validity of our confidence intervals at the expense of some excess conservativeness when $\mu^{\prime\prime\prime}(c)=0$. We note that existing methods, including Calonico et al. (2014) and Imbens and Kalyanaraman (2012), also do not offer guidance on how to potentially benefit from “superefficiency” in similar critical cases; i.e., our approach is again in line with current best practices.
3 Relationship to Existing Methods
Regression discontinuity designs were originally introduced by Thistlethwaite and Campbell (1960). The modern econometric literature goes back to Hahn et al. (2001), who emphasized the identification result (1) and recommended estimation $\tau_{c}$ via local linear regression,
$(\widehat{a}_{w},\widehat{\beta}_{w})=\mathop{\rm argmin}_{(a,\beta)\in\mathbb{R}^{2}}\ \left\{\sum_{i\,:\,W_{i}=w}K\left(\frac{|X_{i}-c|}{h_{n}}\right)\left(Y_{i}-a-\beta(X_{i}-c)\right)^{2}\right\},\ \ \ \ \widehat{\tau}_{\texttt{LLR}}=\widehat{a}_{1}-\widehat{a}_{0},$ (7)
where $h_{n}\rightarrow 0$ is a bandwidth that “localizes” the regression and $K(\cdot)$ is a kernel function with support on [1, 1]11[-1,\,1][ - 1 , 1 ]. In the decade following Hahn et al. (2001), local linear (or polynomial) regression remained the prevalent approach approach to inference in RDDs, with confidence intervals produced by applying off-the-shelf tools for heteroskedasticity-robust inference directly to the above regression (Imbens and Lemieux, 2008; Lee and Lemieux, 2010). A limitation of this approach, however, is that it leads to contradictory guidance on the bandwidth choice under a non-parametric specification as given in, e.g., Assumption 1. The accuracy of $\widehat{\tau}_{\texttt{LLR}}$ depends delicately on the bandwidth $h_{n}$ (Imbens and Kalyanaraman, 2012) and, unfortunately, for choices of $h_{n}$ that optimize this accuracy the bias and standard error of $\widehat{\tau}_{\texttt{LLR}}$ are of the same order—so off the shelf local linear regression inference will not have valid coverage for this choice of $h_{n}$. The practitioner is thus forced to invoke “undersmoothing”, whereby they voluntarily use a non-accuracy-optimizing choice of $h_{n}$ to justify validity of their inferential procedure.
In a major advance, Calonico et al. (2014) and Armstrong and Kolesár (2018) introduced paradigms for inference in RDDs that formally account for the bias of local linear regression. As discussed in Stommes et al. (2023), these methods offer material improvements in the credibility of regression-discontinuity analyses over the earlier practice. The main idea in Calonico et al. (2014) is to estimate and correct for the bias of local linear regression by leveraging higher-order smoothness; we refer to this approach as the “bias-corrected” approach. Calonico, Cattaneo, Farrell, and Titiunik (2019) discuss the use of covariates under the bias-corrected approach to RDDs while Cattaneo and Titiunik (2022) provide a recent review. In contrast, the “bias-aware” approach developed by Armstrong and Kolesár (2018) involves conservatively widening intervals to account for worst-case bias (without needing to actually estimate the bias); see Kolesár and Rothe (2018), Imbens and Wager (2019) and Noack and Rothe (2024) for further applications of this paradigm.
Interestingly, not only do the above papers under the bias-corrected vs. bias-aware paradigms explore different methodological strategies for inference in RDDs; they also focus on proving formal results under different formal paradigms. Calonico et al. (2014) and follow-ups focus entirely on asymptotic validity results—as we do here. They fix a data-generating process, and guarantee that they eventually get coverage to this data-generating process as the sample size gets large (as is also done in (6)). In contrast, most papers written under the bias-aware paradigm specify a quantitative smoothness class for the $\mu_{w}(x)$ (e.g., all functions with a second derivative bounded by a pre-specified constant $B$), and then seek guarantees that hold in finite samples and uniformly over this whole class. Getting uniform guarantees in finite samples of course requires stronger assumptions than getting pointwise asymptotic guarantees; and for this reason papers associated with the bias-aware paradigm often start by making stronger assumptions than those associated with the bias-corrected paradigm. We emphasize, however, that this is due to the sought formal guarantees—and it’s not that bias-aware methods inherently require stronger assumptions to work than bias-corrected methods.
Our proposed method, PLRD, straddles these two literatures. Our confidence intervals are rooted in the bias-aware paradigm of Armstrong and Kolesár (2018), and we use numerical convex optimization to construct our estimator as in Imbens and Wager (2019). In contrast, our formal results are focused on pointwise asymptotics, and the assumptions and guarantee type given in our Theorem 1 closely mirror those in Theorem 1 of Calonico et al. (2014). One notable paper that also straddles these literatures in Armstrong and Kolesár (2020), who propose bias-aware confidence intervals for local linear regression with pointwise asymptotic guarantees. In particular, they show that if one runs local linear regression as in (7) with the mean-square-error optimal bandwidth choice $h_{n}$ following Imbens and Kalyanaraman (2012), then under generic conditions we can build valid, bias-aware 95% confidence intervals for $\tau_{c}$ as $\widehat{\tau}_{\texttt{LLR}}\pm 2.18\times\text{standard errors}$ rather than the conventional $\widehat{\tau}_{\texttt{LLR}}\pm 1.96\times\text{standard errors}$. We provide a comparison of PLRD to this simple rule of thumb in Appendix A.
Wrapping up, PLRD can best be understood within the context of earlier developments in the bias-aware paradigm; and in fact steps (2) and (3) of our algorithmic outline fall squarely within the roadmap for minimax linear inference via numerical convex optimization as presented in Imbens and Wager (2019). However, in a deviation from existing work:
  • We rely on higher-order smoothness and a preliminary cubic fit to initialize tuning parameters of our algorithm; in Armstrong and Kolesár (2018), Kolesár and Rothe (2018) and Imbens and Wager (2019), a parameter analogous to the $B$ derived in our step (1) would need to be pre-specified using subject matter knowledge or chosen using a heuristic.
  • Because of this data-driven choice of $B$ in step (1), we are not able to provide uniform, finite-sample inferential guarantees; [^8] And, as shown in Armstrong and Kolesár (2018), this would in fact be impossible. thus, like in Calonico et al. (2014), we focus on pointwise asymptotic inferential statements.
  • The specific worst-case class we use in our step (2), i.e., with $B$-Lipschitz curvature bounds and a linear treatment effect assumption, has not been previously considered in the literature. The most standard choice in applications of Armstrong and Kolesár (2018), Kolesár and Rothe (2018) and/or Imbens and Wager (2019) has been to only assume that the $\mu_{w}(x)$ have second derivatives bounded by $M$.
Regarding this last point, the earlier focus on the $M$-bounded curvature class appears to have been largely guided by intuitive simplicity of this class rather than any conceptual or empirical arguments. Given the empirical performance of PLRD on calibrated simulations, we suggest that our $B$-Lipschitz curvature class with linear treatment effects may be a good default choice in many applications. We conduct an ablation analysis to examine the effect of different choices on the behavior of our estimator in Appendix A.
4 Empirical Study
We now describe in more detail the experiment presented in the introduction (Table 1). As context for this experiment, Table 3 shows a replication of a number of high-profile RDD applications using a variety of methods; citations for each application are given in the leftmost column and further description of each benchmark dataset is provided Appendix B. We immediately see that the confidence intervals given by rdrobust, especially with the ${\it(b.c.)}$ interval choice, are much shorter than those given by RDHonest, thus naturally making the method attractive to practitioners. But how much can we trust these intervals?
Dataset rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
Lee (2008) $0.059\pm 0.022$ $0.059\pm 0.059$ $0.058\pm 0.032$ $0.079\pm 0.020$
Cattaneo et al. (2015) $7.507\pm 2.859$ $7.507\pm 3.413$ $8.030\pm 4.210$ $6.440\pm 2.374$
Matsudaira (2008) (read.) $0.030\pm 0.028$ $0.030\pm 0.032$ $0.031\pm 0.046$ $0.029\pm 0.033$
Matsudaira (2008) (math) $0.072\pm 0.025$ $0.072\pm 0.030$ $0.057\pm 0.041$ $0.077\pm 0.026$
Jacob and Lefgren (2004) (read.) $0.143\pm 0.042$ $0.143\pm 0.048$ $0.127\pm 0.046$ $0.093\pm 0.037$
Jacob and Lefgren (2004) (math) $0.203\pm 0.035$ $0.203\pm 0.039$ $0.187\pm 0.058$ $0.152\pm 0.032$
Oreopoulos (2006) $0.099\pm 0.104$ $0.099\pm 0.143$ $0.070\pm 0.125$ $0.072\pm 0.115$
Ludwig and Miller (2007) $-2.781\pm 2.363$ $-2.781\pm 2.682$ $-3.153\pm 2.828$ $-1.824\pm 2.019$
Meyersson (2014) $2.983\pm 2.797$ $2.983\pm 3.293$ $3.024\pm 3.109$ $3.162\pm 2.294$
Table 3: 95% confidence intervals produced by different methods on cited benchmark datasets.
Figure 1: The original data from one of our benchmark datasets, Lee (2008), as well as a simulated dataset produced using a WGAN trained to the original data.
A fundamental challenge in assessing methods for causal inference on real-world data is that we typically do not have access to a ground truth we can use for evaluation. In light of this challenge, Athey et al. (2024) proposed using Wasserstein Generative Adversarial Networks (WGANs) to generate synthetic data that mimic benchmark datasets—and thus obtaining simulation instances that enable us to compare methods in settings that are as realistic as possible. Our WGAN evaluation pipeline involves the following three steps:
  1. 1.
    Training: We train WGANs on each benchmark dataset to mimic the underlying data generating process. This involves generating running variables and outcomes for both treated and control units.
  2. 2.
    Generation: Once trained, the WGANs generate multiple synthetic datasets that reflect the characteristics of the original data. These synthetic datasets serve as a proxy for the real data in our simulation study.
  3. 3.
    Evaluation: We apply the various methods, including our proposed PLRD method, to the synthetic datasets. The performance of these methods is then evaluated based on the empirical coverage and the average width of the confidence intervals.
As illustrated in Figure 1, a trained WGAN is able to simulate data whose distribution is very similar to that of the original data; thus suggesting that if a method gets reliable coverage on the simulator we can also trust it on the real data.
The WGANs take as input random Gaussian noise $Z\in\mathbb{R}^{3}$. We train one generator $\widehat{g}_{x}(x)$ for the unconditional distribution of the running variable $X$, and two generators $\widehat{g}_{w}(x,z)$ ($w=0,1$) for the conditional distribution of the outcomes for treated and control units, given the running variable. We can calculate the ground truth $\widetilde{\tau}$ for the GAN estimated distribution by sampling random noise $Z_{q}\in\mathbb{R}^{3}$ for $q=1,\dots,Q$, where $Q$ is a large number ($10^{6}$ in our simulations), and generating outcomes at the threshold, as follows:
$\widetilde{\tau}:=Q^{-1}\sum_{q=1}^{Q}(\widehat{g}_{1}(0,Z_{q})-\widehat{g}_{0}(0,Z_{q})).$
Finally, we generate $i=1,\dots,N=10^{3}$ simulated datasets with confidence intervals $[a_{i,k},b_{i,k}]$, where the index $k$ stands for the $k$-th method we use to produce these CIs. The empirical coverage for the $k$-th method is given by $\widehat{p}_{k}=\frac{1}{N}\sum_{i=1}^{n}\mathbf{1}\{a_{i,k}\leq\widetilde{\tau}\leq b_{i,k}\}$. Table 1 shows the obtained results.
4.1 Revisiting existing simulation designs
For completeness, we also benchmark the PLRD estimator against other existing methods on the simulation settings originally used in Calonico et al. (2014, Section 6). [^9] Of these, Setting 1 was first used by Imbens and Kalyanaraman (2012), while Settings 2 and 3 were original to Calonico et al. (2014). In each of these settings, we draw the running variable $X_{i}\sim 2\,\text{Beta}(2,4)-1$ for $n=500$ samples, and the outcome is given by $Y_{i}=m(X_{i})+\varepsilon_{i}$, where $\varepsilon_{i}$ are mean-zero Gaussian noise with variance $0.1295^{2}$. The form of $m(\cdot)$ varies for the different simulation settings; see Calonico et al. (2014) for details on Settings 1 through 3. Setting 4 uses the function $m=m_{\text{quad}}$ as in Imbens and Kalyanaraman (2012, Section 4.6). Figure 2 plots one realization for each of these four different simulation settings.
Table 4 presents the empirical coverage and average width of 95% CIs produced by different methods across these simulation settings. As before, rdrobust falls short of nominal coverage; note that these numbers are in line with those originally reported by Calonico et al. (2014, Table 1). RDHonest delivers valid coverage at the expense of longer coverage. And, in Settings 1, 3 and 4, plrd delivers valid coverage with intervals whose length is competitive with rdrobust.
Our method, however, does not perform well in Setting 2. The reason for this poor performance is immediately revealed by Fig. 2: Our method relies on the linear treatment effect assumption (Assumption 2), whereas Setting 2 has strong curvature in the treatment effect that is manifested in a sharp jump in curvature at the cutoff. If one believes that such a sharp jump in the curvature change is realistic, one should not rely on Assumption 2; and in Appendix A we detail a variant of plrd without this assumption (and verify that it performs well in Setting 2 of Calonico et al. (2014)). Overall, however, we still rely on Assumption 2 in the default specification of plrd as we believe such sharp curvature effects in $\tau(x)$ to be unusual in most applications. [^10] The Setting 2 of Calonico et al. (2014) was originally derived by fitting a 5th-order polynomial to data from Ludwig and Miller (2007). However, unlike with our WGAN approach, the final simulation distribution as displayed in Figure 2(b) does not resemble the original data as summarized in Ludwig and Miller (2007, Figure IV(A)) in terms of, e.g., its signal-to-noise ratio. In the original dataset, the datapoints are counties; the running variable is poverty rate; the cutoff is associated with head-start eligibility; and the outcome is child mortality rate (per 100,000 children) from causes potentially associated with Head Start.
(a) Setting 1
(b) Setting 2
(c) Setting 3
(d) Setting 4
Figure 2: Visualization of four simulation settings from Calonico et al. (2014, Section 6) and Imbens and Kalyanaraman (2012, Section 4.6). In each setting, the conditional regression surface $m(\cdot)$ is displayed as a solid line, and the points represent data for one simulation realization.
rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
coverage width coverage width coverage width coverage width
Setting 1 86.3% 0.207 91.5% 0.244 95.2% 0.290 94.6% 0.235
Setting 2 91.1% 0.313 93.2% 0.342 96.9% 0.436 51.5% 0.309
Setting 3 88.6% 0.218 93.1% 0.250 95.7% 0.301 95.9% 0.239
Setting 4 87.8% 0.207 93.0% 0.246 96.8% 0.275 97.5% 0.236
Table 4: Empirical coverage and average width of 95% confidence intervals produced by different methods for the simulation settings in Calonico et al. (2014) and Imbens and Kalyanaraman (2012). Methods are run as in Table 1.
These simulation results highlight the strengths and limitations of our plrd method: It performs well in settings that align with its underlying assumptions and can be adapted to handle more complex scenarios. However, users should be cautious when applying plrd to settings with potential highly nonlinear treatment effects $\tau(x)$ (non-linearity in the baseline response function $\mu_{0}(x)$ is not a problem, as long as it is captured by the standard conditions required in Assumption 1).
5 Formal Results
Our point estimator, i.e., the $\widehat{\gamma}_{i}$-weighted average of outcomes $Y_{i}$ with weights as given in (3) is, ignoring cross-fitting, equivalent to a minimax linear estimator for $\tau_{c}$ over the class $\mathcal{M}_{\widehat{B}}$ of all functions satisfying Assumption 2 and for which $\mu^{\prime\prime}_{0}(x)$ is globally $\widehat{B}$-Lipschitz, i.e., [^11] The conditional variance of a linear estimator as in (8) is $\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}$ with $\sigma_{i}^{2}=\operatorname{Var}[Y_{i}\,|\,X_{i}]$. If we had $\sigma_{i}^{2}=\widehat{\sigma}^{2}$ for all $i=1,\,\ldots,\,n$, then (8) would exactly be the minimax linear estimation problem for $\tau_{c}$. In the case of heteroskedasticity, the term $\widehat{\sigma}^{2}\lVert\gamma\rVert_{2}^{2}$ should be interpreted as a variance proxy that yields a practical estimator—and one that still maintains provable validity guarantees under heteroskedasticity; see Imbens and Wager (2019, p. 272) for further discussion.
$\widehat{\tau}_{\texttt{PLRD}}:=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i},\quad\widehat{\gamma}:=\mathop{\rm argmin}_{\gamma}\underbrace{\widehat{\sigma}^{2}\lVert\gamma\rVert_{2}^{2}}_{\text{variance proxy}}+\underbrace{\sup_{\mu_{0},\,\mu_{1}\in\,\mathcal{M}_{\widehat{B}}}\left(\sum_{i=1}^{n}\widehat{\gamma}_{i}\mu_{W_{i}}(X_{i})-(\mu_{1}(c)-\mu_{0}(c))\right)^{2}}_{\text{worst-case squared bias over}\mathcal{M}_{\widehat{B}}}.$ (8)
If $\widehat{B}$ were a constant chosen a-priori and $\mu^{\prime\prime}_{0}(x)$ were in fact is globally $\widehat{B}$-Lipschitz, we could use existing results for minimax linear regression-discontinuity inference (e.g., Armstrong and Kolesár, 2018; Imbens and Wager, 2019) to directly provide guarantees for $\widehat{\tau}_{\mathtt{PLRD}}$. But here, of course, this is not the case: Our choice of $\widehat{B}$ is learned in step 1 of our algorithm; and Assumption 1 only implies local—not global—Lipschitz continuity in $\mu^{\prime\prime}_{0}(x)$
The goal of this section is to provide technical tools to address these challenges, culminating in the proof of Theorem 1. We start by verifying validity of the cubic regression in the first step of our algorithm. The following result is a direct consequence of standard results for local polynomial regression (e.g., Fan and Gijbels, 1996, Theorem 3.1), and implies that $\widehat{B}_{(1)}$ and $\widehat{B}_{(2)}$ in Algorithm 1 satisfy $\widehat{B}_{(k)}\rightarrow_{p}B$ as $n\to\infty$.
Proposition 2.
Suppose that Assumptions 1 and 2 hold and that $X_{i}$ has a density bounded away from 0 in a neighborhood of $c$, and $B:=\max\{|\mu^{\prime\prime\prime}(c)|,\varepsilon\}$. Estimate $\widehat{B}$ via local cubic regression: For $w\in\{0,1\}$ run
$\widehat{\beta}_{w}:=\mathop{\rm argmin}_{\beta_{w}\in\mathbb{R}^{4}}\sum_{\begin{subarray}{c}W_{i}=w\\ |X_{i}-c|\leq\ell\end{subarray}}\left(Y_{i}-\beta_{w,0}-\beta_{w,1}(X_{i}-c)-\frac{\beta_{w,2}}{2}(X_{i}-c)^{2}-\frac{\beta_{w,3}}{6}(X_{i}-c)^{3}\right)^{2},$ (9)
with bandwidth $\ell$ scaling as in (5), and set $\widehat{B}=\max\{|\beta_{0,3}|,|\beta_{1,3}|,\varepsilon\}$. Then, $\widehat{B}\rightarrow_{p}B$ as $n\to\infty$.
Our core analytic result is the following central limit theorem for our PLRD estimator. Asymptotic validity of the PLRD confidence intervals as claimed in Theorem 1 then follows immediately.
Theorem 3.
Under the conditions of Theorem 1, $\widehat{\tau}_{\mathtt{PLRD}}$ as constructed in Algorithm 1 satisfies
[…]
where $b(\widehat{\gamma})$ is interpreted as the conditional bias of $\widehat{\tau}_{\mathtt{PLRD}}$, and $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$. Furthermore,
$\limsup_{n\to\infty}\left(\left|b(\widehat{\gamma})\right|-\widehat{b}\right)\,\big{/}\,{s}(\widehat{\gamma})\leq_{p}0,$ (11)
i.e., the learned bias bound $\widehat{b}$ in (4) is conservative for the true bias; and ${s}(\widehat{\gamma})=\mathcal{O}_{p}(n^{-3/7})$.
5.1 Proof of Theorem 1
Define $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\widehat{\sigma}_{i}^{2}$ as in Algorithm 1. It follows from White (1980) that $\widehat{s}(\widehat{\gamma})/s(\widehat{\gamma})\rightarrow_{p}1$; see Lemma 14 for details. Next, fix any $\delta\in(0,1)$. In view of (11) and Slustky’s lemma, $\mathbb{P}(A_{n})\to 1$ where $A_{n}:=\{|b(\widehat{\gamma})|\leq\widehat{b}+\delta\widehat{s}(\widehat{\gamma})\}$. We show in Lemma 13 that for any $s$, $h$, $t$ and $\delta>0$,
$\inf_{|b|\,\leq\,t+\delta s}\mathbb{P}(|b+sZ|\leq h)\geq\inf_{|b|\,\leq\,t}\mathbb{P}(|b+sZ|\leq h)-\delta,$
where $Z\sim\mathcal{N}(0,1)$. Using this with $s=\widehat{s}(\widehat{\gamma})$, $h=\widehat{h}_{\alpha}$ and $t=\widehat{b}$, we get, for every $n$,
$\mathbb{P}\left(|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\}\cap A_{n}\right)\geq\inf_{|b|\,\leq\,\widehat{b}}\mathbb{P}(|b+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha})-\delta\geq 1-\alpha-\delta,$ (12)
where the last inequality follows from the definition of $\widehat{h}_{\alpha}$. On the other hand, Theorem 3 and Slutsky’s lemma imply that $Z_{n}:=(\widehat{\tau}_{\texttt{PLRD}}-\tau_{c}-b(\widehat{\gamma}))/\widehat{s}(\widehat{\gamma})\stackrel{{\scriptstyle d}}{{\longrightarrow}}Z\sim\mathcal{N}(0,1)$. Consequently,
$\liminf_{n\to\infty}\mathbb{P}(\tau_{c}\in\mathcal{I}_{\alpha})$ $=\liminf_{n\to\infty}\mathbb{P}(|\widehat{\tau}_{\texttt{PLRD}}-\tau_{c}|\leq\widehat{h}_{\alpha})$
$=\liminf_{n\to\infty}\mathbb{P}(\{|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z_{n}|\leq\widehat{h}_{\alpha}\}\cap A_{n})$ (since $\mathbb{P}(A_{n})\to 1$)
$\geq 1-\alpha-\delta.$ (using (12) and Glivenko–Cantelli lemma)
Since $\delta\in(0,1)$ is arbitrary, we are through.
5.2 Proof of Theorem 3
The optimization problem in Algorithm 1 is inherently related to its convex dual (see Boyd and Vandenberghe (2004) for a definition). We show in Appendix C that the dual problem is given by
$\begin{split}&\widehat{\theta}_{n}(\sigma^{2},B)=\mathop{\rm argmin}_{\theta=({\rho},\lambda)\in\Theta}\widehat{L}_{n}(\theta;\sigma^{2},B),\\ &\widehat{L}_{n}(\theta;\sigma^{2},B):=\frac{1}{4n\sigma^{2}}\sum_{i=1}^{n}G^{2}(\theta;X_{i}-c,W_{i})+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n},\end{split}$ (13)
where $\theta=({\rho},\lambda)$ varies in the set
$\Theta:=\left\{(f,\lambda)\mid f:[-\ell,\ell]\to\mathbb{R},\,f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,\,f^{\prime\prime}\text{ is $\lambda_{1}$-Lipschitz},\,\lambda\in[0,\infty)\times\mathbb{R}^{5}\right\},$ (14)
and
$G({\rho},\lambda;x,w):={\rho}(x)+\lambda_{2}w+\lambda_{3}(1-w)+\lambda_{4}wx+\lambda_{5}(1-w)x+\lambda_{6}x^{2}.$
In Algorithm 1, we use sample splitting to estimate $B$ (cf. Proposition 2) and the homoscedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$. However, for ease of exposition, we assume throughout this section that we have $\widehat{B}\rightarrow_{p}B$ and $\widehat{\sigma}^{2}\rightarrow_{p}\sigma^{2}$ independent of the data, and
$\widehat{\theta}_{n}:=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})=\mathop{\rm argmin}_{\theta\,\in\,\Theta}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B}).$
We can recover the primal solution $(\widehat{\gamma}_{n},\widehat{t}_{n}\,)$ from the dual solution $\widehat{\theta}_{n}$ as:
$\widehat{\gamma}_{n,i}=-\frac{1}{2\widehat{\sigma}_{i}^{2}}\,G(\widehat{\theta}_{n};\,X_{i}-c,W_{i}),\quad\widehat{t}_{n}=\frac{1}{2\widehat{B}^{2}}\widehat{\lambda}_{n,1}.$
Next, define the population-level dual problem as:
$\begin{split}&\theta_{n}^{*}(\sigma^{2},B)=\mathop{\rm argmin}_{\theta\,\in\,\Theta}L_{n}(\theta;\sigma^{2},B),\\ &L_{n}(\theta;\sigma^{2},B):=\frac{1}{4\sigma^{2}}\mathbb{E}\left[G^{2}(\theta;X-c,W)\right]+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n}.\end{split}$ (15)
Our proof strategy hinges on establishing fast-enough convergence of $\widehat{\theta}_{n}$ to $\theta_{n}^{*}(\sigma^{2},B)$ to enable us to read-off large-sample properties of $\widehat{\tau}_{\mathtt{PLRD}}$ from the solution to the population dual (15).
Although the empirical minimization problem (13) looks similar to a moment-estimation problem, it has some technical complications, e.g., (i) the dual parameter $\theta=(\rho,\lambda)$ involves a function $\rho$ and a vector $\lambda$, (ii) the set $\Theta$ in (14) is not the Cartesian product of a function space and a vector space, and (iii) the population dual problem in (15) involves the sample size $n$. To tackle these issues, we borrow tools from empirical process theory (see, e.g., van der Vaart and Wellner (1996) for a textbook treatment). The following result, proven in Appendix D, characterizes the rate of convergence of the empirical dual optimizer $\widehat{\theta}_{n}$ to the population dual optimizer $\theta_{n}^{*}$; and implies (Corollary 5) that the weights underlying $\widehat{\tau}_{\mathtt{PLRD}}$ satisfy a Lindeberg-type condition.
Lemma 4.
Let $(\widehat{\rho}_{n},\widehat{\lambda}_{n})=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})$ be the optimizer of the empirical dual problem (13), and $({\rho}_{n}^{*},\lambda_{n}^{*})=\theta_{n}^{*}(\widehat{\sigma}^{2},B)$ be the optimizer of the population dual problem (15). Assume that the pilot bandwidth $\ell$ in Algorithm 1 satisfies (5), and $\widehat{B}\rightarrow_{p}B$. Then, under the conditions of Theorem 1,
[…]
Corollary 5.
Under the conditions of Theorem 1, the weights $\widehat{\gamma}_{i}$ underlying $\widehat{\tau}_{\mathtt{PLRD}}$ satisfy
$\frac{\|\widehat{\gamma}\|_{\infty}}{\|\widehat{\gamma}\|_{2}}=\frac{\max_{1\leq i\leq n}|\widehat{\gamma}_{i}|}{(\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2})^{1/2}}\rightarrow_{p}0.$ (16)
Corollary 5 immediately implies that the central limit theorem claimed in Theorem 3 holds; see the proof of Imbens and Wager (2019, Theorem 1) for details. We delegate showing that $s^{2}(\widehat{\gamma})=\mathcal{O}_{p}(n^{-6/7})$ (Lemma 11) and $\widehat{b}/{s}(\widehat{\gamma})=\mathcal{O}_{p}(1)$ (Lemma 12) to Appendix D. Finally, to verify (11), we note that Assumptions 2 and 1 yield $b(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}\mu_{W_{i}}(X_{i})=B\sum_{i=1}^{n}\widehat{\gamma}_{i}\rho(X_{i}-c)$, where
$\rho(x-c):=(\mu_{0}(x)-\mu_{0}(c)-\mu^{\prime}_{0}(c)(x-c)-\mu^{\prime\prime}_{0}(c)(x-c)^{2}/2)/B.$
It follows from Algorithm 1 (with cross-fitted $\widehat{B}$) that $|b(\widehat{\gamma})|\leq B\widehat{t}$ where $\widehat{t}=\widehat{b}/\widehat{B}$. Consequently,
$\limsup_{n\to\infty}\frac{\left|b(\widehat{\gamma})\right|-\widehat{b}}{{s}(\widehat{\gamma})}-\frac{(B/\widehat{B}-1)\widehat{b}}{{s}(\widehat{\gamma})}\leq_{p}0.$ (17)
Furthermore, since $\widehat{B}\rightarrow_{p}B$ and $\widehat{b}/{s}(\widehat{\gamma})=\mathcal{O}_{p}(1)$, we have
$\lim_{n\to\infty}\frac{(B/\widehat{B}-1)\widehat{b}}{{s}(\widehat{\gamma})}=_{p}0,$
and so (17) implies (11).
Appendix A Details on Baseline Methods
In this section, we give further details on the baseline methods rdrobust and RDHonest considered in our experiments, and provide further discussion of how our modeling assumptions differ from those used to-date in software for bias-aware inference in RDDs.
A.1 Bias-corrected confidence intervals
Define the local polynomial estimator of order $p$ as $\widehat{\tau}_{p}(h_{n}):=\widehat{\mu}_{1,\,p}(h_{n})-\widehat{\mu}_{0,\,p}(h_{n}),$ where $\widehat{\mu}_{1,\,p}(h_{n})$ and $\widehat{\mu}_{0,\,p}(h_{n})$ are the intercepts from local polynomial regressions of order $p$ fitted separately on each side of the threshold. More precisely, for $w\in\{0,1\}$, compute
$\widehat{\mu}_{w,\,p}(h_{n})$ $:=\mathbf{e}_{0}^{\prime}\widehat{\boldsymbol{\beta}}_{w,\,p}(h_{n}),$
$\widehat{\boldsymbol{\beta}}_{w,\,p}(h_{n})$ $:=\mathop{\rm argmin}_{\boldsymbol{\beta}\in\mathbb{R}^{p+1}}\sum_{i\,:\,W_{i}=w}K(|X_{i}-c|/h_{n})\left(Y_{i}-\mathbf{r}_{p}(X_{i}-c)^{\top}\boldsymbol{\beta}\right)^{2},$
where $\mathbf{e}_{0}=(1,0,\ldots,0)\in\mathbb{R}^{p+1}$ is the first unit vector, $\mathbf{r}_{p}(x)=\left(1,x,\ldots,x^{p}\right)^{\top}$, $K(\cdot)$ is a kernel, and $h_{n}$ is a bandwidth. The first-order bias of $\widehat{\tau}_{p}(h_{n})$ is (Calonico et al., 2014; Imbens and Kalyanaraman, 2012)
$B_{p}(h_{n})=\frac{h^{p+1}}{(p+1)!}\left[\mu_{1}^{(p+1)}(c+)-\mu_{0}^{(p+1)}(c-)\right],$
where $\mu_{w}^{(p+1)}$ denotes the derivative of $\mu_{w}(\cdot)$ of order $p+1$. Calonico et al. (2014) estimate the above bias using a higher-order local polynomial of order $q>p$ with a separate pilot bandwidth $b_{n}$, thus compute the bias-corrected estimator as
$\widehat{\tau}^{\mathrm{bc}}_{p,q}(h_{n},b_{n}):=\widehat{\tau}_{p}(h_{n})-h_{n}^{p+1}\widehat{B}_{n,p,q}(h_{n},b_{n}),$
where $\widehat{B}_{n,p,q}(h_{n},b_{n})$ is the estimated bias term $B_{p}(h_{n})$. The confidence interval they construct based on the bias-corrected estimator takes the following form:
$\widehat{\text{CI}}_{1-\alpha}^{\text{bc}}(h_{n},b_{n})=\left(\widehat{\tau}^{\mathrm{bc}}_{p,q}(h_{n},b_{n})\pm z_{1-\alpha/2}\widehat{\sigma}(h_{n})\right),$ (18)
where $z_{1-\alpha/2}$ is the critical value from the standard normal distribution, and $\widehat{\sigma}(h_{n})$ is the standard error (for which Calonico et al. (2014) provide two options: plug-in and fixed matches). The above CI construction is implemented in the rdrobust package, with $p=1$ and $q=2$ as default.
The CI constructed in (18) might not perform well in finite samples because the estimator of the bias introduces additional variability. To address this, Calonico et al. (2014) also propose an alternative, more robust CI construction using the bias-corrected estimator $\widehat{\tau}^{\mathrm{bc}}_{p,q}(h_{n},b_{n})$ as a starting point. They construct this robust CI by inflating the CI in (18) with a more conservative estimator of the standard error:
$\widehat{\text{CI}}_{1-\alpha}^{\text{rbc}}(h_{n},b_{n})=\left(\widehat{\tau}^{\mathrm{bc}}_{p,q}(h_{n},b_{n})\pm z_{1-\alpha/2}\sqrt{\widehat{V}^{\text{bc}}_{n,p,q}}\right),$ (19)
where $\widehat{V}^{\text{bc}}_{n,p,q}$ is an estimator of the asymptotic variance that accounts for both the variability in estimation and the bias-correction. In Tables 1, 2, 3 and 4, the columns marked as b.c. and rob correspond to the confidence intervals in (18) and (19), respectively.
A.2 Bias-aware confidence intervals
The bias-aware inference approach (e.g., Armstrong and Kolesár, 2018, 2020; Imbens and Wager, 2019) provides a different approach for obtaining confidence intervals for $\tau_{c}$. Note that the LLR estimator takes the form $\widehat{\tau}_{\texttt{LLR}}=\sum_{i=1}^{n}\gamma_{i,\texttt{LLR}}Y_{i}$ for some weights $\gamma_{i,\texttt{LLR}}$ that are functions of the running variable. In fact, the same also holds more generally for the estimator $\widehat{\tau}_{p}(h_{n})$ defined in Section A.1. Next, one makes the observation that although $\widehat{\tau}_{\texttt{LLR}}$ is motivated by a regression problem, the rest of our inference procedure hinges upon on properties that hold for any linear estimators of the form $\widehat{\tau}(\gamma)=\sum_{i=1}^{n}\gamma_{i}Y_{i}$. Given by this observation, we can systematically do better by directly finding weights $\gamma_{i}$ that minimize the mean squared error (MSE) of the estimator $\widehat{\tau}(\gamma)$ uniformly over a class $\mathcal{M}$ of conditional response functions $\mu_{w}$, i.e.,
$\widehat{\gamma}:=\mathop{\rm argmin}_{\gamma}\sup_{\mu_{w}\,\in\,\mathcal{M}}\text{MSE}\left(\widehat{\tau}(\gamma)\mid\{X_{i}\}_{i=1}^{n}\right).$ (20)
The two most widely considered classes to date (Armstrong and Kolesár, 2018; Kolesár and Rothe, 2018; Imbens and Wager, 2019) are
$\mathcal{M}_{\text{T2},\,M}=\left\{\mu_{w}(\cdot):|\mu_{w}(x)-\mu_{w}(c)-\mu_{w}^{\prime}(c)x|\leq M(x-c)^{2}/2\text{ for every }x,\text{ and }w=0,1\right\},$ (21)
$\mathcal{M}_{\text{H2},\,M}=\left\{\mu_{w}(\cdot):\mu^{\prime}_{w}(x)\text{ is $M$-Lipschitz for all $x$ and }w=0,1\right\},$ (22)
where $M>0$ is a constant. The first option, considered in early work by Armstrong and Kolesár (2018), justifies accuracy of a 2nd-order Taylor expansion at $c$ and allows for explicit, closed-form solutions to the minimax linear estimation problem. The second option, which simply takes all functions to belong to a 2nd-order Hölder class, was used by Kolesár and Rothe (2018) and Imbens and Wager (2019) and also adopted for the RDHonest package by Armstrong and Kolesár (2018).
Fixing a value of $M$, we numerically solve the optimization problem (20) and obtain an upper bound on the worst-case bias of the resulting estimator as a by-product. We then follow Imbens and Manski (2004) to construct a bias-aware confidence interval by inflating the Gaussian confidence interval accounting for the worst-case bias of $\tau(\widehat{\gamma})$ under the function class $\mathcal{M}$. These bias-aware confidence intervals provide uniformly valid coverage in finite samples, provided the function class $\mathcal{M}$ and the distribution of $Y_{i}$ given $X_{i}$ is Gaussian (Armstrong and Kolesár, 2018). [^12] If the conditional distribution of $Y_{i}$ is non-Gaussian, one can recover exact finite-sample results using the bias-aware approach by using finite-sample concentration results instead of Gaussian approximation. In practice, however, smoothness constant $M$ is typically unknown; both Armstrong and Kolesár (2020) and Imbens and Wager (2019) suggest using some rule-of-thumb estimates.
A.3 The choice of function class
One of the key insights our paper provides is that the choice of the function class $\mathcal{M}$ in the minimax problem (8) plays a crucial role in determining the conservativeness (and thus the practical usefulness) of the resulting bias-aware method. The function classes used in the existing bias-aware literature (e.g., the classes $\mathcal{M}_{\text{T2},\,M}$ and $\mathcal{M}_{\text{H2},\,M}$ defined in (21) and (22) respectively) are much modest in terms of assumed smoothness as compared to the the bias-corrected methods used in practice (see, e.g., Calonico et al. (2014, Theorem 1)). Here, on the other hand, use the class of all functions $\mu_{w}(\cdot)$ that satisfy Assumption 2 and for which $\mu_{0}^{\prime\prime}(\cdot)$ is $\widehat{B}$-Lipschitz, where we obtain $\widehat{B}$ via a global cubic regression and cross-fitting. While this might appear as a minor modification from a theoretical point of view, it materially affects the length of our intervals in practice, as we demonstrate in Tables 1, 2, 3 and 4.
To illustrate the difference between the conservativeness of our function class versus the function classes used by existing bias-aware approaches, we plot in Fig. 3 the regression functions corresponding to the worst-case bias obtained by our method and $\mathcal{M}_{\text{H2},\,M}$ as used in RDHonest. This plot demonstrates that the classes $\mathcal{M}_{\text{H2},\,M}$ can often be unreasonably broad, and include regression functions that are ex ante implausible or rarely occur in practice. This explains why the resulting bias-aware confidence intervals can be substantially more conservative than the bias-corrected ones (see, e.g.Table 3). The empirical success of our PLRD method that we demonstrate in Section 4 is, in fact, rooted in this choice of the function class over which one computes the worst-case MSE.
Figure 3: Functions $\mu_{w}(\cdot)$ with worst-case bias in the RDHonest and PLRD function classes. The top two plots use the $m_{\text{quad}}$ example from Imbens and Kalyanaraman (2012), while the bottom two plots use the data from Lee (2008). In each plot, black lines show $\mu_{w}(\cdot)$ functions with maximum (solid) or minimum (dotted) bias in the respective function class. This figure illustrates how the RDHonest function class can be too modest for many regression discontinuity designs in practice.
A.4 Ablation analysis
Here we present an empirical comparison of our PLRD method with other alternative methods. We summarize our empirical results in Table 5, which complements the results in Table 1 of the main paper. Table 5 compares our PLRD method with four other alternative approaches:
  1. (a)
    Bias-aware intervals for local linear regression following Armstrong and Kolesár (2020):
    1. (i)
      Fit a local linear regression using the rdrobust package, with default bandwidth selection rule.
    2. (ii)
      Obtain the LLR estimator $\widehat{\tau}_{\texttt{LLR}}$ with its conventional standard error estimate $\widehat{\text{SE}}_{\texttt{LLR}}$
    3. (iii)
      Build the confidence interval $\widehat{\tau}_{\texttt{LLR}}\pm 2.181\cdot\widehat{\text{SE}}_{\texttt{LLR}}$ (Armstrong and Kolesár, 2020).
  2. (b)
    Bias-aware intervals for local quadratic regression following Armstrong and Kolesár (2020):
    1. (i)
      Fit a local quadratic regression using the rdrobust package (i.e., with $p=2$ in Section A.1), with default bandwidth selection rule.
    2. (ii)
      Obtain the LQR estimator $\widehat{\tau}_{\texttt{LQR}}$ (defined as $\widehat{\tau}_{2}(h_{n})$ in Section A.1) with its conventional standard error estimate $\widehat{\text{SE}}_{\texttt{LQR}}$
    3. (iii)
      Build the confidence interval $\widehat{\tau}_{\texttt{LQR}}\pm 2.113\cdot\widehat{\text{SE}}_{\texttt{LQR}}$ (Armstrong and Kolesár, 2020).
  3. (c)
    PLRD without Assumption 2: Bias-aware inference operating under the class
    $\mathcal{M}_{\text{H3},\,B}=\left\{\mu_{w}(\cdot):\mu^{\prime\prime}_{w}(x)\text{ is $M$-Lipschitz for all $x$ and }w=0,1\right\}.$
    This method uses the additional smoothness just as PLRD, while allowing different curvatures around the treatment threshold.
    1. (i)
      Estimate $B$ using a global cubic regression on either side of the threshold, in the same manner as we estimate $B$ via global cubic regression in Proposition 2.
    2. (ii)
      Follow Armstrong and Kolesár (2018) and Imbens and Wager (2019) to solve the minimax problem (8) with the function class $\mathcal{M}_{\text{H3},\,\widehat{B}}$ via numerical convex optimization and compute the minimax linear estimator for $\tau_{c}$ within this new class.
    3. (iii)
      As a by-product of the optimization problem, obtain a bound on the worst-case bias of our estimator within this class, and construct “bias aware” confidence intervals for $\tau_{c}$ using the Imbens and Manski (2004) construction.
    Algorithmically, this is similar to our Algorithm 1; we have also implemented this as a part of our plrd package.
  4. (d)
    PLRD without Assumption 1: Bias-aware intervals operating under the intersection of the class $\mathcal{M}_{\text{H2},\,M}$ and the class of regression functions satisfying Assumption 2. In essence, this method combines optrdd (Imbens and Wager, 2019) with Assumption 2.
    1. (i)
      Here we estimate $M$ using a global quadratic regression on either side of the threshold, in the same manner as we estimate $B$ via global cubic regression in Proposition 2.
    2. (ii)
      Follow Armstrong and Kolesár (2018) and Imbens and Wager (2019) to solve the minimax problem (8) with the above function class via numerical convex optimization and compute the minimax linear estimator for $\tau_{c}$ within this new class.
    3. (iii)
      Obtain an upper bound on the worst-case bias of our estimator within this class, and construct “bias aware” confidence intervals for $\tau_{c}$ using the Imbens and Manski (2004) construction.
    Algorithmically, this is similar to the optrdd package by Imbens and Wager (2019). We also provide an implementation of this method as a part of our replication files.
Dataset Alt. (a) Alt. (b) Alt. (c) Alt. (d) plrd
Lee (2008) 0.040 0.047 0.041 0.038 0.035
(96.5%) (96.3%) (97.8%) (99.8%) (98.3%)
Cattaneo et al. (2015) 5.956 7.267 5.259 9.376 4.525
(96.6%) (96.4%) (97.3%) (99.7%) (97.3%)
Matsudaira (2008) (read.) 0.068 0.070 0.071 0.084 0.060
(96.1%) (96.7%) (97.1%) (96.8%) (97.1%)
Matsudaira (2008) (math) 0.050 0.056 0.059 0.073 0.051
(94.4%) (95.7%) (97.3%) (95.2%) (97.0%)
Jacob and Lefgren (2004) (read.) 0.075 0.092 0.078 0.075 0.065
(96.8%) (96.5%) (96.0%) (99.3%) (94.6%)
Jacob and Lefgren (2004) (math) 0.087 0.102 0.082 0.077 0.069
(95.6%) (94.9%) (97.2%) (96.2%) (95.0%)
Oreopoulos (2006) 0.218 0.401 0.225 0.185 0.194
(95.2%) (96.9%) (98.6%) (94.5%) (96.1%)
Ludwig and Miller (2007) 4.516 5.564 4.075 3.846 3.542
(97.2%) (96.5%) (97.6%) (98.2%) (97.4%)
Meyersson (2014) 7.079 8.796 5.905 7.837 5.095
(95.7%) (94.7%) (96.7%) (98.5%) (97.3%)
Pure noise simulation 1.098 1.375 1.612 1.179 0.847
(96.6%) (98.6%) (97.3%) (98.2%) (96.3%)
Simulation Setting 1 0.223 0.278 0.276 0.202 0.235
(92.7%) (97.5%) (97.2%) (92.5%) (94.6%)
Simulation Setting 2 0.297 0.289 0.375 0.493 0.309
(86.4%) (86.1%) (97.7%) (98.2%) (51.5%)
Simulation Setting 3 0.236 0.279 0.280 0.300 0.239
(93.3%) (96.2%) (97.1%) (98.2%) (95.9%)
Simulation Setting 4 0.223 0.276 0.279 0.255 0.236
(93.6%) (96.5%) (96.7%) (98.1%) (97.5%)
Table 5: Comparison of PLRD with other alternatives: first two columns correspond to inflating conventional confidence intervals: instead of $1.96$, we multiply the standard error by a factor of (a) $2.181$ or (b) $2.113$, as suggested by Armstrong and Kolesár (2020); the next two columns correspond to (c) PLRD minus Assumption 2, and (d) PLRD minus Assumption 1; the last column corresponds to our PLRD method. We write the empirical coverages in parentheses and the average width in the rows. Other details of this table are same as for Tables 1, 2 and 4 in the main text.
The empirical results demonstrate that the rule-of-thumb methods (AK20 (a) and AK20 (b)) achieve valid coverage but at the cost of wider intervals, mirroring our earlier findings in Tables 1, 2 and 4. The relaxed variants of PLRD offer insights into the role of each assumption: Dropping either Assumption 1 or Assumption 2 generally leads to more conservative intervals, suggesting that both assumptions contribute to the efficiency of PLRD.
Appendix B Additional details on our real data experiments
Here we provide detailed descriptions of the real datasets analyzed in the main paper. We reference each dataset by the paper that introduced it, same as in Tables 1, 3 and 5.
  1. 1.
    Lee (2008): This dataset, frequently used for illustrating regression discontinuity methods, examines the incumbency advantage in U.S. House elections. The running variable is the margin of victory, defined as the difference in vote share between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$. The outcome of interest is the probability of winning subsequent elections. The dataset consists of $6,558$ observations from U.S. House elections between 1946 and 1998. Lee (2008) reported a significant positive effect of incumbency on the probability of winning subsequent elections, indicating an incumbency advantage; see Table 1 and Figure 2 of Lee (2008) for more details.
  2. 2.
    Cattaneo et al. (2015): This is similar to Lee (2008) except that they study the incumbency advantage in U.S. Senate elections instead of House elections. The running variable is the margin of victory, defined as the difference in vote percentages (ranging from $-100$ to $100$) between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$. The outcome of interest is the percentage of winning subsequent elections. The dataset consists of $1,390$ observations ($1,297$ excluding missing outcomes) from U.S. House elections between 1914 and 2010. Our findings align with the original paper Cattaneo et al. (2015) as well as the replication in Cattaneo et al. (2024).
  3. 3.
    Matsudaira (2008): This dataset records fifth-grade students’ math and reading scores in an unspecified large school district in the northeast from 2001 to 2002. Matsudaira (2008) examined the impact of mandatory summer school on students’ performance. The running variable is the minimum of students’ test scores in math and reading, with a score below the passing threshold for either math or reading determining mandatory attendance in the summer school. The primary outcomes are post-summer school reading and math test scores. We use a version of this dataset shared by the author with K. Kalyanamaran and G. Imbens in 2008. This dataset has a slightly larger sample size $n=68,798$ than the one used in the original paper ($n=66,839$). Our findings mirror the results in Matsudaira (2008, Table 4) for fifth-grade students.
  4. 4.
    Jacob and Lefgren (2004): This dataset contains the record of third-grade student achievement in Chicago Public Schools from 1997 to 1999. Jacob and Lefgren (2004) investigated the effectiveness of remedial education programs on student achievement, using a regression discontinuity design based on test score. We use the same dataset as Imbens and Zajonc (2009), which contains $70,831$ observations after dropping entries with missing information. The running variable is the minimum of the math and reading score for an exam in the spring and the outcome is the post-remedial program reading (resp. math) test scores, reported as Rasch test scores rather than grade equivalents. Rasch test scores are normalized scores (zero mean, unit s.d.) for the population of third-grade students from 1997 to 1999. Our results are aligned with the findings in Jacob and Lefgren (2004) for third-grade students.
  5. 5.
    Oreopoulos (2006): This study concerns the impact of compulsory schooling laws on educational attainment. The running variable is the individual’s year of birth, with thresholds determined by changes in compulsory schooling laws. The outcome here is years of completed education. Our dataset comprises $73,954$ observations, differing slightly from the original study due to variations in preprocessing; again, we use the same dataset also used in Kolesár and Rothe (2018) and focus on the reduced form analysis rather than a fuzzy RD approach (see Table 1, Panel C of the corrigendum to Oreopoulos (2006)). Our findings are in line with Table 2 of Kolesár and Rothe (2018).
  6. 6.
    Ludwig and Miller (2007): This study examines the effectiveness of the Head Start program on children’s health and schooling. The running variable is the 1960 poverty rate, and the threshold corresponds to the qualifying cutoff for program participation. Following Calonico et al. (2014), we consider as outcome variable the mortality rates per 100,000 for children between 5 and 9 years old, with Head Start-related causes, for 1973–1983. After cleaning missing values, this dataset includes $2,783$ observations. Table III and Figure IV (Panel A) in Ludwig and Miller (2007) illustrate that the Head Start program had a negative impact on the outcomes.
  7. 7.
    Meyersson (2014): This study investigates the effect of Islamic party control in local governments on women’s rights, in particular on the educational attainment of young women, focusing on Turkey’s 1994 mayoral elections. The study employs an RD design where the running variable is the Islamic margin of victory—the vote share gap between the leading Islamic and secular parties, thus with a natural cutoff of zero. The analysis is conducted at the municipality level and has $2,629$ observations. We replicate part of the study where the primary outcome is the school attainment for women who were (potentially) in high school during the period 1994-2000, measured with variables extracted from the 2000 census. This data have been used as a running example in Cattaneo et al. (2024). Our results align with the illustrations in Cattaneo et al. (2024) and the original paper Meyersson (2014) — in municipalities where the Islamic party barely won, the educational attainment of women is roughly 3% higher than in municipalities where the party barely lost.
Appendix C Implementation via Convex Optimization
In this section, we discuss how one can solve the optimization problem (8) using readily available software described in, e.g., Boyd and Vandenberghe (2004). Essentially, our task is to solve quadratic minimization problems of the form: [^13] Here, the $\sigma_{i}^{2}$ should simply be taken as observation-level weights in an optimization problem. If we set $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$ to be the conditional variance of the response $Y_{i}$ given the running variable $X_{i}$, then the problem below is exactly the minimax linear estimation problem. When we run this the optimization as specified in Algorithm 1, we set $\sigma_{i}^{2}=\widehat{\sigma}^{2}$ via cross-fitting, as described in Section 2.
$\mathop{\rm{minimize}}_{\gamma}\left[\sup_{\mu_{w}\,\in\,\mathcal{M}_{B}}\left(\sum_{i=1}^{n}\gamma_{i}\mu_{W_{i}}(X_{i})-(\mu_{1}(c)-\mu_{0}(c))\right)^{2}+\sum_{i=1}^{n}\sigma_{i}^{2}\gamma_{i}^{2}\right].$ (23)
Here, we take $\mathcal{M}_{B}$ of all conditional response functions $\mu_{w}(\cdot)$ satisfying Assumption 2 and such that $\mu_{0}^{\prime\prime}(x)$ is globally $B$-Lipschitz. Since $\gamma_{i}$ are measurable functions of $\{X_{i}\}_{i=1}^{n}$, the bias of the estimator $\widehat{\tau}(\gamma)=\sum_{i=1}^{n}\gamma_{i}Y_{i}$ conditional on the running variable is given by
$\mathbb{E}\left[\sum_{i=1}^{n}\gamma_{i}Y_{i}\mid\{X_{i}\}_{i=1}^{n}\right]-\tau_{c}$ $=\sum_{i=1}^{n}\gamma_{i}\mu_{W_{i}}(X_{i})-\tau_{c}$
$=\sum_{i=1}^{n}\gamma_{i}(W_{i}(\tau_{c}+\beta(X_{i}-c))+\mu_{0}(X_{i}))-\tau_{c}$
$=\sum_{i=1}^{n}\gamma_{i}\left(\mu_{0}(X_{i})-\mu_{0}(c)-\mu_{0}^{\prime}(c)(X_{i}-c)-\frac{1}{2}\mu_{0}^{\prime\prime}(c)(X_{i}-c)^{2}\right),$
provided
$\sum_{i=1}^{n}\gamma_{i}=0,\ \sum_{i=1}^{n}\gamma_{i}W_{i}=1,\ \sum_{i=1}^{n}\gamma_{i}W_{i}(X_{i}-c)=0,\ \sum_{i=1}^{n}\gamma_{i}(1-W_{i})(X_{i}-c)=0,\ \sum_{i=1}^{n}\gamma_{i}(X_{i}-c)^{2}=0,$
which are implicit constraints in our optimization procedure. Indeed, at values of $\gamma$ for which these constraints are not all satisfied, we can construct $\mu_{w}(\cdot)\in\mathcal{M}_{B}$ that forces the squared conditional bias to blow up to $\infty$. Thus, the solution $\widehat{\gamma}$ to (23) must satisfy the constraints. Next, define
$\rho(x):=\frac{1}{B}\left(\mu_{0}(x+c)-\mu_{0}(c)-\mu_{0}^{\prime}(c)x-\frac{1}{2}\mu_{0}^{\prime\prime}(c)x^{2}\right).$
[…]
$\widehat{\gamma}:=\mathop{\rm{minimize}}_{\gamma}\left[B^{2}\left(\sup_{\rho\in\mathcal{R}}\sum_{i=1}^{n}\gamma_{i}\rho(X_{i}-c)\right)^{2}+\sum_{i=1}^{n}\sigma_{i}^{2}\gamma_{i}^{2}\right].$
Finally, we make the implicit constraints explicit and use an auxiliary variable $t$ and rewrite the above optimization problem as:
$(\widehat{\gamma},\widehat{t}\,):=$ $\mathop{\rm{minimize}}_{\gamma,t}\left\{B^{2}t^{2}+\sum_{i=1}^{n}\gamma_{i}^{2}\sigma_{i}^{2}\right\}$
$\text{subject to }\ \sup_{\rho\in\mathcal{R}}\sum_{i=1}^{n}\gamma_{i}\rho(X_{i}-c)\leq t,\ \sum_{i=1}^{n}\gamma_{i}W_{i}=1,\ \sum_{i=1}^{n}\gamma_{i}\left(1-W_{i}\right)=-1,$ (24)
$\qquad\qquad\ \sum_{i=1}^{n}\gamma_{i}(X_{i}-c)=0,\ \sum_{i=1}^{n}\gamma_{i}(1-W_{i})(X_{i}-c)=0,\ \sum_{i=1}^{n}\gamma_{i}(X_{i}-c)^{2}=0.$
An immediate consequence of this modified optimization problem is that we get an upper bound on the conditional bias of our PLRD estimator. Next, we discuss how one can solve the above problem using convex duality. The dual problem is given by
$\mathop{\rm{maximize}}_{\rho(\cdot),\lambda}\inf_{\gamma,t}\left\{\sum_{i=1}^{n}\gamma_{i}^{2}\sigma_{i}^{2}+B^{2}t^{2}\right.$
$\qquad\qquad\qquad\quad+\lambda_{1}\left(\sum_{i=1}^{n}\gamma_{i}\rho(X_{i}-c)-t\right)+\lambda_{2}\left(\sum_{i=1}^{n}\gamma_{i}W_{i}-1\right)+\lambda_{3}\left(\sum_{i=1}^{n}\gamma_{i}\left(1-W_{i}\right)+1\right)$
$\qquad\qquad\qquad\quad\left.+\lambda_{4}\sum_{i=1}^{n}\gamma_{i}W_{i}(X_{i}-c)+\lambda_{5}\sum_{i=1}^{n}\gamma_{i}\left(1-W_{i}\right)(X_{i}-c)+\lambda_{6}\sum_{i=1}^{n}\gamma_{i}(X_{i}-c)^{2}\right\}$
$\text{subject to }\lambda_{1}\geq 0,\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5},\lambda_{6}\in\mathbb{R},\text{and},\rho\in\mathcal{R}.$
The inner minimization problem is simply a quadratic one, which we can explicitly solve to reduce the dual problem to the following.
$\widehat{\theta}=\mathop{\rm argmin}_{\theta=(\widetilde{\rho},\lambda)\in\Theta}\left(\frac{1}{4n}\sum_{i=1}^{n}\frac{G^{2}(\theta;X_{i}-c,W_{i})}{\sigma_{i}^{2}}+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n}\right),$ (25)
where $\theta=(\widetilde{\rho},\lambda)$ denotes the dual variable that varies in the set
$\Theta:=\left\{(f,\lambda)\mid f:[-\ell,\ell]\to\mathbb{R},f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is $\lambda_{1}$-Lipschitz},\lambda\in[0,\infty)\times\mathbb{R}^{5}\right\},$ (26)
and
$G(\widetilde{\rho},\lambda;x,w):=\widetilde{\rho}(x)+\lambda_{2}w+\lambda_{3}(1-w)+\lambda_{4}wx+\lambda_{5}(1-w)x+\lambda_{6}x^{2}.$
The function $\widetilde{\rho}(\cdot)$ in the above formulation corresponds to the function $\lambda_{1}\rho(\cdot)$ in the original problem. Thus, we can recover the optimal variables $(\widehat{\gamma},\widehat{t}\,)$ for the primal problem (24) from the dual optimal variable $\widehat{\theta}$ (i.e., solution of (25)), as follows.
$\widehat{\gamma}_{i}=-\frac{1}{2\sigma_{i}^{2}}G(\widehat{\theta};X_{i}-c,W_{i}),\quad\widehat{t}=\frac{1}{2B^{2}}\widehat{\lambda}_{1}.$
Since our optimization problem differs from Imbens and Wager (2019) essentially in the choice of the function class, Imbens and Wager (2019, Proposition 2) provides a basis for approximating (25) using a finite-dimensional quadratic program, in a way that the resulting discrete solutions converge uniformly to the continuous solution as the grid size becomes small. In our software implementation, we adopt this finite-dimensional approximation approach to compute the optimal weights using quadprog, a function for quadratic programming in R (R Core Team, 2023).
Appendix D Main proofs
Proof of Lemma 4.
For simplicity in notation, we assume without loss of generality that the regression discontinuity threshold is $c=0$. We split the proof into some auxiliary results, namely, Lemmas 10, 8, 7, 9 and 11 in Appendix E. First, using tools from empirical process theory (see van der Vaart and Wellner (1996) for a book-level treatment), we prove a uniform concentration bound for the loss functions of the empirical and population dual problems. Recall that the dual parameter space is defined as
$\Theta(\ell):=\left\{(f,\lambda)\mid f:[-\ell,\ell]\to\mathbb{R},\,f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is $\lambda_{1}$-Lipschitz},\lambda\in[0,\infty)\times\mathbb{R}^{5}\right\}.$
Define a neighborhood $\Theta_{n}(\ell,\delta,\varepsilon)$ of the dual optimizer $\theta_{n}^{*}$, as follows:
$\Theta_{n}(\ell,\delta,\varepsilon):=\{(\rho,\lambda)\mid(\rho,\lambda)\in\Theta(\ell),\ |\lambda_{1}-\lambda_{n,1}^{*}|\leq\delta\lambda_{n,1}^{*}\text{ and }\|\lambda_{-1}-\lambda^{*}_{n,-1}\|_{2}\leq\varepsilon\|\lambda_{n,-1}^{*}\|_{2}\},$
where $\ell,\delta,\varepsilon\in(0,1)$ are fixed. We show in Lemma 7 the following maximal inequality.
$\mathbb{E}\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})|\lesssim(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\|\lambda_{n,-1}^{*}\|_{2}^{2}\frac{\sqrt{\varepsilon}}{\sqrt{n}}+\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\frac{\ell^{3}}{\sqrt{n}}.$
Next, Lemma 9 simplifies the above bound to the following.
$\mathbb{E}\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})|\lesssim(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}.$ (27)
On the other hand, Lemma 9 tells us that
$\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},B)|\lesssim\frac{(\lambda_{n,1}^{*})^{2}}{n}\left|\frac{1}{\widehat{B}}-\frac{1}{B}\right|.$
This, in conjunction with (27) tells us that for every $\eta\in(0,1)$, there exists $C=C(\eta)>0$ such that, for all sufficiently large $n$,
$\mathbb{P}\left(\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\sigma^{2},B)|\leq C\left((\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}+\frac{(\lambda_{n,1}^{*})^{2}}{n}\right)\right)\geq 1-\eta.$ (28)
We also show in Lemma 8 that
$L_{n}(\theta;\sigma^{2},B)-L_{n}(\theta_{n}^{*};\sigma^{2},B)\geq\frac{\alpha}{4\sigma^{2}}\left(\|\rho-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\lambda_{-1}-\lambda_{n,-1}^{*}\|_{2}^{2}\right)+\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}-R_{n},$ (29)
for some $\alpha\in(0,1)$, where $0\leq R_{n}\leq\ell^{3}(\lambda_{1}+\lambda_{n,1}^{*})\|\lambda_{-1}-\lambda_{n,-1}^{*}\|_{2}/2\sigma^{2}$. Note, for $(\rho,\lambda)\in\Theta_{n}(\ell,\delta,\varepsilon)$,
$R_{n}\leq\frac{1}{2\sigma^{2}}\ell^{3}(2+\delta)\lambda_{n,1}^{*}\varepsilon\|\lambda_{n,-1}^{*}\|_{2}\leq\frac{3}{2\sigma^{2}}\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\ell^{3}\varepsilon.$ (30)
Next, we show in Lemma 10 that for every $\delta,\eta\in(0,1)$, and $\varepsilon=O(n^{-1/4})$, it holds that
$\mathbb{P}\left(\widehat{\theta}_{n}\in\Theta_{n}(\ell,\delta,\varepsilon)\right)\geq 1-\eta,$
for all sufficiently large $n$. The following inequalities hold whenever the event in (28) holds as well as $\widehat{\theta}_{n}\in\Theta_{n}(\ell,\delta,\varepsilon)$ with $\varepsilon=O(n^{-1/4})$, i.e., the following inequalities hold with probability at least $1-2\eta$ for all sufficiently large $n$.
$\frac{\alpha}{4\sigma^{2}}\left(\|\widehat{\rho}_{n}-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\widehat{\lambda}_{n,-1}-\lambda_{n,-1}^{*}\|_{2}^{2}\right)+\frac{(\widehat{\lambda}_{n,1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$\leq\frac{3}{2\sigma^{2}}\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\ell^{3}\varepsilon+L_{n}(\widehat{\theta};\sigma^{2},B)-L_{n}(\theta_{n}^{*};\sigma^{2},B)$ (using (29) and (30))
$\leq\frac{3}{2\sigma^{2}}\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\ell^{3}\varepsilon+\underbrace{\widehat{L}_{n}(\widehat{\theta};\widehat{\sigma}^{2},\widehat{B})-\widehat{L}_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},\widehat{B})}_{\leq 0}+2\sup_{\theta\in\Theta_{n}(\ell,\delta,\varepsilon)}\left|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\sigma^{2},B)\right|$ (using triangle inequality and the definition of $\widehat{\theta}_{n}$)
$\lesssim\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\frac{\ell^{3}}{n^{1/4}}+(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{1}{n^{5/2+1/8}}+\frac{(\lambda_{n,1}^{*})^{2}}{n}.$ (using (28))
Finally, Lemma 9 tells us that $\|\lambda_{n,-1}^{*}\|_{2}=O(\lambda_{n,1}^{*}\ell^{3}+n^{-1})$, while Lemma 11 tells us that $\lambda_{n,1}^{*}=\mathcal{O}_{p}(n^{-3/7})$. This simplifies the above upper bound to the following.
$\frac{\alpha}{4\sigma^{2}}\left(\|\widehat{\rho}_{n}-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\widehat{\lambda}_{n,-1}-\lambda_{n,-1}^{*}\|_{2}^{2}\right)+\frac{(\widehat{\lambda}_{n,1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$\qquad\qquad\qquad\lesssim_{\eta}(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{n^{1/4}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{(\lambda_{n,1}^{*})^{2}}{n}$
$\qquad\qquad\qquad\lesssim_{\eta}\frac{\ell^{6}}{n^{1/4+6/7}}+\frac{\ell^{3}}{n^{3/2+3/7}}+\frac{1}{n^{1+6/7}},$
with probability at least $1-2\eta$, for all sufficiently large $n$. Finally, we can ignore the last term in the above display, since $\ell$ satisfies (5). This completes the proof. $\square$
Proof of Corollary 5.
For the PLRD weights $\widehat{\gamma}_{i}$ derived in Algorithm 1, we now show that
$\frac{\|\widehat{\gamma}\|_{\infty}}{\|\widehat{\gamma}\|_{2}}=\frac{\max_{1\leq i\leq n}|\widehat{\gamma}_{i}|}{(\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2})^{1/2}}\rightarrow_{p}0.$
Assume without loss of generality that the threshold for treatment is at $c=0$. Recall that we use sample splitting in Algorithm 1 and write $\widehat{\gamma}_{i}=\widehat{\gamma}_{(k),i}/2$ if $i\in I_{k}$ ($k=1,2$). Thus,
$\frac{\|\widehat{\gamma}\|_{\infty}}{\|\widehat{\gamma}\|_{2}}=\frac{\max_{1\leq i\leq n}|\widehat{\gamma}_{i}|}{(\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2})^{1/2}}\lesssim\frac{\|\widehat{\gamma}_{(1)}\|_{\infty}}{\|\widehat{\gamma}_{(1)}\|_{2}}+\frac{\|\widehat{\gamma}_{(2)}\|_{\infty}}{\|\widehat{\gamma}_{(2)}\|_{2}}.$
Since $\widehat{\gamma}_{(1)}$ and $\widehat{\gamma}_{(2)}$ are identically distributed, it suffices to show the following.
$\frac{\|\widehat{\gamma}_{(1)}\|_{\infty}}{\|\widehat{\gamma}_{(1)}\|_{2}}=\frac{\max_{i\in I_{1}}|\widehat{\gamma}_{(1),i}|}{(\sum_{i\in I_{1}}\widehat{\gamma}_{(1),i}^{2})^{1/2}}\rightarrow_{p}0.$ (31)
Towards showing (31), note that by Cauchy-Schwarz inequality,
$|I_{1}|\sum_{i\in I_{1}}\widehat{\gamma}_{(1),i}^{2}=\sum_{i\in I_{1}}(2W_{i}-1)^{2}\sum_{i\in I_{1}}\widehat{\gamma}_{(1),i}^{2}\geq\bigg{(}\sum_{i\in I_{1}}(2W_{i}-1)\widehat{\gamma}_{(1),i}\bigg{)}^{2}=4.$
This yields the bound $\|\widehat{\gamma}_{(1)}\|_{2}\gtrsim n^{-1/2}$. For the numerator in (31) use triangle inequality to write
$\sqrt{n}\max_{i\in I_{1}}|\widehat{\gamma}_{(1),i}|$ $\leq\sqrt{n}\max_{i\in I_{1}}|\gamma_{(1),i}^{*}|+\sqrt{n}\max_{i\in I_{1}}|\widehat{\gamma}_{(1),i}-\gamma_{(1),i}^{*}|,$
where $\gamma_{(1),i}^{*}:=-(\rho_{n}^{*}(X_{i})+\lambda_{n,-1}^{*}v(X_{i},W_{i}))/2\widehat{\sigma}_{(2)}^{2}$, with $v(x,w)=(w,1-w,xw,x(1-w),x^{2})^{\top}$, and
$(\rho_{n}^{*},\lambda_{n}^{*}):=\mathop{\rm argmin}_{(\rho,\lambda)\in\Theta}\frac{1}{4\sigma^{2}_{(2)}}\mathbb{E}\left[G^{2}(\rho,\lambda;X,W)\right]+\frac{\lambda_{1}^{2}}{4nB^{2}_{(2)}}+\frac{\lambda_{2}-\lambda_{3}}{n},$
with $\Theta$ as defined in (26). For any $(\rho,\lambda)\in\Theta$, $\rho^{(j)}(0)=0$ for $j=0,1,2$, and $\rho^{\prime\prime}$ is $\lambda_{n,1}^{*}$-Lipschitz, so a Taylor expansion of ${\rho}(x)$ around $x=0$ gives us
$|{\rho}(x)|=|\rho(x)-\rho(0)-\rho^{\prime}(0)x|=\frac{|\rho^{\prime\prime}(\xi)-\rho^{\prime\prime}(0)|}{|\xi-x|}\left|\xi-x\right|\frac{x^{2}}{2}\leq\lambda_{n,1}^{*}\frac{\ell^{3}}{2}.$ (32)
Consequently,
$\max_{i\in I_{1}}|\gamma_{(1),i}^{*}|\lesssim\sup_{x}\left|\rho_{n}^{*}(x)+v(x,1(x>0))^{\top}\lambda_{n,-1}^{*}\right|\lesssim\lambda_{n,1}^{*}\ell^{3}+\|\lambda_{n,-1}^{*}\|_{2},$
where we used (32) to bound $\sup_{x}|\rho_{n}^{*}(x)|$ and Cauchy-Schwarz inequality to bound the other term. Since $B_{(2)}$ and $\sigma_{(2)}^{2}$ are independent of the data in the fold $I_{1}$, Lemma 9 applies here and tells us that $\|\lambda_{n,-1}^{*}\|_{2}=O(\lambda_{n,1}^{*}\ell^{3}+n^{-1})$. Consequently, we have
$\sqrt{n}\max_{i\in I_{1}}|\gamma_{(1),i}^{*}|\lesssim\sqrt{n}\left(\lambda_{n,1}^{*}\ell^{3}+n^{-1}\right).$
On the other hand, $\max_{i\in I_{1}}|\widehat{\gamma}_{(1),i}-\gamma_{(1),i}^{*}|\lesssim\|\widehat{\rho}_{n}-\rho_{n}^{*}\|_{\infty}+\|\widehat{\lambda}_{n,-1}-\lambda_{n,-1}^{*}\|_{2}$. It follows from (33) that $\|\widehat{\rho}_{n}-\rho_{n}^{*}\|_{\infty}\leq(\widehat{\lambda}_{n,1}+\lambda_{n,1}^{*})\ell^{3}$. Applying Lemma 4, we deduce that
$\|\widehat{\lambda}_{n,-1}-\lambda_{n,-1}^{*}\|_{2}=\mathcal{O}_{p}\left(\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{1/8}}+\sqrt{\lambda_{n,1}^{*}}\frac{\ell^{3/2}}{\sqrt{n}}+\frac{\lambda_{n,1}^{*}}{n^{3/4}}\right).$
Putting this all together,
$\sqrt{n}\max_{i\in I_{1}}|\widehat{\gamma}_{(1),i}-\gamma_{(1),i}^{*}|\lesssim(\widehat{\lambda}_{n,1}+\lambda_{n,1}^{*})\ell^{3}\sqrt{n}+\sqrt{\lambda_{n,1}^{*}}\ell^{3/2}+\lambda_{n,1}^{*}n^{-1/4}.$
Further, we show in Lemma 11 that $\widehat{\lambda}_{n,1}$ and $\lambda_{n,1}^{*}$ are both $O(n^{-3/7})$. Consequently,
$\sqrt{n}\max_{i\in I_{1}}|\widehat{\gamma}_{(1),i}|$ $\leq\sqrt{n}\max_{i\in I_{1}}|\gamma_{(1),i}^{*}|+\sqrt{n}\max_{i\in I_{1}}|\widehat{\gamma}_{(1),i}-\gamma_{(1),i}^{*}|=\mathcal{O}_{p}\left(n^{-3/7}\ell^{3}\sqrt{n}\right)=\mathcal{O}_{p}\left(n^{1/14}\ell^{3}\right),$
which is $o_{p}(1)$ since $\ell=o(n^{-1/12})$. $\square$
Appendix E Auxiliary results
[…]
Following van der Vaart and Wellner (1996, Section 2.7), we consider the class of functions $f:\mathcal{Z}\to\mathbb{R}$ defined on a bounded set $\mathcal{Z}\subset\mathbb{R}$ such that the derivatives of $f$ up to order $k$ are uniformly bounded, and $f^{(k)}$ is Lipschitz continuous. We can define a norm on this class as follows:
$\|f\|_{k+1}:=\max_{j\leq k}\sup_{x}|f^{(j)}(x)|+\sup_{x,y}\frac{|f^{(k)}(x)-f^{(k)}(y)|}{|x-y|},$
where the last supremum is over all $x,y$ in the interior of $\mathcal{Z}$ with $x\neq y$. Finally, denote by $\mathcal{C}_{M}^{k}(\mathcal{Z})$ the set of all continuous functions $f:\mathcal{Z}\to\mathbb{R}$ with $\|f\|_{k}\leq M$. Throughout, we use $A\lesssim_{\eta}B$ to denote that $A\leq CB$ for some arbitrary constant $C>0$ that can depend on $\eta$. For a vector $\lambda=(\lambda_{1},\lambda_{2},\dots,\lambda_{p})$ we denote $\lambda_{-1}:=(\lambda_{2},\dots,\lambda_{p})$. We use $\|x\|_{\infty}$ to denote $\max_{i}|x_{i}|$ when $x$ is a vector, and $\sup_{t}|x(t)|$ when $x$ is a function.
Lemma 6.
Consider the following class of functions:
[…]
where $k\geq 0$ is an integer and $0<\ell\leq 1$. It holds that
$\log N(\varepsilon,\mathcal{F}_{B,\,\ell,\,k},\|\cdot\|_{\infty})\lesssim\ell(B/\varepsilon)^{1/(k+1)},\quad\text{and}\quad\log N_{[\,]}(\varepsilon,\mathcal{F}_{B,\,\ell,\,k},L_{r}(P))\lesssim\ell(B/\varepsilon)^{1/(k+1)},$
for any $r\geq 1$ and probability measure $P$ on $\mathbb{R}$. Thus, $\mathcal{F}_{B,\,\ell,\,k}$ is universally Donsker.
Proof of Lemma 6.
For any $\rho\in\mathcal{F}_{B,\,\ell,\,k}$, the derivatives of $\rho$ upto order $k$ are all bounded by $B\ell$. Thus, $\|\rho\|_{k+1}\leq B(1+\ell)\leq 2B$. Consequently,
$N(\varepsilon,\mathcal{F}_{B,\,\ell,\,k},\|\cdot\|_{\infty})$ $\leq N(\varepsilon,\mathcal{C}^{k+1}_{2B}([-\ell,\ell]),\|\cdot\|_{\infty})$
$\leq N(\varepsilon/2B,\mathcal{C}^{k+1}_{1}([-\ell,\ell]),\|\cdot\|_{\infty}).$
For any $\delta>0$, note that $N(\delta,[-\ell,\ell],|\cdot|)\lesssim\ell/\delta$. We can thus slightly tweak the proof of van der Vaart and Wellner (1996, Theorem 2.7.1), with $\alpha=k+1$ and $\lambda(\mathcal{X}_{1})=2\ell$ (in their notation), to conclude that
$\log N(\varepsilon/2B,\mathcal{C}^{k+1}_{1}([-\ell,\ell]),\|\cdot\|_{\infty}))\lesssim\ell(\varepsilon/2B)^{-1/(k+1)},$
as desired to show. This gives us the first entropy bound (without bracketing).
Next, we can adapt the proof technique of van der Vaart and Wellner (1996, Corollary 2.7.2) to deduce that the last display also yields the desired bound on entropy with bracketing. Finally, the bound on $\log N_{[\,]}(\varepsilon,\mathcal{F}_{B,\,\ell,\,k},L_{2}(P))$ implies that $\mathcal{F}_{B,\,\ell,\,k}$ is $P$-Donsker for any probability distribution $P$ (see, e.g., van der Vaart (1998, Theorem 19.5)). $\square$
For the rest of this section, we assume without loss of generality that the treatment threshold is at $c=0$. Before stating the next few lemmas, we recall some notation. The empirical dual problem is given by
$\widehat{\theta}_{n}(\sigma^{2},B)=\mathop{\rm argmin}_{\theta=({\rho},\lambda)\in\Theta(\ell)}\widehat{L}_{n}(\theta;B),\quad\widehat{L}_{n}(\theta;B):=\frac{1}{4n}\sum_{i=1}^{n}\frac{G^{2}(\theta;X_{i},W_{i})}{\sigma_{i}^{2}}+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n},$
where $G(\theta;x,w)=G(\rho,\lambda;x,w)={\rho}(x)+\lambda_{2}w+\lambda_{3}(1-w)+\lambda_{4}wx+\lambda_{5}(1-w)x+\lambda_{6}x^{2}$, with $w=1(x>0)$, and
$\Theta(\ell):=\left\{(f,\lambda)\mid f:[-\ell,\ell]\to\mathbb{R},f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is $\lambda_{1}$-Lipschitz},\lambda\in[0,\infty)\times\mathbb{R}^{5}\right\}.$
Define the dual problem at the population level as
$\theta_{n}^{*}(\sigma^{2},B)=\mathop{\rm argmin}_{\theta\in\Theta(\ell)}L_{n}(\theta;\sigma^{2},B),\quad\text{where}\quad L_{n}(\theta;\sigma^{2},B):=\frac{1}{4\sigma^{2}}\mathbb{E}\left[G^{2}(\theta;X,W)\right]+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n}.$
In Algorithm 1, we sample splitting to estimate the unknown Lipschitz constant $B$ and the homoscedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$. However, for ease of exposition, we assume throughout this proof that an estimate $\widehat{B}$ of $B$ and an estimate $\widehat{\sigma}^{2}$ of $\sigma^{2}$ is available independently of the data using which we solve the empirical dual problem. We define
$\widehat{\theta}_{n}=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B}),\quad\text{and}\quad\theta_{n}^{*}=\theta_{n}^{*}(\widehat{\sigma}^{2},B).$
Lemma 7.
Define
$\Theta_{n}(\ell,\delta,\varepsilon):=\{(\rho,\lambda)\mid(\rho,\lambda)\in\Theta(\ell),\ |\lambda_{1}-\lambda_{n,1}^{*}|\leq\delta\lambda_{n,1}^{*}\text{ and }\|\lambda_{-1}-\lambda^{*}_{n,-1}\|_{2}\leq\varepsilon\|\lambda_{n,-1}^{*}\|_{2}\},$
where $\ell,\delta,\varepsilon\in(0,1)$ are fixed. Then, the following maximal inequality holds.
$\mathbb{E}\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})|\lesssim(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\|\lambda_{n,-1}^{*}\|_{2}^{2}\frac{\sqrt{\varepsilon}}{\sqrt{n}}+\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\frac{\ell^{3}}{\sqrt{n}}.$
Proof of Lemma 7.
We begin by noting that
$\left|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})\right|=\frac{1}{4\widehat{\sigma}^{2}}|(\mathbb{P}_{n}-\mathbb{P})G^{2}(\theta;X,W)|.$
We prove a maximal inequality for the above RHS using tools from empirical process theory. Write $G(\rho,\lambda;x,w)={\rho}(x)+\lambda_{-1}^{\top}v(x,w)$ where $v(x,w):=(w,1-w,wx,(1-w)x,x^{2})^{\top}$, and note that $\|v(x,w)\|_{2}\leq\sqrt{1+\ell+\ell^{2}}\leq 2$. To reduce notation, we write $\Theta_{n}=\Theta_{n}(\ell,\delta,\varepsilon)$ for the rest of this proof. Using triangle inequality,
$\sup_{\theta\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})G^{2}(\theta;X,W)|$ $\leq\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})\rho^{2}(X)|+\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})(\lambda_{-1}^{\top}v(X,W))^{2}|$
$\qquad+2\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})\rho(X)\lambda_{-1}^{\top}v(X,W)|.$
We will bound each of the above three quantities in order. First, note that for any $(\rho,\lambda)\in\Theta_{n}$, $\rho^{(j)}(0)=0$ for $j=0,1,2$, and $\rho^{\prime\prime}$ is $B_{1}:=(1+\delta)\lambda_{n,1}^{*}$-Lipschitz. This gives $|\rho^{\prime\prime}(x)|=|\rho^{\prime\prime}(x)-\rho^{\prime\prime}(0)|\leq B_{1}\ell$. Moreover, a Taylor expansion of ${\rho}(x)$ about $x=0$ yields the following
$|{\rho}(x)|=|\rho(x)-\rho(0)-\rho^{\prime}(0)x|=\frac{|\rho^{\prime\prime}(\xi)-\rho^{\prime\prime}(0)|}{|\xi-x|}\left|\xi-x\right|\frac{x^{2}}{2}\leq B_{1}\frac{\ell^{3}}{2}.$ (33)
A similar argument gives $|\rho^{\prime}(x)|\leq B_{1}\ell^{2}$. Putting these all together, we can say that the function $r=\rho^{2}$ satisfies $r(0)=0$, $r^{\prime}(0)=0$, $r^{\prime\prime}(x)=2\rho^{\prime\prime}(x)\rho(x)+2\rho^{\prime}(x)^{2}$, and thus
$\|r^{\prime\prime}\|_{\infty}\lesssim B_{1}^{2}\ell^{4}\lesssim(\lambda_{n,1}^{*})^{2}\ell^{4}.$
Consequently, using the notation of Lemma 6, we can say that $\mathcal{A}:=\{\rho^{2}:(\rho,\lambda)\in\Theta_{n}\text{ for some }\lambda\}$ is a subset of the class $\mathcal{F}_{\ell,h,1}$ with Lipschitz constant $B\lesssim(\lambda_{n,1}^{*})^{2}\ell^{4}$. Therefore, Lemma 6 implies that
$\log N_{[\,]}(\varepsilon^{\prime},\mathcal{A},L_{2}(P))\leq\log N_{[\,]}(\varepsilon^{\prime},\mathcal{F}_{\ell,h,1},L_{2}(P))\leq\ell(B/\varepsilon^{\prime})^{1/2}\lesssim\lambda_{n,1}^{*}\ell^{3}(\varepsilon^{\prime})^{-1/2}.$
Finally, note that (33) implies that the constant function $F=(\lambda_{n,1}^{*}\ell^{3})^{2}$ works as an envelope to the class $\mathcal{A}$. We can now apply (van der Vaart, 1998, Corollary 19.35) to conclude that
$\mathbb{E}\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})\rho^{2}(X)|$ $=\mathbb{E}\sup_{r\in\mathcal{A}}|(\mathbb{P}_{n}-\mathbb{P})r|$
[…]
$\lesssim\frac{1}{\sqrt{n}}F^{3/4}F^{1/4}=(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}.$
Next, we deal with
$\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})(\lambda_{-1}^{\top}v(X,W))^{2}|=\sup_{\lambda\in\Lambda_{n}}|(\mathbb{P}_{n}-\mathbb{P})(\lambda_{-1}^{\top}v(X,W))^{2}|,$
where $\Lambda_{n}=\Lambda_{n}(\varepsilon)=\{v\in\mathbb{R}^{5}:\|v-\lambda_{n,-1}^{*}\|\leq\varepsilon\|\lambda_{n,-1}^{*}\|\}\subseteq[0,\infty)\times\mathbb{R}^{5}$. Note, for any $\lambda\in\Lambda_{n}$,
$|\lambda_{-1}^{\top}v(x,w)|\leq\|\lambda_{-1}\|_{2}\|v(x,w)\|_{2}\leq(1+\varepsilon)\|\lambda_{n,-1}^{*}\|_{2}\sqrt{1+\ell+\ell^{2}}\leq 4\|\lambda_{n,-1}^{*}\|_{2}.$
Therefore for any fixed $x,w$, and $\lambda,\widetilde{\lambda}\in\Lambda_{n}$,
$|\langle\lambda,v(x,w))\rangle^{2}-\langle\widetilde{\lambda},v(x,w))\rangle^{2}|$ $\leq 8\|\lambda_{n,-1}^{*}\|_{2}|\langle\lambda,v(x,w))\rangle-\langle\widetilde{\lambda},v(x,w))\rangle|$
$\leq 16\|\lambda_{n,-1}^{*}\|_{2}\|\lambda-\widetilde{\lambda}\|_{2},$
showing that the function $\lambda\mapsto(\lambda^{\top}v(x,w))^{2}$ (where $\lambda\in\Lambda_{n}$) is $c$-Lipschitz, for $c=16\|\lambda_{n,-1}^{*}\|$. Consequently,
$\log N_{[\,]}(\varepsilon^{\prime},\Lambda_{n},L_{2}(P))$ $\leq 5\log\left(1+\frac{2c\mathop{\rm diam}(\Lambda_{n})}{\varepsilon^{\prime}}\right)$
$\leq\frac{10c\mathop{\rm diam}(\Lambda_{n})}{\varepsilon^{\prime}}\lesssim\varepsilon\|\lambda_{n,-1}^{*}\|_{2}^{2}(\varepsilon^{\prime})^{-1}.$
We can now apply (van der Vaart, 1998, Corollary 19.35) with envelop $(4\|\lambda_{n,-1}^{*}\|_{2})^{2}$ to conclude that
[…]
$\lesssim\|\lambda_{n,-1}^{*}\|_{2}^{2}\frac{\sqrt{\varepsilon}}{\sqrt{n}}.$
It only remains to prove a maximal inequality for the following supremum:
$\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})\rho(X)\lambda_{-1}^{\top}v(X,W)|.$
Recall from (33) that $|\rho(x)|\leq B_{1}\ell^{3}/2\leq\lambda_{n,1}^{*}\ell^{3}$ for any $(\rho,\lambda)\in\Theta_{n}$. Also recall from above that $|\langle\lambda_{-1},v(x,w)\rangle|\leq 4\|\lambda_{n,-1}^{*}\|_{2}$. Consequently, for any $(\rho,\lambda)\in\Theta_{n}$, and for every $x,w$, we have
$|\rho(x)\lambda_{-1}^{\top}v(x,w)|\leq 4\lambda_{n,1}^{*}\|\lambda_{n,-1}\|_{2}\ell^{3},$
proving that the constant function $4\lambda_{n,1}^{*}\|\lambda_{n,-1}\|_{2}\ell^{3}$ works as an envelope for this class.
Next, we fix $\varepsilon^{\prime}>0$. Suppose that $\mathcal{R}_{\varepsilon}:=\{\rho_{1},\dots,\rho_{N_{1}}\}$ is a minimal $\varepsilon_{1}$-cover for $\pi_{1}(\Theta_{n})=\{\rho:(\rho,\lambda)\in\Theta_{n}\text{ for some }\lambda\}$ under the sup-norm, and $\mathcal{L}_{\varepsilon}:=\{B_{1},\dots,\ell_{N_{2}}\}$ is a minimal $\varepsilon_{2}$-cover for $\Lambda_{n}=\{v\in\mathbb{R}^{5}:\|v-\lambda_{n,-1}^{*}\|_{2}\leq\varepsilon\|\lambda_{n,-1}^{*}\|_{2}\}$ under the $\ell_{2}$-norm, where $\varepsilon_{1}:={\varepsilon^{\prime}}/{8\|\lambda_{n,-1}^{*}\|_{2}}$ and $\varepsilon_{2}={\varepsilon^{\prime}}/{4\lambda_{n,1}^{*}\ell^{3}}$. It is well-known (see, e.g., van der Vaart (1998)) that,
$\log N_{2}\leq 5\log\left(1+\frac{c\varepsilon\|\lambda_{n,-1}^{*}\|}{\varepsilon^{\prime}/4\lambda_{n,1}^{*}\ell^{3}}\right)\lesssim\varepsilon\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|\ell^{3}(\varepsilon^{\prime})^{-1}.$
On the other hand, Lemma 6 yields
$\log N_{1}\lesssim(\varepsilon^{\prime}/8\|\lambda_{n,-1}\|_{2})^{-1/3}(2\lambda_{n,1}^{*})^{1/3}\ell=(\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2})^{1/3}\ell(\varepsilon^{\prime})^{-1/3}.$
Now, for any $\theta=(\rho,\lambda)\in\Theta_{n}$, pick $\theta^{0}=(\rho^{0},\lambda^{0})\in\mathcal{R}_{\varepsilon}\times\mathcal{L}_{\varepsilon}$ such that $\|\rho-\rho^{0}\|_{\infty}\leq\varepsilon_{1}$ and $\|\lambda-\lambda^{0}\|_{2}\leq\varepsilon_{2}$. Note that,
$|\rho(x)\langle\lambda_{-1},v(x,w)\rangle-\rho^{0}(x)\langle\lambda_{-1}^{0},v(x,w)\rangle|$
$\leq|\rho(x)|\cdot|\langle\lambda_{-1},v(x,w)\rangle-\langle\lambda_{-1}^{0},v(x,w)\rangle|+|\langle\lambda_{-1}^{0},v(x,w)\rangle|\cdot|\rho(x)-\rho^{0}(x)|$
$\leq\lambda_{n,1}^{*}\ell^{3}\cdot\|\lambda_{-1}-\lambda_{-1}^{0}\|_{2}\|v(x,w)\|_{2}+4\|\lambda_{n,-1}^{*}\|_{2}\cdot|\rho(x)-\rho^{0}(x)|$
$\leq\lambda_{n,1}^{*}\ell^{3}\cdot\frac{\varepsilon^{\prime}}{4\lambda_{n,1}^{*}\ell^{3}}\cdot 2+4\|\lambda_{n,-1}^{*}\|_{2}\cdot\frac{\varepsilon^{\prime}}{8\|\lambda_{n,-1}^{*}\|_{2}}\leq\varepsilon^{\prime}.$
Therefore, the class $\mathcal{B}$ of functions $(x,w)\mapsto\rho(x)\langle\lambda_{-1},v(x,w)\rangle$, where $(\rho,\lambda)\in\Theta_{n}$, satisfies
$\log N(\varepsilon^{\prime},\mathcal{B},\|\cdot\|_{\infty})$ $\leq\log N_{1}N_{2}\lesssim(\varepsilon^{\prime})^{-1/3}(\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2})^{1/3}\ell+(\varepsilon^{\prime})^{-1}\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|\ell^{3}\varepsilon.$
This also gives us a bound for the entropy with bracketing (cf. van der Vaart and Wellner (1996, Corollary 2.7.2)), as follows.
$\log N_{[\,]}(\varepsilon^{\prime},G(\Theta_{n}),L_{2}(P))$ $\leq\log N(2\varepsilon^{\prime},G(\Theta_{n}),|\cdot|_{\infty})$
$\lesssim(\varepsilon^{\prime})^{-1/3}(\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\ell^{3})^{1/3}+(\varepsilon^{\prime})^{-1}\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|\ell^{3}\varepsilon.$
Recall from above that the constant function $F_{ab}:=4\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\ell^{3}$ acts as an envelop function for the class $\mathcal{B}$. We now apply (van der Vaart, 1998, Corollary 19.35) once more to conclude that
[…]
$\lesssim\frac{1}{\sqrt{n}}\left((F_{ab})^{5/6}(F_{ab})^{1/6}+(F_{ab})^{1/2}\varepsilon^{1/2}(F_{ab})^{1/2}\right)$
$\lesssim\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\frac{\ell^{3}}{\sqrt{n}},$
which completes the proof. $\square$
Lemma 8.
Let $G(\rho,\lambda;x,w):=\rho(x)+\lambda_{-1}^{\top}v(x,w)$, where $v(x,w):=(w,1-w,xw,x(1-w),x^{2})$. Assume that the assumptions in Theorem 1 hold. Then, there exists $0<\alpha\leq 1$ such that, for any $u\in\mathbb{R}^{5}$,
$u^{\top}\mathbb{E}\left[v(X,W)v(X,W)^{\top}\right]u\geq\alpha\|u\|_{2}^{2}.$
Further, $\theta_{n}^{*}(\sigma^{2},B)=\mathop{\rm argmin}_{\theta\,\in\,\Theta(\ell)}L_{n}(\theta;\sigma^{2},B)$ satisfies the following.
$L_{n}(\theta;\sigma^{2},B)-L_{n}(\theta_{n}^{*};\sigma^{2},B)$ $\geq\frac{1}{4\sigma^{2}}\mathbb{E}\left(\rho(X)-\rho_{n}^{*}(X)+\langle\lambda_{-1}-\lambda_{n,-1}^{*},v(X,W)\rangle\right)^{2}+\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$\geq\frac{\alpha}{4\sigma^{2}}\left(\|\rho-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\lambda_{-1}-\lambda_{n,-1}^{*}\|_{2}^{2}\right)+\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}-R_{n},$
where $R_{n}=\frac{2}{4\sigma^{2}}\mathbb{E}|(\rho(X)-\rho_{n}^{*}(X))\langle\lambda_{-1}-\lambda_{n,-1}^{*},v(X,W)\rangle|$.
Proof of Lemma 8.
Define $A:=\mathbb{E}[v(X,W)v(X,W)^{\top}]$. Since $A$ is positive semidefinite, it has non-negative eigenvalues. We first show that the smallest eigenvalue of $A$, which we denote as $\lambda_{\min}(A)$, is strictly positive. To prove this, assume to the contrary that $\lambda_{\min}(A)=0$. Then, there exists $u\in\mathbb{R}^{5}$, $u\neq 0$, such that $u^{\top}Au=0$. We can conclude from $u^{\top}Au=\mathbb{E}(v(X,W)^{\top}u)^{2}=0$ that $v(X,W)^{\top}u=0$ with probability $1$. This implies that, either $X$ is a solution of $u_{1}+u_{3}x+u_{5}x^{2}=0$ (when $w=\text{$\mathbf{1}$}{\left\{x>0\right\}}=1$), or a solution of $u_{2}+u_{4}x+u_{5}x^{2}=0$ (when $w=\text{$\mathbf{1}$}{\left\{x>0\right\}}=0$). But, $X$ having a finite support contradicts the assumption that the running variable has a strictly positive density at the cutoff. So we must have $\lambda_{\min}(A)>0$. Therefore,
$u^{\top}\mathbb{E}\left[v(X,W)v(X,W)^{\top}\right]u\geq\alpha\|u\|_{2}^{2},$
for any $u\in\mathbb{R}^{5}$, where $\alpha:=\min\{1,\lambda_{\min}(A)\}$.
For proving the second part, we write $L_{n}=L_{n}(\,\cdot\,;\sigma^{2},B)$ for notational convenience. First, we simplify $L_{n}(\theta)-L_{n}(\theta_{n}^{*})$ using the fact that gradients of the loss $L_{n}$ at $\theta_{n}^{*}$ are zero. To remove the dependence of the domain on the parameter $\lambda_{1}$, we write
$L_{n}(\rho,\lambda)=\frac{1}{4\sigma^{2}}\mathbb{E}\left(\lambda_{1}\widetilde{\rho}(X)+\lambda_{-1}^{\top}v(X,W)\right)^{2}+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n}=:\widetilde{L}_{n}(\widetilde{\rho},\lambda),$
where $\widetilde{\rho}(\cdot)=\rho(\cdot)/\lambda_{1}\in\mathcal{R}$. Since $\widetilde{L}_{n}$ is minimized at $(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})$, we can deduce the following.
  1. 1.
    $\nabla_{\lambda_{-1}}\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})=0$. This implies that,
    $\frac{1}{2\sigma^{2}}\mathbb{E}\left[\left(\lambda_{n,1}^{*}\widetilde{\rho}_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)v(X,W)\right]+\frac{e_{1}-e_{2}}{n}=0.$ (34)
    where $e_{1}=(1,0,0,0,0)$ and $e_{2}=(0,1,0,0,0)$.
  2. 2.
    $\nabla_{\lambda_{1}}\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})=0$. This implies that,
    $\frac{1}{2\sigma^{2}}\mathbb{E}\left[\left(\lambda_{n,1}^{*}\widetilde{\rho}_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\widetilde{\rho}_{n}^{*}(X)\right]+\frac{2\lambda_{n,1}^{*}}{4nB^{2}}=0.$ (35)
  3. 3.
    $\nabla_{\rho}\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})=0$. We can say that the convex function $t\mapsto\widetilde{L}_{n}(t\widetilde{\rho}+(1-t)\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})$ is minimized at $t=0$, which gives us the following.
    $\frac{2\lambda_{n,1}^{*}}{4\sigma^{2}}\mathbb{E}\left[\left(\widetilde{\rho}(X)-\widetilde{\rho}_{n}^{*}(X)\right)\left(\lambda_{n,1}^{*}\widetilde{\rho}_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\right]\geq 0.$ (36)
Equipped with Eqs. 34, 35 and 36, we are ready to simplify $L_{n}(\theta)-L_{n}(\theta_{n}^{*})=\widetilde{L}_{n}(\widetilde{\rho},\lambda)-\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})$, as follows.
$L_{n}(\theta)-L_{n}(\theta_{n}^{*})-\frac{1}{4\sigma^{2}}\mathbb{E}\left(\rho(X)-\rho_{n}^{*}(X)+\langle\lambda_{-1}-\lambda_{n,-1}^{*},v(X,W)\rangle\right)^{2}$
$=\frac{2}{4\sigma^{2}}\mathbb{E}\left[(\rho-\rho_{n}^{*})(X)\left(\rho_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\right]+\frac{\lambda_{1}^{2}-(\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$\qquad+\frac{2}{4\sigma^{2}}\mathbb{E}\left[\langle\lambda_{-1}-\lambda_{n,-1}^{*},v(X,W)\rangle\left(\rho_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\right]+\frac{\lambda_{2}-\lambda_{3}-\lambda_{n,2}^{*}+\lambda_{n,3}^{*}}{n}.$
Using (34), we deduce that the second line in the above display is zero. Next, set $\widetilde{\rho}=\rho/\lambda_{1}$, $\widetilde{\rho}_{n}^{*}=\rho_{n}^{*}/\lambda_{n,1}^{*}$. Continuing from above,
$=\frac{2}{4\sigma^{2}}\mathbb{E}\left[(\lambda_{1}\widetilde{\rho}(X)-\lambda_{n,1}^{*}\widetilde{\rho}_{n}^{*}(X))\left(\rho_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\right]+\frac{\lambda_{1}^{2}-(\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$=\frac{2\lambda_{1}}{4\sigma^{2}}\mathbb{E}\left[(\widetilde{\rho}(X)-\widetilde{\rho}_{n}^{*}(X))\left(\rho_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\right]$
$\qquad\qquad+\frac{2}{4\sigma^{2}}\mathbb{E}\left[(\lambda_{1}-\lambda_{n,1}^{*})\widetilde{\rho}_{n}^{*}(X)\left(\rho_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\right]+\frac{\lambda_{1}^{2}-(\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$\geq 0-\frac{2\lambda_{n,1}^{*}}{4nB^{2}}(\lambda_{1}-\lambda_{n,1}^{*})+\frac{\lambda_{1}^{2}-(\lambda_{n,1}^{*})^{2}}{4nB^{2}}$ (using (36) and (35))
$\geq\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}.$
Therefore,
$L_{n}(\theta)$ $-L_{n}(\theta_{n}^{*})\geq\frac{1}{4\sigma^{2}}\mathbb{E}\left(\rho(X)-\rho_{n}^{*}(X)+\langle\lambda_{-1}-\lambda_{n,-1}^{*},v(X,W)\rangle\right)^{2}+\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$\geq\frac{1}{4\sigma^{2}}\left[\mathbb{E}(\rho(X)-\rho_{n}^{*}(X))^{2}+(\lambda_{-1}-\lambda_{n,-1}^{*})^{\top}A(\lambda_{-1}-\lambda_{n,-1}^{*})\right]-R_{n}+\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}$
$\geq\frac{\alpha}{4\sigma^{2}}\left(\|\rho-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\lambda_{-1}-\lambda_{n,-1}^{*}\|_{2}^{2}\right)+\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}-R_{n},$
as desired to show.
$\square$
Lemma 9.
The optimizer $\theta_{n}^{*}=(\rho_{n}^{*},\lambda_{n}^{*})$ of the dual population problem (15) satisfies
$\|\lambda_{n,-1}^{*}\|_{2}=O(\lambda_{n,1}^{*}\ell^{3}+n^{-1}).$
Consequently,
$\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},B)|\lesssim\frac{(\lambda_{n,1}^{*})^{2}}{n}\left|\frac{1}{\widehat{B}}-\frac{1}{B}\right|.$
Proof of Lemma 9.
Recall from (34) in the proof of Lemma 8 that
$\frac{1}{2\sigma^{2}}\mathbb{E}\left[\left(\lambda_{n,1}^{*}\widetilde{\rho}_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)v(X,W)\right]+\frac{e_{1}-e_{2}}{n}=0.$
Rearranging this equation yields
$A\lambda_{n,-1}^{*}=-\lambda_{n,1}^{*}\mathbb{E}\widetilde{\rho}_{n}^{*}(X)v(X,W)-2\sigma^{2}\frac{(e_{1}-e_{2})}{n}$
where $A:=\mathbb{E}[v(X,W)v(X,W)^{\top}]$, as defined in the proof of Lemma 8. Since $\|v(X,W)\|_{2}\leq 2$ and (33) yields $\|\widetilde{\rho}_{n}^{*}\|_{\infty}\leq\ell^{3}/2$, we deduce that the above right hand side is $O(\lambda_{n,1}^{*}\ell^{3}+1/n)$, which finishes the proof of the first part. The second conclusion follows immediately from the definition once we use $\|\lambda_{n,-1}^{*}\|_{2}=O(\lambda_{n,1}^{*}\ell^{3}+1/n)$. $\square$
Lemma 10.
Consider the neighborhood $\Theta_{n}(\ell,\delta,\varepsilon)$ of the optimizer $\theta_{n}^{*}=\theta_{n}^{*}(\widehat{\sigma}^{2},B)$ of the population dual problem, as defined in Lemma 7. Denote by $\widehat{\theta}_{n}=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})$ the optimizer of the empirical dual problem. Assume that $\widehat{B}\rightarrow_{p}B$, and that $\ell=o(n^{-1/12})$. For any $\delta,\eta\in(0,1)$, and $\varepsilon=O(n^{-1/4})$, it holds that
$\mathbb{P}(\widehat{\theta}_{n}\in\Theta_{n}(\ell,\delta,\varepsilon))\geq 1-\eta,$
for all sufficiently large $n$.
Proof of Lemma 10.
It follows from Lemma 7 that for any $n\geq 1$,
$\mathbb{E}\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})|\lesssim(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\|\lambda_{n,-1}^{*}\|_{2}^{2}\frac{\sqrt{\varepsilon}}{\sqrt{n}}+\lambda_{n,1}^{*}\|\lambda_{n,-1}^{*}\|_{2}\frac{\ell^{3}}{\sqrt{n}},$
where $\Theta_{n}(\ell,\delta,\varepsilon):=\{(\rho,\lambda)\in\Theta(\ell)\text{ such that }|\lambda_{1}-\lambda_{n,1}^{*}|\leq\delta\lambda_{n,1}^{*}\text{ and }\|\lambda_{-1}-\lambda^{*}_{n,-1}\|_{2}\leq\varepsilon\|\lambda_{n,-1}^{*}\|_{2}\}$. Invoking Lemma 9, we reduce the above to the following.
$\mathbb{E}\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})|\lesssim(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}.$
Fix $\eta\in(0,1)$. It follows from above that for some absolute constant $C=C(\eta)>0$, we have
$\mathbb{P}\left(\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})|\leq C\left((\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}\right)\right)\geq 1-\frac{\eta}{2},$ (37)
On the other hand, Lemma 9 also tells us that
$\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},B)|\leq C\frac{(\lambda_{n,1}^{*})^{2}}{n}\left|\frac{1}{\widehat{B}}-\frac{1}{B}\right|.$
Consequently,
$\mathbb{P}\left(\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|L_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},B)|\leq\delta^{2}\frac{(\lambda_{n,1}^{*})^{2}}{16nB^{2}}\right)\geq 1-\frac{\eta}{2},$
for all sufficiently large $n$. This, combined with (37) implies that,
$\mathbb{P}\left(\sup_{\theta\,\in\,\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},B)|\leq C\left(\frac{(\lambda_{n,1}^{*})^{2}\ell^{6}}{\sqrt{n}}+\frac{\lambda_{n,1}^{*}\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}\right)+\frac{\delta^{2}(\lambda_{n,1}^{*})^{2}}{16nB^{2}}\right)\geq 1-\eta,$ (38)
for all sufficiently large $n$. Next, we recall from Lemma 8 that
$L_{n}(\theta;\widehat{\sigma}^{2},B)-L_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},B)\geq\frac{\alpha}{4\widehat{\sigma}^{2}}\left(\|\rho-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\lambda_{-1}-\lambda_{n,-1}^{*}\|_{2}^{2}\right)+\frac{(\lambda_{1}-\lambda_{n,1}^{*})^{2}}{4nB^{2}}-R_{n},$ (39)
for some $\alpha\in(0,1)$, where
$R_{n}=\frac{2}{4\widehat{\sigma}^{2}}\mathbb{E}|(\rho(X)-\rho_{n}^{*}(X))\langle\lambda_{-1}-\lambda_{n,-1}^{*},v(X,W)\rangle|.$
We define another neighborhood of $\theta_{n}^{*}=(\rho_{n}^{*},\lambda_{n}^{*})$, as follows:
$\widetilde{\Theta}_{n}:=\{(\rho,\lambda)\in\Theta(\ell):\|\rho-\rho_{n}^{*}\|_{L^{2}(P)}^{2}+\|\lambda_{-1}-\lambda_{n,-1}^{*}\|_{2}^{2}\leq\varepsilon^{2}\|\lambda_{n,-1}^{*}\|_{2}^{2}\text{ and }|\lambda_{1}-\lambda_{n,1}^{*}|\leq\delta\lambda_{n,1}^{*}\}.$
For $(\rho,\lambda)\in\widetilde{\Theta}_{n}$, it follows from the Cauchy-Schwarz inequality that
$R_{n}\leq\frac{1}{4\widehat{\sigma}^{2}}\left(\|\rho-\rho_{n}^{*}\|_{L^{2}(P)}^{2}+\|\langle\lambda_{n,-1}-\lambda_{n,-1}^{*},v(X,W)\rangle\|_{2}^{2}\right)\leq\frac{\beta}{4\widehat{\sigma}^{2}}\varepsilon^{2}\|\lambda_{n,-1}^{*}\|_{2}^{2},$ (40)
where $\beta=\max\{1,\lambda_{\max}\}$, where $\lambda_{\max}$ is the largest eigenvalue of $A=\mathbb{E}[v(X,W)v(X,W)^{\top}]$. It is immediate from definition that $\widetilde{\Theta}_{n}\subseteq\Theta_{n}(\ell,\delta,\varepsilon)$, and that for every $\theta$ on the boundary $\partial\widetilde{\Theta}_{n}$,
$\|\rho-\rho_{n}^{*}\|_{L^{2}(P)}^{2}+\|\lambda_{-1}-\lambda_{n,-1}^{*}\|_{2}^{2}=\varepsilon^{2}\|\lambda_{n,-1}^{*}\|_{2}^{2},\quad\text{and}\quad|\lambda_{1}-\lambda_{n,1}^{*}|=\delta\lambda_{n,1}^{*}.$
We can now combine (38), (39) and (40) to conclude that, for any $\theta\in\partial\widetilde{\Theta}_{n}$,
$\frac{\alpha}{4\widehat{\sigma}^{2}}\varepsilon^{2}\|\lambda_{n,-1}^{*}\|_{2}^{2}+\frac{\delta^{2}(\lambda_{n,1}^{*})^{2}}{4nB^{2}}-\frac{\beta}{4\widehat{\sigma}^{2}}\varepsilon^{2}\|\lambda_{n,-1}^{*}\|_{2}^{2}$
$\leq L_{n}(\theta;\widehat{\sigma}^{2},B)-L_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},B)$ (using (39) and (40))
$\leq\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-\widehat{L}_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},\widehat{B})+2\sup_{\theta\in\Theta_{n}(\ell,\delta,\varepsilon)}|\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-L_{n}(\theta;\widehat{\sigma}^{2},B)|$ (41)
$\leq\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})-\widehat{L}_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},\widehat{B})+C\left((\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}\right)+\delta^{2}\frac{(\lambda_{n,1}^{*})^{2}}{8nB^{2}},$ (using (38))
with probability at least $1-\eta$, for any $n\geq 1$. We can rewrite the above inequality as $\widehat{L}_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},\widehat{B})\leq\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})+\Delta_{n}$ and we show below that $\Delta_{n}<0$ for all sufficiently large $n$. This implies, by taking infimum over $\theta\in\partial\widetilde{\Theta}_{n}$, that
$\mathbb{P}\left(\widehat{L}_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},\widehat{B})\leq\inf_{\theta\in\partial\widetilde{\Theta}_{n}}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})\right)\geq 1-\eta,$
for all sufficiently large $n$. Finally, note that since the empirical loss function $\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})$ is convex in $\theta$, the above event implies that $\widehat{\theta}_{n}=\mathop{\rm argmin}_{\theta}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})$ must be inside $\widetilde{\Theta}_{n}$. Thus,
$\mathbb{P}\left(\widehat{\theta}_{n}\in\widetilde{\Theta}_{n}\right)\geq 1-\eta,$
for all sufficiently large $n$, as desired to show.
To complete the proof, we need to show that $\Delta_{n}<0$ for all sufficiently large $n$. Observe that,
$\Delta_{n}$ $=C\left((\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}\right)+\delta^{2}\frac{(\lambda_{n,1}^{*})^{2}}{8nB^{2}}-\left(\frac{\delta^{2}(\lambda_{n,1}^{*})^{2}}{4nB^{2}}-\frac{\beta-\alpha}{4\widehat{\sigma}^{2}}\varepsilon^{2}\|\lambda_{n,-1}^{*}\|_{2}^{2}\right)$
$\leq C\left((\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}+\lambda_{n,1}^{*}\frac{\ell^{3}}{n^{3/2}}+\frac{\sqrt{\varepsilon}}{n^{5/2}}\right)-\delta^{2}\frac{(\lambda_{n,1}^{*})^{2}}{8nB^{2}}+c_{0}\varepsilon^{2}\left(\lambda_{n,1}^{*}\ell^{3}+\frac{1}{n}\right)^{2}$
$=\frac{(\lambda_{n,1}^{*})^{2}}{n}\left(C\ell^{6}\sqrt{n}+c_{0}\varepsilon^{2}\ell^{6}n-\frac{\delta^{2}}{8B^{2}}+\frac{C\ell^{3}}{\lambda_{n,1}^{*}\sqrt{n}}+2c_{0}\varepsilon^{2}\frac{\ell^{3}}{\lambda_{n,1}^{*}}+\frac{C\sqrt{\varepsilon}}{n^{3/2}(\lambda_{n,1}^{*})^{2}}+\frac{c_{0}\varepsilon^{2}}{n(\lambda_{n,1}^{*})^{2}}\right).$
We show in Lemma 11 that $\lambda_{n,1}^{*}=O(n^{-3/7})$. In light of this, we can write
$\limsup_{n\to\infty}\frac{\Delta_{n}}{(\lambda_{n,1}^{*})^{2}/n}\leq\lim_{n\to\infty}\left(C\ell^{6}\sqrt{n}+c_{0}\varepsilon^{2}\ell^{6}n-\frac{\delta^{2}}{8B^{2}}\right)\leq-\frac{\delta^{2}}{8B^{2}}<0,$
since $\ell^{6}\sqrt{n}=o(1)$ and $\varepsilon^{2}\ell^{6}n=\varepsilon^{2}\sqrt{n}\cdot\ell^{6}\sqrt{n}=o(1)$ as well. This completes the proof. $\square$
Lemma 11.
Under the conditions of Theorem 1,
$\widehat{\lambda}_{n,1}(\widehat{\sigma}^{2},\widehat{B})=\mathcal{O}_{p}(n^{-3/7}),\quad\lambda_{n,1}^{*}(\widehat{\sigma}^{2},{B})=\mathcal{O}_{p}(n^{-3/7}),\quad\text{and}\quad\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}=\mathcal{O}_{p}(n^{-6/7})$
Proof of Lemma 11.
[…]
$\left(\widehat{\tau}(h_{n}),\widehat{\alpha},\widehat{\beta}\right)=\mathop{\rm argmin}_{\tau,\alpha,\beta_{1},\beta_{0}\in\mathbb{R}}\ \sum_{i=1}^{n}K\left(\frac{X_{i}}{h_{n}}\right)\left(Y_{i}-\alpha-\tau W_{i}-\beta_{0}X_{i}^{-}-\beta_{1}X_{i}^{+}-\beta_{2}X_{i}^{2}\right)^{2}.$
For simplicity, we use the triangular kernel $K(x)=1\{|x|\leq 1\}$. It follows along the lines of the proof of Wager (2024, Proposition 8.1) that $\widehat{\tau}(h_{n})$ has worst-case squared curvature bias bounded above by $\widehat{B}h_{n}^{6}$, and sampling noise of the order of $1/nh_{n}$. With $h_{n}=O(n^{-1/7})$, this estimator has MSE of the order of $n^{-6/7}$. Consequently, $\widehat{\lambda}_{n,1}=\mathcal{O}_{p}(n^{-3/7})$ and $\|\widehat{\gamma}_{n}\|_{2}^{2}=\mathcal{O}_{p}(n^{-6/7})$. The proof for $\lambda_{n,1}^{*}$ is analogous. $\square$
Lemma 12.
Under the conditions of Theorem 3, it holds that $\widehat{b}/{s}(\widehat{\gamma})=\mathcal{O}_{p}(1)$ as $n\to\infty$, where ${s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}$, with $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$.
Proof of Lemma 12.
Assume $c=0$ without loss of generality. First, we can replace ${s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}$ with $\widehat{V}_{n}:=\widehat{\sigma}^{2}\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}$, using the fact that $0<\sigma_{\min}\leq\sigma_{i}\leq\sigma_{\max}<\infty$. Recall that, ignoring cross-fitting notation as mentioned in the beginning of this section, we define
$(\widehat{\gamma}_{n},\widehat{t}_{n}):=\mathop{\rm argmin}_{\gamma,\,t}\left(\widehat{B}^{2}t^{2}+\widehat{\sigma}^{2}\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\right)\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\sum_{i=1}^{n}\gamma_{i}\rho(X_{i})\leq t,$
[…]
$(\gamma^{*}_{n},t_{n}^{*}):=\mathop{\rm argmin}_{\gamma,t}\left\{\widehat{B}^{2}t^{2}+\frac{1}{n}\widehat{\sigma}^{2}\mathbb{E}[\gamma^{2}(X)]\right\}\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\mathbb{E}[\gamma(X)\rho(X)]\leq t.$
[…]
To prove this by contradiction, assume that for any $c>0$, we have $V_{n}^{*}\leq_{p}cn^{-6/7}$ infinitely often. We examine the MSE of $\gamma_{n}^{*}$ for the objective
$R_{m}(\gamma):=B^{2}\left(\sup_{\rho\in\mathcal{R}}\mathbb{E}\gamma(X)\rho(X)\right)^{2}+\frac{1}{m}\mathbb{E}[\gamma^{2}(X)],$
where $m=n/k$ (assume without loss of generality that $n$ is divisible by $k$). Note that
$R_{m}(\gamma_{n}^{*})$ $=\widehat{B}^{2}\left(\sup_{\rho\in\mathcal{R}}\mathbb{E}\gamma_{n}^{*}(X)\rho(X)\right)^{2}+\frac{1}{m}\mathbb{E}[\gamma^{*,2}_{n}(X)]$
$=\widehat{B}^{2}(t_{n}^{*})^{2}+kV_{n}^{*}\leq_{p}(a_{+}+kc)k^{-6/7}m^{-6/7}.$
[…]
Lemma 13.
Suppose that $Z\sim N(0,1)$. Then, for any $s$, $h$, $t$ and $\delta>0$, we have
$\inf_{|b|\,\leq\,t+\delta s}\mathbb{P}(|b+sZ|\leq h)\geq\inf_{|b|\,\leq\,t}\mathbb{P}(|b+sZ|\leq h)-\delta.$
Proof of Lemma 13.
Fix any $s$, $h$, $t$ and $\delta>0$. Denote
$g(b)=\mathbb{P}(|b+sZ|\leq h)=\Phi\left(\frac{h-b}{s}\right)-\Phi\left(\frac{-h-b}{s}\right).$
Note that,
$g^{\prime}(b)=-\frac{1}{s}\left(\phi\left(\frac{h-b}{s}\right)-\phi\left(\frac{-h-b}{s}\right)\right)\implies\|g^{\prime}\|_{\infty}\leq\frac{2\phi(0)}{s}\leq\frac{1}{s}.$
Next, fix any $b_{1}$ with $|b_{1}|\leq t+\delta s$. There exists $b_{0}$ such that $|b_{0}|\leq t$ and $|b_{1}-b_{0}|\leq\delta s$. Consequently, $|g(b_{1})-g(b_{0})|\leq\|g^{\prime}\|_{\infty}|b_{1}-b_{0}|\leq\delta$. This implies that $g(b_{1})\geq g(b_{0})-\delta\geq\inf_{|b|\leq t}g(b)-\delta$. Taking infinimum over $b_{1}$, we get the desired conclusion. $\square$
Lemma 14.
Under the conditions of Theorem 3, it holds that $\widehat{s}(\widehat{\gamma})/{s}(\widehat{\gamma})\rightarrow_{p}1$ as $n\to\infty$, where ${s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}$ with $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$, and $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\widehat{\sigma}_{i}^{2}$ with $\widehat{\sigma}_{i}^{2}=(Y_{i}-\widehat{\mu}_{W_{i}}(X_{i}))^{2}$ as defined in Algorithm 1.
Proof of Lemma 14.
In Algorithm 1, we use sample-splitting and fit regression on either side of the threshold; here we assume for simplicity in notation that we fit the regression of $Y_{i}$ on $X_{i}$ globally (our argument applied on either side of the threshold will then complete the proof). Also assume without loss of generality that $c=0$. Write $\widetilde{X}_{i}=(1,X_{i})$, $\widehat{\varepsilon}_{i}=Y_{i}-\widehat{\mu}(X_{i})=Y_{i}-\widetilde{X}_{i}\widehat{\beta}$ and $\varepsilon_{i}^{*}=Y_{i}-\mu(X_{i})=Y_{i}-\widetilde{X}_{i}\beta^{*}$ where $\widetilde{X}\beta^{*}$ is the best linear projection of $\mu(X)=\mathbb{E}[Y\mid X]$ onto $\text{span}\{1,X\}$. It follows from White (1980) that $\widehat{\beta}\rightarrow_{p}\beta^{*}$. Next, write
$\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\widehat{\varepsilon}_{i}^{2}=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\varepsilon_{i}^{*2}+2\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\varepsilon_{i}^{*}\widetilde{X}_{i}^{\top}(\beta^{*}-\widehat{\beta})+\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}(\widetilde{X}_{i}^{\top}(\beta^{*}-\widehat{\beta}))^{2}.$ (42)
On the other hand, note that $\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}\geq\sigma^{2}_{\min}\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}$. Therefore,
$\frac{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}(\widetilde{X}_{i}^{\top}(\beta^{*}-\widehat{\beta}))^{2}}{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}}\lesssim\left\|(\widehat{\beta}-\beta^{*})^{\top}\frac{|\widehat{\gamma}_{i}|\widetilde{X}_{i}}{\|\widehat{\gamma}_{i}\|_{2}}\right\|_{2}^{2}=o_{p}(1),$
since $\widehat{\gamma}_{i}=0$ for $|X_{i}|>\ell$. For the cross-term, we apply Cauchy-Schwarz inequality to write
$\frac{\left|\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\varepsilon_{i}^{*}\widetilde{X}_{i}^{\top}(\beta^{*}-\widehat{\beta})\right|}{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}}\leq\left(\frac{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\varepsilon_{i}^{*2}}{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}}\right)^{1/2}\left\|(\widehat{\beta}-\beta^{*})^{\top}\frac{|\widehat{\gamma}_{i}|\widetilde{X}_{i}}{\|\widehat{\gamma}_{i}\|_{2}}\right\|_{2}.$
We can handle the second term exactly as above. For the first part, we show below that
$\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\varepsilon_{i}^{*2}/\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}\rightarrow_{p}1.$ (43)
As a consequence, we can continue from (42) to write
$\frac{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\widehat{\varepsilon}_{i}^{2}}{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}}=\frac{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\varepsilon_{i}^{*2}}{\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2}}+o_{p}(1)+o_{p}(1),$
which completes the proof. To prove (43), fix any $\varepsilon>0$ and $\eta>0$. Define $a_{n,i}=\widehat{\gamma}_{i}^{2}/\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}$, $1\leq i\leq n$, and write $\Delta_{i}=\varepsilon_{i}^{*2}-\sigma_{i}^{2}$. It follows from the assumption $\sup_{x}\mathbb{E}[|Y_{i}-\mu(X_{i})|^{q}\mid X_{i}=x]<\infty$ and Jensen’s inequality that $\mathbb{E}[|\Delta_{i}|^{q/2}|X_{i}=x]\leq C$ for some universal constant $C\in(0,\infty)$. Then,
$\mathbb{E}\left[\left|\sum_{i=1}^{n}a_{n,i}\Delta_{i}\text{$\mathbf{1}$}{\left\{|\Delta_{i}|>M\right\}}\right|\mid X_{1},\dots,X_{n}\right]\leq\sum_{i=1}^{n}a_{n,i}\mathbb{E}\left[\left|\Delta_{i}\right|\text{$\mathbf{1}$}{\left\{|\Delta_{i}|>M\right\}}\mid X_{1},\dots,X_{n}\right]$
$\leq\sum_{i=1}^{n}a_{n,i}\left(\mathbb{E}\left[\left|\Delta_{i}\right|^{q/2}\mid X_{i}\right]\right)^{2/q}\mathbb{P}\left(|\Delta_{i}|>M\mid X_{i}\right)^{1-2/q}$ (Hölder’s inequality)
$\leq\sum_{i=1}^{n}a_{n,i}\frac{1}{M^{q/2-1}}\mathbb{E}\left[\left|\Delta_{i}\right|^{q/2}\mid X_{i}\right]$ (Markov’s inequality)
$\leq\frac{C}{M^{q/2-1}}.$
Since $q>2$, we can pick $M$ large enough so that
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{n,i}\Delta_{i}\text{$\mathbf{1}$}{\left\{|\Delta_{i}|>M\right\}}\right|>\varepsilon/2\right)<\eta/2,$
and $|\mathbb{E}[\Delta_{i}\text{$\mathbf{1}$}{\left\{|\Delta_{i}|\leq M\right\}}|\leq\varepsilon/4$ (the latter is possible since $\lim_{M\to\infty}\mathbb{E}[\Delta_{i}\text{$\mathbf{1}$}{\left\{|\Delta_{i}|\leq M\right\}}]=\mathbb{E}[\Delta_{i}]=0$ by Dominated Convergence Theorem). Consequently, by Chebyshev’s inequality,
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{n,i}\Delta_{i}\text{$\mathbf{1}$}{\left\{|\Delta_{i}|\leq M\right\}}\right|>\varepsilon/2\,\bigg{|}\,\{X_{i}\}\right)\lesssim\sum_{i=1}^{n}a_{n,i}^{2}M^{2}\leq M^{2}\max_{1\leq i\leq n}a_{n,i}<\eta/2,$
for all sufficiently large $n$, since we have $\max_{i\leq n}a_{n,i}=o_{p}(1)$ from Corollary 5. Putting these together, we obtain $\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}(\varepsilon_{i}^{*2}-\sigma_{i}^{2})/\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}=o_{p}(1)$. Since $0\leq\sigma^{2}_{\min}\leq\sigma_{i}^{2}$, this finishes the proof of (43). $\square$

### v2

@@
 The fractionWe are grateful to Evan Munro for sharing results on the behavior of existing methods for RDDs under generative adversarial simulation following Athey et al.
-(2024), and to P.
-Aronow for drawing our attention to their behavior on pure noise.
-This work was supported by the National Science Foundation under grant SES-2242876, the Office of Naval Research under grants N00014-17-1-2131 and N00014-19-1-2468, and gift from Amazon.
-of empirical NBER working papers using regression discontinuity designs (RDDs) increased from nearly zero in 1995 to more than 10% by 2015 (Currie, Kleven, and Zwiers, 2020, Figure 4B).
+(2024).
+We are also grateful to Tim Armstrong, P.
+Aronow, Max Farrell, Paul Goldsmith-Pinkham, Michal Kolesár, Christoph Rothe and Rocío Titiunik for insightful conversations and for comments on a previous version of this paper.
+This work was supported by the National Science Foundation under grant SES-2242876, the Office of Naval Research under grants N00014-17-1-2131 and N00014-19-1-2468, and a gift from Amazon.
+Software and replication files are available at https://github.com/ghoshadi/plrd.
+of empirical NBER working papers using regression discontinuity designs (RDDs) increased from nearly zero in 1995 to between 15% and 20% by 2015 (Currie et al., 2020, Figure 4B).
+[^1] Goldsmith-Pinkham (2024) maintains up-to-date numbers at https://paulgp.com/econlit-pipeline.
 Despite the ubiquity of RDDs in applied work, however, there are still open questions as to econometric best practices for inference in RDDs.
-In particular, although standard interpretation of inferential statements relies on confidence intervals achieving nominal coverage of true treatment effects, we find that widely used methods systematically fail to achieve nominal coverage.
-rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
+In particular, we find that widely used methods systematically fail to achieve nominal coverage.
+conventional rdrobust RDHonest plrd
 Dataset EC width EC width EC width EC width
-Lee (2008) 91.0% 0.036 95.4% 0.042 97.6% 0.052 98.3% 0.035
-Cattaneo et al. (2015) 90.7% 5.353 95.1% 6.281 97.4% 6.980 97.3% 4.525
-Matsudaira (2008) (read.) 93.8% 0.061 96.9% 0.069 98.4% 0.096 97.1% 0.060
-Matsudaira (2008) (math) 88.6% 0.045 93.0% 0.054 97.8% 0.072 97.0% 0.051
-Jacob and Lefgren (2004) (read.) 88.4% 0.067 95.0% 0.080 98.2% 0.104 96.5% 0.065
-Jacob and Lefgren (2004) (math) 87.8% 0.078 93.4% 0.093 97.1% 0.120 95.0% 0.069
-Oreopoulos (2006) 84.1% 0.196 94.6% 0.264 98.9% 0.316 96.1% 0.194
-Ludwig and Miller (2007) 91.2% 4.058 95.7% 4.786 97.6% 5.238 97.4% 3.542
-Meyersson (2014) 89.5% 6.362 93.3% 7.531 96.3% 8.099 97.3% 5.095
-Table 1: Empirical coverage (EC) and width of 95% confidence intervals from different methods, on simulated RDDs calibrated to nine applications.
+Lee (2008) 94.4% 0.037 94.8% 0.044 97.3% 0.056 98.1% 0.045
+Cattaneo et al. (2015) 94.6% 5.003 95.0% 5.884 97.5% 6.796 96.1% 3.870
+Ludwig and Miller (2007) 94.4% 5.954 94.9% 7.053 97.8% 7.808 96.8% 4.272
+Meyersson (2014) 93.8% 6.301 94.5% 7.431 96.8% 7.869 96.3% 4.618
+Matsudaira (2008) (read.) 86.4% 0.050 92.2% 0.059 97.9% 0.103 97.7% 0.059
+Matsudaira (2008) (math) 92.0% 0.046 93.4% 0.054 98.3% 0.094 97.6% 0.052
+Jacob and Lefgren (2004) (read.) 94.2% 0.068 94.3% 0.082 97.9% 0.114 97.5% 0.076
+Jacob and Lefgren (2004) (math) 94.4% 0.059 95.3% 0.071 98.2% 0.124 97.9% 0.061
+Oreopoulos (2006) 93.5% 0.189 94.6% 0.255 98.8% 0.303 96.2% 0.124
+Table 1: Empirical coverage (EC) and (average) width of 95% confidence intervals from different methods, on simulated RDDs calibrated to nine applications.
 Following Athey et al.
 (2024) we generate our simulation distributions by training a Wasserstein Generative Adversarial Neural Network (WGAN) on the original data.
-We report results for rdrobust (Calonico et al., 2014) (both the “b.c.” method justified by asymptotics and the “rob” method with a finite-sample variance correction) and RDHonest (Armstrong and Kolesár, 2018).
+We report results for the conventional local linear regression with bandwidth chosen as recommended in Imbens and Kalyanaraman (2012); robust bias-corrected (rdrobust) confidence intervals of Calonico et al.
+(2014); and bias-aware (RDHonest) confidence intervals as proposed in Armstrong and Kolesár (2018).
 We also report results for our proposed plrd method.
 We highlight (in green) the entries whose coverage is not statistically significantly (i.e., at the 5% level using a one-sided binomial test) below nominal 95% coverage and (in blue) the entries with the shortest widths among those delivering valid coverage.
-All results are aggregated across 1000 simulation replications.
-We assess the performance of existing methods by evaluating them on simulations calibrated to nine high-profile applications of RDDs.
-We focus on approaches proposed by Calonico, Cattaneo, and Titiunik (2014) and Armstrong and Kolesár (2018), as implemented in the R functions rdrobust and RDHonest respectively.
-These methods are representative of current state of the art for inference in RDDs (Stommes, Aronow, and Sävje, 2023), and both offer automated end-to-end confidence interval constructions that can directly be applied to raw data.
+All results are aggregated across 10,000 simulation replications.
+We assess the performance of three commonly used methods by evaluating them on simulations calibrated to nine high-profile applications of RDDs.
+First, we consider conventional local linear regression (Hahn, Todd, and der Klaauw, 2001) with bandwidth chosen as in Imbens and Kalyanaraman (2012).
+We also consider two more recent proposals by Calonico, Cattaneo, and Titiunik (2014) and Armstrong and Kolesár (2018), as implemented in the R functions rdrobust and RDHonest respectively.
+These methods are representative of the current state of the art for inference in RDDs (Stommes, Aronow, and Sävje, 2023), and both offer automated end-to-end confidence interval constructions that can directly be applied to raw data.
 The method rdrobust by Calonico et al.
-(2014) is the dominant approach to regression-discontinuity inference in current econometric practice.
-[^1] This is the only general purpose regression discontinuity package listed in the “Econometrics Task View” for highly used R packages on CRAN (https://cran.r-project.org/web/views/Econometrics.html); and both the R (Calonico et al., 2015) and Stata (Calonico et al., 2017) versions of the function are widely used and cited.
+(2014) is a dominant approach to regression-discontinuity inference in current econometric practice.
+[^2] This is the only general purpose regression discontinuity package listed in the “Econometrics Task View” for highly used R packages on CRAN (https://cran.r-project.org/web/views/Econometrics.html); and both the R (Calonico et al., 2015) and Stata (Calonico et al., 2017) versions of the function are widely used and the original paper (Calonico et al., 2014) cited is over 4500 times as of now.
 The method starts by running a local linear regression as recommended by Hahn, Todd, and der Klaauw (2001), and then applies a 2nd-order bias correction to mitigate curvature bias.
-The method rdrobust offers two options for building confidence intervals—a basic (b.c.) method which inherits the standard error estimator from the original local linear regression (and is justified by asymptotic arguments), and a robust (rob.) method that incorporates a finite-sample variance inflation to account for the bias correction—and we consider both.
+[^3] The method rdrobust offers two options for building confidence intervals: A basic bias-corrected method which inherits the standard error estimator from the original local linear regression (and is justified by asymptotic arguments), and a robust method that incorporates a finite-sample variance inflation to account for the bias correction.
+The latter is the recommended one in the paper and this is the one we consider here.
 Meanwhile, the method RDHonest by Armstrong and Kolesár (2018) explicitly inflates the width of local linear regression confidence intervals to accommodate worst-case plausible curvature bias (rather than trying to correct for this bias).
@@
 However, the method also has a default fall-back option where curvature bounds are derived using the rule-of-thumb of Armstrong and Kolesár (2020); in our experiments we use this default.
-To assess these methods, we use generative adversarial neural networks to generate simulation instances that are representative of real-world applications, as recommended in Athey, Imbens, Metzger, and Munro (2024).
-[^3] The generative adversarial learning approach provides a simulator whose data-generating distribution closely mimics the moments and idiosyncracies of the original data (see Fig.
-1 for an illustration).
-The generated simulator does not use an explicit parametric model and does not enforce any built-in smoothness assumptions, thus bringing it towards a neutral comparison ground for evaluating methods that were motivating under different assumptions.
-We summarize results for 95% confidence intervals built using the considered methods in Table 1; see Section 4 for details on the simulation design.
-A first observation is that rdrobust systematically undercovers the true regression discontinuity parameter.
-The basic (b.c.) intervals only achieve results consistent with nominal coverage in 1 out of 9 settings.
-The finite-sample adjusted intervals (rob.) are wider, but still only achieve nominal coverage in 6 out of the 9 settings.
+To ensure that the evaluations of these methods are credible, we follow Athey, Imbens, Metzger, and Munro (2024) and use generative adversarial neural networks (Goodfellow et al., 2014) to generate simulation instances that are representative of a set of seven seminal regression discontinuity studies.
+[^5] Two of the studies have two outcomes, so we have a total of nine applications.
+The generative adversarial learning approach yields simulators whose data-generating distribution closely mimics the moments and idiosyncrasies of the original datasets; see Figure 1 for an illustration.
+The generated simulators do not use explicit parametric models and do not enforce any built-in smoothness assumptions, thus bringing us towards a neutral comparison ground for evaluating methods that were motivated under different assumptions.
+Figure 1: The original data from one of our benchmark datasets, Lee (2008), as well as a simulated dataset produced using a WGAN trained to the original data.
+This semi-synthetic simulation setup is an important part of our argument.
+It comes with two qualifications.
+First, it relies on the availability of datasets where regression discontinuity methods are applicable.
+We apply it to nine such data sets, but we encourage researchers to apply it to additional ones.
+Second, it relies on our generative adversarial nets yielding realistic data generating processes given the provided datasets.
+Well-known results in non-parametric inference imply that, fixing a sample size and given any dataset where standard regression-discontinuity estimators work well, one can construct another adversarial dataset that is statistically indistinguishable from the first but where the same estimators fail to get coverage (Low, 1997).
+Our approach to evaluation implicitly relies on an assumption that the real datasets we are calibrated to aren’t adversarial in this sense.
+Furthermore, our estimation process involves some choices of tuning parameters for which different researchers may make different choices; in our experiments we seek to mitigate concerns of researcher bias by using a single set of tuning parameters across all datasets.
+We summarize results for 95% confidence intervals in Table 1; see Section 4 for details on the simulation design.
+A first observation is that the conventional local linear regression intervals systematically undercover the true parameter.
+The method rdrobust achieves better—although not perfect—coverage, with nominal rates in 4 out of the 9 settings.
+However, it does so at the expense of wider confidence intervals than conventional local linear regression (18% to 36% wider in our applications).
+RDHonest in turn achieves reliable coverage throughout all designs, but at the cost of still-wider intervals: The median ratio between the width of these intervals and the rdrobust ones varies between 1.06 to 1.76 in our simulations, and between 1.25 to 2.11 relative to the conventional ones.
+Furthermore, these confidence intervals are often excessively conservative, thus suggesting that some of this additional width is not strictly needed.
+[^6] In several settings, RDHonest achieves 98% coverage or above—and this may come at a substantial cost in width.
+For example, in the classical setting with no bias adjustment, Gaussian 98% confidence intervals are 18.7% wider than the conventional 95% intervals.
 These qualitative findings do not appear to be specific to our simulation designs; for example, in Table 3 we observe similar coverage properties when applying the method to pure noise with no relationship between the outcome and the running variable.
-rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
-Dataset EC width EC width EC width EC width
-pure noise 88.7% 0.997 92.8% 1.181 94.9% 1.293 96.3% 0.847
-[…]
-Meanwhile, the method RDHonest achieves valid coverage, but at the cost of substantially wider intervals: The median ratio between the length of these intervals and the basic bias-corrected (b.c.) ones varies between 1.27 to 1.60 in our simulations.
-Furthermore, these confidence intervals are often excessively conservative, thus suggesting that some of this additional width isn’t strictly needed.
-[^4] In several settings, RDHonest achieves 98% coverage or above—and this may come at a substantial cost in width.
-For example, in the classical setting with no bias adjustment, Gaussian 98% confidence intervals are 19% wider than the conventional 95% intervals.
-Our findings leave open a pragmatic question: Is it possible to design automated confidence intervals that reliably achieve nominal coverage in realistic settings, all while matching rdrobust in terms of interval length?
+These semi-synthetic simulation findings leave open a pragmatic question: Is it possible to design automated confidence intervals that reliably achieve nominal coverage in realistic settings, while remaining competitive with the conventional intervals in terms of interval width?
+We emphasize here the “realistic” qualifier.
+Even if it is not feasible to achieve uniformly better properties, it may be that most real-world applications are sufficiently smooth to allow for nominal coverage and narrower confidence intervals than the existing methods.
 The main contribution of this paper is a method, partially linear regression discontinuity inference (PLRD), that provides an affirmative answer to this question—at least in the context of the numerical settings we consider.
@@
 We implement our method in the package plrd for R (R Core Team, 2023), which provides end-to-end functionality for regression discontinuity inference without relying on the user to specify hard-to-choose tuning parameters.
-As seen in Tables 1 and 2, plrd achieves nominal and near-exact coverage throughout, and in every instance yields the shortest intervals among those with valid coverage.
+As illustrated in Tables 1 and 3, our proposed plrd method achieves nominal and near-exact coverage throughout, and in nearly every instance yields the shortest intervals among those with valid coverage.
 2 The PLRD Estimator
@@
 (2014), PLRD starts by estimating $B$ via a cubic regression.
+[^7] We use sample splitting Algorithm 1 to justify technical arguments underlying formal results given in Section 5.
+It is possible, however, that an alternative argument could be used to justify similar results without sample splitting; see, e.g., Armstrong and Kolesár (2020, Appendix E) for results of this type.
   2. 2.
@@
 Our full procedure—including some extra algorithmic details such as cross fitting—is detailed as Algorithm 1.
-Our main formal result about the PLRD estimator is that it yields valid confidence intervals under our assumptions in the large-sample limit.
+As a pragmatic safeguard against failures of Assumption 2, we start our procedure with a hypothesis test, and if we reject this assumption with high confidence then we proceed with a variant of our method that only relies on Assumption 1; see Appendix A for further discussion.
 Algorithm 1 Partially Linear Regression Discontinuity Inference (PLRD)
@@
 Input: Data $\{(X_{i},Y_{i})\}_{i=1}^{n}$, the cutoff $c$, and the confidence level $1-\alpha$ (default is set as $95\%$).
-[…]
+Tuning Parameters: Window size $\ell$ for estimating the smoothness parameter $B$ (default is set as the whole range), a lower bound $\varepsilon$ for $\widehat{B}$ (default is $\varepsilon={\operatorname{SD}[Y]}\,/\,100$), and a confidence level $1-\alpha^{\prime}$ for curvature detection (default is set as $99.9\%$).
 Output: The PLRD estimator $\widehat{\tau}_{\texttt{PLRD}}$ with associated $1-\alpha$ confidence interval $\mathcal{I}_{\alpha}$.
 1:Discard all observations with $|X_{i}-c|>\ell$.
-2:Randomly split the data into two near-equally sized folds, with index sets $I_{1}$ and $I_{2}$.
-3:Use the observations in fold $I_{2}$ to estimate the Lipschitz constant $B$ by fitting cubic polynomials before and after the threshold as in (9), and denote this estimate as $\widehat{B}_{(2)}$.
-4:Enforce a lower bound $\widehat{B}_{(2)}\leftarrow\max\{\widehat{B}_{(2)},\,\varepsilon\}$.
-5:Estimate $\sigma^{2}$ by running the ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $W_{i}$ for $i\in I_{2}$.
+2:Perform an ANOVA test between cubic polynomials fitted with or without different curvatures at significance level $1-\alpha^{\prime}$ to detect whether different curvatures should be used.
+3:Randomly split the data into two near-equally sized folds, with index sets $I_{1}$ and $I_{2}$.
+4:Estimate $\sigma^{2}$ by running the ordinary least-squares regression of $Y_{i}$ on the interaction of $X_{i}$ and $W_{i}$ for $i\in I_{2}$.
 Denote this estimate as $\widehat{\sigma}^{2}_{(2)}$ and record the residuals from this regression as:
 $r_{i}:=Y_{i}-\widehat{\mu}_{W_{i}}(X_{i})\text{ for each }i\in I_{2}.$ (2)
+5:if the test for different curvatures in Step 2 is not rejected then
+6:  Use the observations in fold $I_{2}$ to estimate the Lipschitz constant $B$ by fitting cubic
+7:   polynomials with the same curvature as in (10), and denote this estimate as $\widehat{B}_{(2)}$.
+8:  Enforce a lower bound $\widehat{B}_{(2)}\leftarrow\max\{\widehat{B}_{(2)},\,\varepsilon\}$.
+9:  Let $\mathcal{R}:=\{f:f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is }1\text{-Lipschitz}\}$, and solve the following optimization problem via quadratic programming (see Appendix D for details):
 […]
-$(\widehat{\gamma}_{(1)},\widehat{t}_{(1)}):=\mathop{\rm{minimize}}_{\gamma,t}\left(\widehat{B}^{2}_{(2)}t^{2}+\widehat{\sigma}^{2}_{(2)}\sum_{i\in I_{1}}\gamma_{i}^{2}\right)\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\ \sum_{i\in I_{1}}\gamma_{i}\rho(X_{i}-c)\leq t.$ (3)
-7:Repeat the steps 3 to 5 with the roles of $I_{1}$ and $I_{2}$ swapped.
-8:Set $\widehat{\gamma}_{i}=\gamma_{(k),i}/2$ if $i\in I_{k}$ ($k=1,2$), and $\widehat{\gamma}_{i}=0$ if $|X_{i}-c|>\ell$.
-Compute $\widehat{\tau}_{\texttt{PLRD}}=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i}$, and $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}r_{i}^{2}$ where the $r_{i}$ are the residuals from the least squares regressions in step 4.
-9:Report the following $100(1-\alpha)\%$ confidence interval for $\tau_{c}$:
+10:  Repeat Steps 6 to 9 with the roles of $I_{1}$ and $I_{2}$ swapped.
+11:else
+12:  Use the observations in fold $I_{2}$ to estimate the Lipschitz constant $B$ by fitting cubic
+13:   polynomials with different curvatures (cf. Appendix A), and denote this estimate as $\widehat{B}_{(2)}$.
+14:  Enforce a lower bound $\widehat{B}_{(2)}\leftarrow\max\{\widehat{B}_{(2)},\,\varepsilon\}$.
+15:  Let $\mathcal{R}:=\{f:f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is }1\text{-Lipschitz}\}$, and solve the following optimization problem via quadratic programming (see Appendix D for details):
+[…]
+16:  Repeat Steps 12 to 15 with the roles of $I_{1}$ and $I_{2}$ swapped.
+17:end if
+18:Set $\widehat{\gamma}_{i}=\gamma_{(k),i}/2$ if $i\in I_{k}$ ($k=1,2$), and $\widehat{\gamma}_{i}=0$ if $|X_{i}-c|>\ell$.
+Compute $\widehat{\tau}_{\texttt{PLRD}}=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i}$, and $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}r_{i}^{2}$ where the $r_{i}$ are the residuals from the least squares regressions in (2).
+19:Report the following $100(1-\alpha)\%$ confidence interval for $\tau_{c}$:
 $\mathcal{I}_{\alpha}:=[\widehat{\tau}_{\texttt{PLRD}}-\widehat{h}_{\alpha},\,\widehat{\tau}_{\texttt{PLRD}}+\widehat{h}_{\alpha}],\quad\widehat{h}_{\alpha}:=\min\,\{h:\mathbb{P}(|b+\widehat{s}(\widehat{\gamma})Z|\leq h)\geq 1-\alpha\text{ for all }|b|\leq\widehat{b}\,\}$, (5)
-where $Z\sim\mathcal{N}(0,1)$, and $\widehat{b}:=(\widehat{B}_{(2)}\widehat{t}_{(1)}+\widehat{B}_{(1)}\widehat{t}_{(2)})/2$.
+where $Z\sim\mathcal{N}(0,1)$, and $\widehat{b}:=\big(\widehat{B}_{(2)}\widehat{t}_{(1)}+\widehat{B}_{(1)}\widehat{t}_{(2)}\big)/2$.
+Our main formal result about the PLRD estimator is that it yields valid confidence intervals under our assumptions in the large-sample limit.
+[^9] In our formal results, we assume a continuous running variable with support near the cutoff.
+In the case of discrete running variables, steps 2 and 3 of our procedure remain valid, i.e., given an appropriate $B$ they yield partial identification intervals for $\tau_{c}$ with a guarantee of the type (7) (Imbens and Wager, 2019; Kolesár and Rothe, 2018).
+However, when $X_{i}$ has discrete support, step 1 is no longer guaranteed to consistently recover $B$ under Assumption 1 alone; and so further assumptions would be required to justify data-driven detection of $B$.
 Theorem 1.
-Consider a sharp regression discontinuity design under Assumptions 1 and 2, where the running variable $X_{i}$ has a density strictly positive at the threshold $c$, $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$ is bounded away from 00, i.e., $\sigma_{i}\geq\sigma_{\min}$ for some $\sigma_{\min}>0$, and $\mathbb{E}[|Y_{i}-\mu_{W_{i}}(X_{i})|^{q}\mid X_{i}=x]\leq C$ uniformly over all $x$, for some $q>2$ and $C\in(0,\infty)$.
-Suppose furthermore that we execute the initial cubic regression used to pick $B$ over a pilot bandwidth of length $\ell$ with
+Consider a sharp regression discontinuity design under Assumptions 1 and 2, where the running variable $X_{i}$ has a strictly positive density at the threshold $c$, $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$ is bounded away from 0, i.e., $\sigma_{i}\geq\sigma_{\min}$ for some $\sigma_{\min}>0$, and $\mathbb{E}[|Y_{i}-\mu_{W_{i}}(X_{i})|^{q}\mid X_{i}=x]\leq C$ uniformly over all $x$, for some $q>2$ and $C\in(0,\infty)$.
+Suppose furthermore that we execute the initial cubic regression used to pick $B$ over a pilot bandwidth of width $\ell$ with
 $\lim_{n\rightarrow\infty}n^{1/12}\ell=0,\ \ \ \ \lim_{n\rightarrow\infty}n^{1/7}\ell=+\infty,$ (6)
-and choose any $\varepsilon>0$ in Algorithm 1. Then, level-$(1-\alpha)$ PLRD intervals $\mathcal{I}_{\alpha}$ are asymptotically valid: For any $10<\alpha<$ 1,
+[…]
 $\liminf_{n\to\infty}\,\mathbb{P}[\tau_{c}\in\mathcal{I}_{\alpha}]\geq 1-\alpha.$ (7)
@@
 (2014) and Imbens and Kalyanaraman (2012).
-First, regarding $\ell$, this bandwidth can in practice be chosen simply to be “large”: We only need $\ell$ to decay faster than $n^{-1/12}$, whereas the mean-squred error optimal bandwidth for local linear regression in our setting (i.e., under Assumption 1) scales as $n^{-1/7}$ (Cheng, Fan, and Marron, 1997).
-In our implementation, we by default simply set $\ell$ to contain the full range of the data, and this is choice is what we use in all our experiments.
+First, regarding $\ell$, this bandwidth can in practice be chosen simply to be “large”: We only need $\ell$ to decay faster than $n^{-1/12}$, whereas the mean-squared error optimal bandwidth for local linear regression in our setting (i.e., under Assumption 1) scales as $n^{-1/7}$ (Cheng, Fan, and Marron, 1997).
+In our implementation, we by default simply set $\ell$ to contain the full range of the data, and this choice is what we use in all our experiments.
 [^10] This choice mirrors Imbens and Kalyanaraman (2012, p.
@@
 (2001), who emphasized the identification result (1) and recommended estimation $\tau_{c}$ via local linear regression,
-$(\widehat{a}_{w},\widehat{\beta}_{w})=\mathop{\rm argmin}_{(a,\beta)\in\mathbb{R}^{2}}\ \left\{\sum_{i\,:\,W_{i}=w}K\left(\frac{|X_{i}-c|}{h_{n}}\right)\left(Y_{i}-a-\beta(X_{i}-c)\right)^{2}\right\},\ \ \ \ \widehat{\tau}_{\texttt{LLR}}=\widehat{a}_{1}-\widehat{a}_{0},$ (7)
-where $h_{n}\rightarrow 0$ is a bandwidth that “localizes” the regression and $K(\cdot)$ is a kernel function with support on [1, 1]11[-1,\,1][ - 1 , 1 ].
+$(\widehat{a}_{w},\widehat{\beta}_{w})=\mathop{\rm argmin}_{(a,\beta)\in\mathbb{R}^{2}}\ \left\{\sum_{W_{i}=w}K\left(\frac{|X_{i}-c|}{h_{n}}\right)\left(Y_{i}-a-\beta(X_{i}-c)\right)^{2}\right\},\ \ \ \ \widehat{\tau}_{\texttt{LLR}}=\widehat{a}_{1}-\widehat{a}_{0},$ (8)
+where $h_{n}\rightarrow 0$ is a bandwidth that “localizes” the regression and $K(\cdot)$ is a kernel function with support on [1, 1][-1,\,1].
 In the decade following Hahn et al.
-(2001), local linear (or polynomial) regression remained the prevalent approach approach to inference in RDDs, with confidence intervals produced by applying off-the-shelf tools for heteroskedasticity-robust inference directly to the above regression (Imbens and Lemieux, 2008; Lee and Lemieux, 2010).
+(2001), local linear (or polynomial) regression remained the prevalent approach to inference in RDDs, with confidence intervals produced by applying off-the-shelf tools for heteroskedasticity-robust inference directly to the above regression (Imbens and Lemieux, 2008; Lee and Lemieux, 2010).
 A limitation of this approach, however, is that it leads to contradictory guidance on the bandwidth choice under a non-parametric specification as given in, e.g., Assumption 1.
-The accuracy of $\widehat{\tau}_{\texttt{LLR}}$ depends delicately on the bandwidth $h_{n}$ (Imbens and Kalyanaraman, 2012) and, unfortunately, for choices of $h_{n}$ that optimize this accuracy the bias and standard error of $\widehat{\tau}_{\texttt{LLR}}$ are of the same order—so off the shelf local linear regression inference will not have valid coverage for this choice of $h_{n}$.
+The accuracy of $\widehat{\tau}_{\texttt{LLR}}$ depends delicately on the bandwidth $h_{n}$, with a first systematic proposal given in Imbens and Kalyanaraman (2012).
+For choices of $h_{n}$ that optimize this accuracy the bias and standard error of $\widehat{\tau}_{\texttt{LLR}}$ are of the same order—so off-the-shelf local linear regression inference will not have valid coverage for this choice of $h_{n}$.
 The practitioner is thus forced to invoke “undersmoothing”, whereby they voluntarily use a non-accuracy-optimizing choice of $h_{n}$ to justify validity of their inferential procedure.
@@
 (2014) is to estimate and correct for the bias of local linear regression by leveraging higher-order smoothness; we refer to this approach as the “bias-corrected” approach.
+[^12] When reporting results using the bias-correct approach, we use robust standard error estimates as recommended by Calonico et al.
+(2014) and as implemented in the package rdrobust (Calonico et al., 2015; Calonico et al., 2017).
 Calonico, Cattaneo, Farrell, and Titiunik (2019) discuss the use of covariates under the bias-corrected approach to RDDs while Cattaneo and Titiunik (2022) provide a recent review.
@@
 (2014) and follow-ups focus entirely on asymptotic validity results—as we do here.
-They fix a data-generating process, and guarantee that they eventually get coverage to this data-generating process as the sample size gets large (as is also done in (6)).
+They fix a data-generating process, and guarantee that they eventually get coverage for this data-generating process as the sample size gets large (as is also done in (7)).
 In contrast, most papers written under the bias-aware paradigm specify a quantitative smoothness class for the $\mu_{w}(x)$ (e.g., all functions with a second derivative bounded by a pre-specified constant $B$), and then seek guarantees that hold in finite samples and uniformly over this whole class.
 Getting uniform guarantees in finite samples of course requires stronger assumptions than getting pointwise asymptotic guarantees; and for this reason papers associated with the bias-aware paradigm often start by making stronger assumptions than those associated with the bias-corrected paradigm.
-We emphasize, however, that this is due to the sought formal guarantees—and it’s not that bias-aware methods inherently require stronger assumptions to work than bias-corrected methods.
+We emphasize, however, that this is due to the sought formal guarantees; and it’s not that bias-aware methods inherently require stronger assumptions to work than bias-corrected methods.
 Our proposed method, PLRD, straddles these two literatures.
 Our confidence intervals are rooted in the bias-aware paradigm of Armstrong and Kolesár (2018), and we use numerical convex optimization to construct our estimator as in Imbens and Wager (2019).
-In contrast, our formal results are focused on pointwise asymptotics, and the assumptions and guarantee type given in our Theorem 1 closely mirror those in Theorem 1 of Calonico et al.
+In contrast, our formal results are focused on pointwise asymptotics, and the assumptions and guarantee type given in our Theorem 1 mirror those in Theorem 1 of Calonico et al.
 (2014).
 One notable paper that also straddles these literatures in Armstrong and Kolesár (2020), who propose bias-aware confidence intervals for local linear regression with pointwise asymptotic guarantees.
-In particular, they show that if one runs local linear regression as in (7) with the mean-square-error optimal bandwidth choice $h_{n}$ following Imbens and Kalyanaraman (2012), then under generic conditions we can build valid, bias-aware 95% confidence intervals for $\tau_{c}$ as $\widehat{\tau}_{\texttt{LLR}}\pm 2.18\times\text{standard errors}$ rather than the conventional $\widehat{\tau}_{\texttt{LLR}}\pm 1.96\times\text{standard errors}$.
-We provide a comparison of PLRD to this simple rule of thumb in Appendix A.
+In particular, they show that if one runs local linear regression as in (8) with a mean-square-error optimal bandwidth choice $h_{n}$, then under generic conditions we can build valid, bias-aware 95% confidence intervals for $\tau_{c}$ as $\widehat{\tau}_{\texttt{LLR}}\pm 2.18\times\text{standard errors}$ rather than the conventional $\widehat{\tau}_{\texttt{LLR}}\pm 1.96\times\text{standard errors}$.
+We provide a comparison of PLRD to this approach in Appendix B.
 Wrapping up, PLRD can best be understood within the context of earlier developments in the bias-aware paradigm; and in fact steps (2) and (3) of our algorithmic outline fall squarely within the roadmap for minimax linear inference via numerical convex optimization as presented in Imbens and Wager (2019).
@@
   • We rely on higher-order smoothness and a preliminary cubic fit to initialize tuning parameters of our algorithm; in Armstrong and Kolesár (2018), Kolesár and Rothe (2018) and Imbens and Wager (2019), a parameter analogous to the $B$ derived in our step (1) would need to be pre-specified using subject matter knowledge or chosen using a heuristic.
-  • Because of this data-driven choice of $B$ in step (1), we are not able to provide uniform, finite-sample inferential guarantees; [^8] And, as shown in Armstrong and Kolesár (2018), this would in fact be impossible.
+  • Because of this data-driven choice of $B$ in step (1), we are not able to provide uniform, finite-sample inferential guarantees; [^13] As shown in Armstrong and Kolesár (2018), doing so would in fact be impossible under smoothness assumptions like Assumption 1 alone.
+We do note that proving uniform bounds for methods involving data-driven smoothness guarantees can be possible under more restricted function classes, e.g., Armstrong and Kolesár (2020, Appendix E) consider such results over function classes where global polynomial approximations are guaranteed to capture worst-case local curvature of the underlying functions.
+We leave a study of uniformity guarantees for PLRD over restricted function classes to future work.
 thus, like in Calonico et al.
@@
 Given the empirical performance of PLRD on calibrated simulations, we suggest that our $B$-Lipschitz curvature class with linear treatment effects may be a good default choice in many applications.
-We conduct an ablation analysis to examine the effect of different choices on the behavior of our estimator in Appendix A.
+We conduct an ablation analysis to examine the effect of different choices on the behavior of our estimator in Appendix B.
 4 Empirical Study
 We now describe in more detail the experiment presented in the introduction (Table 1).
-As context for this experiment, Table 3 shows a replication of a number of high-profile RDD applications using a variety of methods; citations for each application are given in the leftmost column and further description of each benchmark dataset is provided Appendix B.
-We immediately see that the confidence intervals given by rdrobust, especially with the ${\it(b.c.)}$ interval choice, are much shorter than those given by RDHonest, thus naturally making the method attractive to practitioners.
-But how much can we trust these intervals?
-Dataset rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
-Lee (2008) $0.059\pm 0.022$ $0.059\pm 0.059$ $0.058\pm 0.032$ $0.079\pm 0.020$
-Cattaneo et al. (2015) $7.507\pm 2.859$ $7.507\pm 3.413$ $8.030\pm 4.210$ $6.440\pm 2.374$
-Matsudaira (2008) (read.) $0.030\pm 0.028$ $0.030\pm 0.032$ $0.031\pm 0.046$ $0.029\pm 0.033$
-Matsudaira (2008) (math) $0.072\pm 0.025$ $0.072\pm 0.030$ $0.057\pm 0.041$ $0.077\pm 0.026$
-Jacob and Lefgren (2004) (read.) $0.143\pm 0.042$ $0.143\pm 0.048$ $0.127\pm 0.046$ $0.093\pm 0.037$
-Jacob and Lefgren (2004) (math) $0.203\pm 0.035$ $0.203\pm 0.039$ $0.187\pm 0.058$ $0.152\pm 0.032$
-Oreopoulos (2006) $0.099\pm 0.104$ $0.099\pm 0.143$ $0.070\pm 0.125$ $0.072\pm 0.115$
-Ludwig and Miller (2007) $-2.781\pm 2.363$ $-2.781\pm 2.682$ $-3.153\pm 2.828$ $-1.824\pm 2.019$
-Meyersson (2014) $2.983\pm 2.797$ $2.983\pm 3.293$ $3.024\pm 3.109$ $3.162\pm 2.294$
+As context for this experiment, Table 2 shows a replication of a number of high-profile RDD applications using a variety of methods; citations for each application are given in the leftmost column and further description of each benchmark dataset is provided Appendix C.
+Dataset conventional rdrobust RDHonest plrd
+Lee (2008) $064\pm 0.022 0.0$$59\pm 0.025 0.0$$58\pm 0.032 0.0$$073\pm 0.024$
+Cattaneo et al. (2015) $7.414\pm 2.859$ $7.507\pm 3.413$ $8.030\pm 4.210$ $5.830\pm 2.127$
+Ludwig and Miller (2007) $2.409\pm 2.363$ $2.781\pm 2.682$ $3.153\pm 2.828$ $1.249\pm 1.507$
+Meyersson (2014) $3.020\pm 2.797$ $2.983\pm 3.293$ $3.024\pm 3.109$ $3.010\pm 2.093$
+Matsudaira (2008) (read.) $036\pm 0.028 0.0$$30\pm 0.032 0.0$$31\pm 0.046 0.0$$017\pm 0.033$
+Matsudaira (2008) (math) $075\pm 0.025 0.0$$72\pm 0.030 0.0$$57\pm 0.041 0.0$$073\pm 0.028$
+Jacob and Lefgren (2004) (read.) $0.129\pm 0.042$ $0.143\pm 0.048$ $127\pm 0.046 0.0$$099\pm 0.040$
+Jacob and Lefgren (2004) (math) $0.195\pm 0.035$ $0.203\pm 0.039$ $0.187\pm 0.058$ $0.168\pm 0.039$
+Oreopoulos (2006) $070\pm 0.104 0.0$$99\pm 0.143 0.0$$70\pm 0.125 0.0$$021\pm 0.064$
 Table 2: 95% confidence intervals produced by different methods on cited benchmark datasets.
-Figure 1: The original data from one of our benchmark datasets, Lee (2008), as well as a simulated dataset produced using a WGAN trained to the original data.
 A fundamental challenge in assessing methods for causal inference on real-world data is that we typically do not have access to a ground truth we can use for evaluation.
@@
   1. 1.
-    Training: We train WGANs on each benchmark dataset to mimic the underlying data generating process. This involves generating running variables and outcomes for both treated and control units.
+    Training: We train WGANs on each benchmark dataset to mimic the underlying data generating process for both treated and control units.
   2. 2.
@@
 As illustrated in Figure 1, a trained WGAN is able to simulate data whose distribution is very similar to that of the original data; thus suggesting that if a method gets reliable coverage on the simulator we can also trust it on the real data.
-The WGANs take as input random Gaussian noise $Z\in\mathbb{R}^{3}$.
-We train one generator $\widehat{g}_{x}(x)$ for the unconditional distribution of the running variable $X$, and two generators $\widehat{g}_{w}(x,z)$ ($w=0,1$) for the conditional distribution of the outcomes for treated and control units, given the running variable.
-We can calculate the ground truth $\widetilde{\tau}$ for the GAN estimated distribution by sampling random noise $Z_{q}\in\mathbb{R}^{3}$ for $q=1,\dots,Q$, where $Q$ is a large number ($10^{6}$ in our simulations), and generating outcomes at the threshold, as follows:
-$\widetilde{\tau}:=Q^{-1}\sum_{q=1}^{Q}(\widehat{g}_{1}(0,Z_{q})-\widehat{g}_{0}(0,Z_{q})).$
-Finally, we generate $i=1,\dots,N=10^{3}$ simulated datasets with confidence intervals $[a_{i,k},b_{i,k}]$, where the index $k$ stands for the $k$-th method we use to produce these CIs.
-The empirical coverage for the $k$-th method is given by $\widehat{p}_{k}=\frac{1}{N}\sum_{i=1}^{n}\mathbf{1}\{a_{i,k}\leq\widetilde{\tau}\leq b_{i,k}\}$.
+Our WGAN training procedure consists of the following steps.
+First, apply marginal CDF transforms to the data to convert $(X_{i},Y_{i})$ to $(U_{i},V_{i})$, where $U_{i}=F_{x}(X_{i})$ and $V_{i}=F_{y}(Y_{i})$.
+For discrete data we break ties by randomization, and for additional stability we use a Gaussian transformation.
+We train two generators $\widehat{g}_{w}(u,z)$ ($w=0,1$), one for treated units and one for controls, that mimic the conditional distribution of the CDF-transformed outcome $V$ given the CDF-transformed running variable $U$.
+Finally, we invert the CDF-transforms to obtain $(\widetilde{X}_{i},\widetilde{Y}_{i})$ from the pair $(\widetilde{U}_{i},\widetilde{V}_{i})$ generated by the trained WGAN.
+We calculate the ground truth $\widetilde{\tau}$ for the GAN estimated distribution as follows.
+We first calculate the cutoff $c^{\prime}$ under the applied CDF transforms.
+We then sample random noise $Z_{0,q},Z_{1,q}\in\mathbb{R}$ for $q=1,\dots,Q$, where $Q$ is a large number ($10^{7}$ in our simulations), generate CDF-transformed outcomes at the threshold $c^{\prime}$, and invert them:
+$\widetilde{\tau}:=Q^{-1}\sum_{q=1}^{Q}\left(F_{y}^{-1}(\widehat{g}_{1}(c^{\prime},Z_{1,q}))-F_{y}^{-1}(\widehat{g}_{0}(c^{\prime},Z_{0,q}))\right).$
+We generate $N=10^{4}$ simulated datasets indexed by $i=1,\dots,N$, and compute confidence intervals $a_{i,k},b_{i,k}]$, for each method $k$.
+The empirical coverage for the $k$-th method is given by $\widehat{p}_{k}=N^{-1}\sum_{i=1}^{n}\mathbf{1}\{a_{i,k}\leq\widetilde{\tau}\leq b_{i,k}\}$.
 Table 1 shows the obtained results.
 4.1 Revisiting existing simulation designs
-For completeness, we also benchmark the PLRD estimator against other existing methods on the simulation settings originally used in Calonico et al.
+conventional rdrobust RDHonest plrd
+coverage width coverage width coverage width coverage width
+Pure noise 92.6% 1.419 93.1% 1.681 96.0% 1.854 97.2% 1.011
+Setting 1 88.4% 0.203 90.9% 0.240 95.4% 0.288 94.8% 0.237
+Setting 2 79.7% 0.266 90.5% 0.296 96.2% 0.440 97.3% 0.389
+Setting 3 90.1% 0.213 92.7% 0.246 96.2% 0.301 99.6% 0.290
+Setting 4 91.8% 0.201 92.3% 0.238 96.6% 0.276 96.1% 0.178
+Table 3: Empirical coverage and average width of 95% confidence intervals produced by different methods on pure noise and on simulations from Calonico et al.
+(2014).
+All simulations have sample size $n=500$, and results are aggregated across 10,000 replications.
+Methods are run as in Table 1.
+For completeness, we also benchmark the PLRD estimator against other existing methods on some standard simulation designs.
+First, we consider a pure noise setting where we generate the running variables as $X_{i}\sim\text{Unif}([-1,\,1])$, the outcomes as $Y_{i}\sim\mathcal{N}(0,\,1)$, and set treatment to $W_{i}=1(\{X_{i}\geq 0\})$.
+Next, we consider simulation settings used in Calonico et al.
 (2014, Section 6).
-[^9] Of these, Setting 1 was first used by Imbens and Kalyanaraman (2012), while Settings 2 and 3 were original to Calonico et al.
-(2014).
-In each of these settings, we draw the running variable $X_{i}\sim 2\,\text{Beta}(2,4)-1$ for $n=500$ samples, and the outcome is given by $Y_{i}=m(X_{i})+\varepsilon_{i}$, where $\varepsilon_{i}$ are mean-zero Gaussian noise with variance $0.1295^{2}$.
+In each of these settings, we draw the running variable $X_{i}\sim 2\,\text{Beta}(2,4)-1$, and the outcome is given by $Y_{i}=m(X_{i})+\varepsilon_{i}$, where $\varepsilon_{i}$ are mean-zero Gaussian noise with variance $0.1295^{2}$.
 The form of $m(\cdot)$ varies for the different simulation settings; see Calonico et al.
@@
 Setting 4 uses the function $m=m_{\text{quad}}$ as in Imbens and Kalyanaraman (2012, Section 4.6).
-Figure 2 plots one realization for each of these four different simulation settings.
+We note that the linear treatment effect condition (Assumption 2) is not satisfied in several of these designs, so we expect our pragmatic safeguard (given in Algorithm 1) to play an important role in delivering robust performance.
 Table 3 presents the empirical coverage and average width of 95% CIs produced by different methods across these simulation settings.
@@
 (2014, Table 1).
-RDHonest delivers valid coverage at the expense of longer coverage.
-And, in Settings 1, 3 and 4, plrd delivers valid coverage with intervals whose length is competitive with rdrobust.
-Our method, however, does not perform well in Setting 2.
-The reason for this poor performance is immediately revealed by Fig.
-2: Our method relies on the linear treatment effect assumption (Assumption 2), whereas Setting 2 has strong curvature in the treatment effect that is manifested in a sharp jump in curvature at the cutoff.
-If one believes that such a sharp jump in the curvature change is realistic, one should not rely on Assumption 2; and in Appendix A we detail a variant of plrd without this assumption (and verify that it performs well in Setting 2 of Calonico et al.
-(2014)).
-Overall, however, we still rely on Assumption 2 in the default specification of plrd as we believe such sharp curvature effects in $\tau(x)$ to be unusual in most applications.
-[^10] The Setting 2 of Calonico et al.
-(2014) was originally derived by fitting a 5th-order polynomial to data from Ludwig and Miller (2007).
-However, unlike with our WGAN approach, the final simulation distribution as displayed in Figure 2(b) does not resemble the original data as summarized in Ludwig and Miller (2007, Figure IV(A)) in terms of, e.g., its signal-to-noise ratio.
-In the original dataset, the datapoints are counties; the running variable is poverty rate; the cutoff is associated with head-start eligibility; and the outcome is child mortality rate (per 100,000 children) from causes potentially associated with Head Start.
-(a) Setting 1
-(b) Setting 2
-(c) Setting 3
-(d) Setting 4
-Figure 2: Visualization of four simulation settings from Calonico et al.
-(2014, Section 6) and Imbens and Kalyanaraman (2012, Section 4.6).
-In each setting, the conditional regression surface $m(\cdot)$ is displayed as a solid line, and the points represent data for one simulation realization.
-rdrobust (b.c.) rdrobust (rob.) RDHonest plrd
-coverage width coverage width coverage width coverage width
-Setting 1 86.3% 0.207 91.5% 0.244 95.2% 0.290 94.6% 0.235
-Setting 2 91.1% 0.313 93.2% 0.342 96.9% 0.436 51.5% 0.309
-Setting 3 88.6% 0.218 93.1% 0.250 95.7% 0.301 95.9% 0.239
-Setting 4 87.8% 0.207 93.0% 0.246 96.8% 0.275 97.5% 0.236
-Table 4: Empirical coverage and average width of 95% confidence intervals produced by different methods for the simulation settings in Calonico et al.
-(2014) and Imbens and Kalyanaraman (2012).
-Methods are run as in Table 1.
-These simulation results highlight the strengths and limitations of our plrd method: It performs well in settings that align with its underlying assumptions and can be adapted to handle more complex scenarios.
-However, users should be cautious when applying plrd to settings with potential highly nonlinear treatment effects $\tau(x)$ (non-linearity in the baseline response function $\mu_{0}(x)$ is not a problem, as long as it is captured by the standard conditions required in Assumption 1).
+RDHonest delivers valid coverage at the expense of wider intervals, and plrd delivers valid coverage with intervals whose width is competitive with rdrobust.
 5 Formal Results
@@
 272) for further discussion.
-$\widehat{\tau}_{\texttt{PLRD}}:=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i},\quad\widehat{\gamma}:=\mathop{\rm argmin}_{\gamma}\underbrace{\widehat{\sigma}^{2}\lVert\gamma\rVert_{2}^{2}}_{\text{variance proxy}}+\underbrace{\sup_{\mu_{0},\,\mu_{1}\in\,\mathcal{M}_{\widehat{B}}}\left(\sum_{i=1}^{n}\widehat{\gamma}_{i}\mu_{W_{i}}(X_{i})-(\mu_{1}(c)-\mu_{0}(c))\right)^{2}}_{\text{worst-case squared bias over}\mathcal{M}_{\widehat{B}}}.$ (8)
+$\widehat{\tau}_{\texttt{PLRD}}:=\sum_{i=1}^{n}\widehat{\gamma}_{i}Y_{i},\quad\widehat{\gamma}:=\mathop{\rm argmin}_{\gamma}\underbrace{\widehat{\sigma}^{2}\lVert\gamma\rVert_{2}^{2}}_{\text{variance proxy}}+\underbrace{\sup_{\mu_{0},\,\mu_{1}\in\,\mathcal{M}_{\widehat{B}}}\left(\sum_{i=1}^{n}{\gamma}_{i}\mu_{W_{i}}(X_{i})-(\mu_{1}(c)-\mu_{0}(c))\right)^{2}}_{\text{worst-case squared bias over }\mathcal{M}_{\widehat{B}}}$.
+(9)
 If $\widehat{B}$ were a constant chosen a-priori and $\mu^{\prime\prime}_{0}(x)$ were in fact is globally $\widehat{B}-$Lipschitz, we could use existing results for minimax linear regression-discontinuity inference (e.g., Armstrong and Kolesár, 2018; Imbens and Wager, 2019) to directly provide guarantees for $\widehat{\tau}_{\mathtt{PLRD}}$.
-But here, of course, this is not the case: Our choice of $\widehat{B}$ is learned in step 1 of our algorithm; and Assumption 1 only implies local—not global—Lipschitz continuity in $\mu^{\prime\prime}_{0}(x)$
+But here, of course, this is not the case: Our choice of $\widehat{B}$ is learned in step 1 of our algorithm, and Assumption 1 only implies local—not global—Lipschitz continuity in $\mu^{\prime\prime}_{0}(x)$ .
 The goal of this section is to provide technical tools to address these challenges, culminating in the proof of Theorem 1.
@@
 Proposition 2.
-Suppose that Assumptions 1 and 2 hold and that $X_{i}$ has a density bounded away from 0 in a neighborhood of $c$, and $B:=\max\{|\mu^{\prime\prime\prime}(c)|,\varepsilon\}$.
-Estimate $\widehat{B}$ via local cubic regression: For $w\in\{0,1\}$ run
-$\widehat{\beta}_{w}:=\mathop{\rm argmin}_{\beta_{w}\in\mathbb{R}^{4}}\sum_{\begin{subarray}{c}W_{i}=w\\ |X_{i}-c|\leq\ell\end{subarray}}\left(Y_{i}-\beta_{w,0}-\beta_{w,1}(X_{i}-c)-\frac{\beta_{w,2}}{2}(X_{i}-c)^{2}-\frac{\beta_{w,3}}{6}(X_{i}-c)^{3}\right)^{2},$ (9)
-with bandwidth $\ell$ scaling as in (5), and set $\widehat{B}=\max\{|\beta_{0,3}|,|\beta_{1,3}|,\varepsilon\}$. Then, $\widehat{B}\rightarrow_{p}B$ as $n\to\infty$.
+Suppose that Assumptions 1 and 2 hold and that $X_{i}$ has a density bounded away from 0 in a neighborhood of $c$, and $B:=\max\{|\mu_{0}^{\prime\prime\prime}(c)|,|\mu_{1}^{\prime\prime\prime}(c)|,\varepsilon\}$.
+Run the local cubic regression:
+$\widehat{\beta}:=\mathop{\rm argmin}_{\beta\in\mathbb{R}^{6}}\sum_{\begin{subarray}{c}|X_{i}-c|\leq\ell\end{subarray}}\left(Y_{i}-\beta_{W_{i},0}-\beta_{W_{i},1}(X_{i}-c)-\frac{\beta_{2}}{2}(X_{i}-c)^{2}-\frac{\beta_{3}}{6}(X_{i}-c)^{3}\right)^{2},$ (10)
+with bandwidth $\ell$ scaling as in (6), and set [^15] As an additional safe-guard against small sample sizes, we calculate a standard error estimate $\mathrm{s.e.}(\widehat{\beta}_{3})$ and use the conservative yet asymptotically consistent estimate $\widehat{B}=\max\{|\widehat{\beta}_{3}\pm 1.96\cdot\mathrm{s.e.}(\widehat{\beta}_{3})|,\varepsilon\}$ in our software.
+$\widehat{B}=\max\{|\beta_{3}|,\varepsilon\}$.
+Then, $\widehat{B}\rightarrow_{p}B$ as $n\to\infty$.
 Our core analytic result is the following central limit theorem for our PLRD estimator. Asymptotic validity of the PLRD confidence intervals as claimed in Theorem 1 then follows immediately.
@@
 Under the conditions of Theorem 1, $\widehat{\tau}_{\mathtt{PLRD}}$ as constructed in Algorithm 1 satisfies
+$\left(\widehat{\tau}_{\mathtt{PLRD}}-b(\widehat{\gamma})-\tau_{c}\right)\,\big/\,{s}(\widehat{\gamma})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,\,1),\ \ \ \ {s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\sigma_{i}^{2},\ \ \ \ b(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}\,\mu_{W_{i}}(X_{i})-\tau_{c},$ (11)
 […]
-where $b(\widehat{\gamma})$ is interpreted as the conditional bias of $\widehat{\tau}_{\mathtt{PLRD}}$, and $\sigma_{i}^{2}={\rm Var}(Y_{i}\mid X_{i})$. Furthermore,
 $\limsup_{n\to\infty}\left(\left|b(\widehat{\gamma})\right|-\widehat{b}\right)\,\big/\,{s}(\widehat{\gamma})\leq_{p}0,$ (12)
@@
 5.1 Proof of Theorem 1
-Define $\widehat{s}^{2}(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\widehat{\sigma}_{i}^{2}$ as in Algorithm 1.
-It follows from White (1980) that $\widehat{s}(\widehat{\gamma})/s(\widehat{\gamma})\rightarrow_{p}1$; see Lemma 14 for details.
-Next, fix any $\delta\in(0,1)$.
-In view of (11) and Slustky’s lemma, $\mathbb{P}(A_{n})\to 1$ where $A_{n}:=\{|b(\widehat{\gamma})|\leq\widehat{b}+\delta\widehat{s}(\widehat{\gamma})\}$.
-We show in Lemma 13 that for any $s$, $h$, $t$ and $\delta>0$,
+[…]
 $\inf_{|b|\,\leq\,t+\delta s}\mathbb{P}(|b+sZ|\leq h)\geq\inf_{|b|\,\leq\,t}\mathbb{P}(|b+sZ|\leq h)-\delta,$
-where $Z\sim\mathcal{N}(0,1)$. Using this with $s=\widehat{s}(\widehat{\gamma})$, $h=\widehat{h}_{\alpha}$ and $t=\widehat{b}$, we get, for every $n$,
-$\mathbb{P}\left(|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\}\cap A_{n}\right)\geq\inf_{|b|\,\leq\,\widehat{b}}\mathbb{P}(|b+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha})-\delta\geq 1-\alpha-\delta,$ (12)
+[…]
+$\begin{split}\text{$\mathbf{1}$}_{A_{n}}\mathbb{P}\left(|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\,\Big|\,(X_{i},Y_{i})_{i=1}^{n}\right)&\geq\text{$\mathbf{1}$}_{A_{n}}\inf_{|b|\,\leq\,\widehat{b}\,+\,\delta\widehat{s}(\widehat{\gamma})}\mathbb{P}\left(|b+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\right)\\ &\geq\text{$\mathbf{1}$}_{A_{n}}\left(\inf_{|b|\,\leq\,\widehat{b}}\mathbb{P}(|b+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha})-\delta\right)\\ &\geq\text{$\mathbf{1}$}_{A_{n}}(1-\alpha-\delta),\end{split}$ (13)
 where the last inequality follows from the definition of $\widehat{h}_{\alpha}$.
@@
 $\liminf_{n\to\infty}\mathbb{P}(\tau_{c}\in\mathcal{I}_{\alpha})$ $=\liminf_{n\to\infty}\mathbb{P}(|\widehat{\tau}_{\texttt{PLRD}}-\tau_{c}|\leq\widehat{h}_{\alpha})$
-$=\liminf_{n\to\infty}\mathbb{P}(\{|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z_{n}|\leq\widehat{h}_{\alpha}\}\cap A_{n})$ (since $\mathbb{P}(A_{n})\to 1$)
-$\geq 1-\alpha-\delta.$ (using (12) and Glivenko–Cantelli lemma)
+$=\liminf_{n\to\infty}\mathbb{P}\left(|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\right)$
+$\geq\liminf_{n\to\infty}\mathbb{P}(\{|b(\widehat{\gamma})+\widehat{s}(\widehat{\gamma})Z|\leq\widehat{h}_{\alpha}\}\cap A_{n})$
+$\geq 1-\alpha-\delta.$ (using (13) and $\mathbb{P}(A_{n})\to 1$)
 Since $\delta\in(0,1)$ is arbitrary, we are through.
 5.2 Proof of Theorem 3
-The optimization problem in Algorithm 1 is inherently related to its convex dual (see Boyd and Vandenberghe (2004) for a definition). We show in Appendix C that the dual problem is given by
+The optimization problem in Algorithm 1 is inherently related to its convex dual (see Boyd and Vandenberghe (2004) for a definition). We show in Appendix D that the dual problem is given by
 $\begin{split}&\widehat{\theta}_{n}(\sigma^{2},B)=\mathop{\rm argmin}_{\theta=({\rho},\lambda)\in\Theta}\widehat{L}_{n}(\theta;\sigma^{2},B),\\ &\widehat{L}_{n}(\theta;\sigma^{2},B):=\frac{1}{4n\sigma^{2}}\sum_{i=1}^{n}G^{2}(\theta;X_{i}-c,W_{i})+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n},\end{split}$ (14)
@@
 In Algorithm 1, we use sample splitting to estimate $B$ (cf.
-Proposition 2) and the homoscedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$.
-However, for ease of exposition, we assume throughout this section that we have $\widehat{B}\rightarrow_{p}B$ and $\widehat{\sigma}^{2}\rightarrow_{p}\sigma^{2}$ independent of the data, and
-$\widehat{\theta}_{n}:=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})=\mathop{\rm argmin}_{\theta\,\in\,\Theta}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B}).$
+Proposition 2) and the conditional variance $\sigma^{2}=\mathbb{E}[{\rm Var}(Y_{i}\mid X_{i})]$.
+However, for ease of exposition, we assume throughout this section that we have $\widehat{B}\rightarrow_{p}B$ and $\widehat{\sigma}^{2}\rightarrow_{p}\sigma^{2}$ independently of the data, and
+[…]
 We can recover the primal solution $(\widehat{\gamma}_{n},\widehat{t}_{n}\,)$ from the dual solution $\widehat{\theta}_{n}$ as:
@@
 To tackle these issues, we borrow tools from empirical process theory (see, e.g., van der Vaart and Wellner (1996) for a textbook treatment).
-The following result, proven in Appendix D, characterizes the rate of convergence of the empirical dual optimizer $\widehat{\theta}_{n}$ to the population dual optimizer $\theta_{n}^{*}$; and implies (Corollary 5) that the weights underlying $\widehat{\tau}_{\mathtt{PLRD}}$ satisfy a Lindeberg-type condition.
+The following result, proven in Appendix E, characterizes the rate of convergence of the empirical dual optimizer $\widehat{\theta}_{n}$ to the population dual optimizer $\theta_{n}^{*}$; and implies (Corollary 5) that the weights underlying $\widehat{\tau}_{\mathtt{PLRD}}$ satisfy a Lindeberg-type condition.
 Lemma 4.
-Let $(\widehat{\rho}_{n},\widehat{\lambda}_{n})=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})$ be the optimizer of the empirical dual problem (13), and $({\rho}_{n}^{*},\lambda_{n}^{*})=\theta_{n}^{*}(\widehat{\sigma}^{2},B)$ be the optimizer of the population dual problem (15).
-Assume that the pilot bandwidth $\ell$ in Algorithm 1 satisfies (5), and $\widehat{B}\rightarrow_{p}B$.
-Then, under the conditions of Theorem 1,
 […]
+$\|\widehat{\rho}_{n}-\rho_{n}^{*}\|^{2}_{L^{2}(P)}+\|\widehat{\lambda}_{n,-1}-\lambda_{n,-1}^{*}\|_{2}^{2}=\mathcal{O}_{p}\left(\frac{\ell^{\,6}}{n^{1/4+6/7}}+\frac{\ell^{\,3}}{n^{1+3/7}}\right).$
 Corollary 5.
@@
 Corollary 5 immediately implies that the central limit theorem claimed in Theorem 3 holds; see the proof of Imbens and Wager (2019, Theorem 1) for details.
-We delegate showing that $s^{2}(\widehat{\gamma})=\mathcal{O}_{p}(n^{-6/7})$ (Lemma 11) and $\widehat{b}/{s}(\widehat{\gamma})=\mathcal{O}_{p}(1)$ (Lemma 12) to Appendix D.
+We delegate showing that $s^{2}(\widehat{\gamma})=\mathcal{O}_{p}(n^{-6/7})$ (Lemma 12) and $\widehat{b}/{s}(\widehat{\gamma})=\mathcal{O}_{p}(1)$ (Lemma 13) to Appendix E.
 Finally, to verify (12), we note that Assumptions 2 and 1 yield $b(\widehat{\gamma})=\sum_{i=1}^{n}\widehat{\gamma}_{i}\mu_{W_{i}}(X_{i})=B\sum_{i=1}^{n}\widehat{\gamma}_{i}\rho(X_{i}-c)$, where
@@
 and so (18) implies (12).
-Appendix A Details on Baseline Methods
+Appendix A Safeguard against Curvature Changes
+While Assumption 2 provides a parsimonious and empirically realistic model for treatment effect heterogeneity in many applications, it can in principle be violated.
+To safeguard against such violations, our software incorporates an automatic diagnostic that tests for difference in curvatures.
+The diagnostic works as follows.
+Before running the main PLRD procedure, we conduct an ANOVA test comparing a restricted cubic polynomial model (where treatment interacts only linearly with $X_{i}-c$, consistent with Assumption 2) against an unrestricted model (where treatment interacts with all polynomial terms in $X_{i}-c$).
+If the test rejects at significance level $\alpha^{\prime}$ (default 0.1%), the method proceeds with a more flexible specification that only assumes Assumption 1 but not Assumption 2.
+Otherwise, it uses the default specification under both assumptions.
+This adaptive approach helps explain the strong empirical performance of plrd across all simulation settings in Table 3, including those where Assumption 2 fails.
+When the flexible specification is selected, estimation of the Lipschitz constant $B$ changes to accommodate potentially different curvature on either side of the threshold.
+Rather than fitting a single cubic polynomial as in (10), we fit separate cubic regressions for treated and control units:
+Proposition 6 (Consistent estimation of the Lipschitz constant $B$ under different curvatures).
+Suppose that Assumption 1 holds and that $X_{i}$ has a density bounded away from 0 in a neighborhood of $c$, and $B:=\max\{|\mu_{0}^{\prime\prime\prime}(c)|,|\mu_{1}^{\prime\prime\prime}(c)|,\varepsilon\}$.
+Estimate $\widehat{B}$ via local cubic regression: For $w\in\{0,1\}$ run
+$\widehat{\beta}_{w}:=\mathop{\rm argmin}_{\beta_{w}\in\mathbb{R}^{4}}\sum_{\begin{subarray}{c}W_{i}=w\\ |X_{i}-c|\leq\ell\end{subarray}}\left(Y_{i}-\beta_{w,0}-\beta_{w,1}(X_{i}-c)-\frac{\beta_{w,2}}{2}(X_{i}-c)^{2}-\frac{\beta_{w,3}}{6}(X_{i}-c)^{3}\right)^{2},$ (19)
+with bandwidth $\ell$ scaling as in (6), and set $\widehat{B}=\max\{|\beta_{0,3}|,|\beta_{1,3}|,\varepsilon\}$. Then, $\widehat{B}\rightarrow_{p}B$ as $n\to\infty$.
+With this estimated $\widehat{B}$ in hand, the subsequent minimax optimization in Algorithm 1 adjusts the constraint set to allow for different curvature functions $\rho_{0}$ and $\rho_{1}$ on either side of the cutoff (see (4) versus (3)).
+The optimization problem remains computationally tractable via quadratic programming, and the remainder of the PLRD procedure—including cross-fitting, variance estimation, and confidence interval construction—proceeds similarly to the default case.
+We note that a formal proof of asymptotic validity under this flexible specification would follow similar arguments to those in Theorem 1, with nearly identical technical steps that are omitted here.
+Appendix B Details on Baseline Methods
 In this section, we give further details on the baseline methods rdrobust and RDHonest considered in our experiments, and provide further discussion of how our modeling assumptions differ from those used to-date in software for bias-aware inference in RDDs.
-A.1 Bias-corrected confidence intervals
+B.1 Bias-Corrected Confidence Intervals
 Define the local polynomial estimator of order $p$ as $\widehat{\tau}_{p}(h_{n}):=\widehat{\mu}_{1,\,p}(h_{n})-\widehat{\mu}_{0,\,p}(h_{n}),$ where $\widehat{\mu}_{1,\,p}(h_{n})$ and $\widehat{\mu}_{0,\,p}(h_{n})$ are the intercepts from local polynomial regressions of order $p$ fitted separately on each side of the threshold.
@@
 where $\widehat{V}^{\text{bc}}_{n,p,q}$ is an estimator of the asymptotic variance that accounts for both the variability in estimation and the bias-correction.
-In Tables 1, 2, 3 and 4, the columns marked as b.c.
-and rob correspond to the confidence intervals in (18) and (19), respectively.
-A.2 Bias-aware confidence intervals
+In Tables 1, 2 and 3, the columns marked as rdrobut correspond to the confidence intervals computed using (21).
+B.2 Bias-Aware Confidence Intervals
 The bias-aware inference approach (e.g., Armstrong and Kolesár, 2018, 2020; Imbens and Wager, 2019) provides a different approach for obtaining confidence intervals for $\tau_{c}$.
 Note that the LLR estimator takes the form $\widehat{\tau}_{\texttt{LLR}}=\sum_{i=1}^{n}\gamma_{i,\texttt{LLR}}Y_{i}$ for some weights $\gamma_{i,\texttt{LLR}}$ that are functions of the running variable.
-In fact, the same also holds more generally for the estimator $\widehat{\tau}_{p}(h_{n})$ defined in Section A.1.
+In fact, the same also holds more generally for the estimator $\widehat{\tau}_{p}(h_{n})$ defined in Section B.1.
 Next, one makes the observation that although $\widehat{\tau}_{\texttt{LLR}}$ is motivated by a regression problem, the rest of our inference procedure hinges upon on properties that hold for any linear estimators of the form $\widehat{\tau}(\gamma)=\sum_{i=1}^{n}\gamma_{i}Y_{i}$.
@@
 The second option, which simply takes all functions to belong to a 2nd-order Hölder class, was used by Kolesár and Rothe (2018) and Imbens and Wager (2019) and also adopted for the RDHonest package by Armstrong and Kolesár (2018).
-Fixing a value of $M$, we numerically solve the optimization problem (20) and obtain an upper bound on the worst-case bias of the resulting estimator as a by-product.
-We then follow Imbens and Manski (2004) to construct a bias-aware confidence interval by inflating the Gaussian confidence interval accounting for the worst-case bias of $\tau(\widehat{\gamma})$ under the function class $\mathcal{M}$.
-These bias-aware confidence intervals provide uniformly valid coverage in finite samples, provided the function class $\mathcal{M}$ and the distribution of $Y_{i}$ given $X_{i}$ is Gaussian (Armstrong and Kolesár, 2018).
-[^12] If the conditional distribution of $Y_{i}$ is non-Gaussian, one can recover exact finite-sample results using the bias-aware approach by using finite-sample concentration results instead of Gaussian approximation.
-In practice, however, smoothness constant $M$ is typically unknown; both Armstrong and Kolesár (2020) and Imbens and Wager (2019) suggest using some rule-of-thumb estimates.
-A.3 The choice of function class
+[…]
+B.3 The Choice of Function Class
 One of the key insights our paper provides is that the choice of the function class $\mathcal{M}$ in the minimax problem (9) plays a crucial role in determining the conservativeness (and thus the practical usefulness) of the resulting bias-aware method.
@@
 Here, on the other hand, use the class of all functions $\mu_{w}(\cdot)$ that satisfy Assumption 2 and for which $\mu_{0}^{\prime\prime}(\cdot)$ is $\widehat{B}-$Lipschitz, where we obtain $\widehat{B}$ via a global cubic regression and cross-fitting.
-While this might appear as a minor modification from a theoretical point of view, it materially affects the length of our intervals in practice, as we demonstrate in Tables 1, 2, 3 and 4.
-To illustrate the difference between the conservativeness of our function class versus the function classes used by existing bias-aware approaches, we plot in Fig.
-3 the regression functions corresponding to the worst-case bias obtained by our method and $\mathcal{M}_{\text{H2},\,M}$ as used in RDHonest.
+While this might appear as a minor modification from a theoretical point of view, it materially affects the width of our intervals in practice, as we demonstrate in Tables 1, 2 and 3.
+To illustrate the difference between the conservativeness of our function class versus the function classes used by existing bias-aware approaches, we plot in Figure 2 the regression functions corresponding to the worst-case bias obtained by our method and $\mathcal{M}_{\text{H2},\,M}$ as used in RDHonest.
 This plot demonstrates that the classes $\mathcal{M}_{\text{H2},\,M}$ can often be unreasonably broad, and include regression functions that are ex ante implausible or rarely occur in practice.
-This explains why the resulting bias-aware confidence intervals can be substantially more conservative than the bias-corrected ones (see, e.g.Table 3).
+This explains why the resulting bias-aware confidence intervals can be substantially more conservative than the bias-corrected ones (see, e.g.Tables 1, 2 and 3).
 The empirical success of our PLRD method that we demonstrate in Section 4 is, in fact, rooted in this choice of the function class over which one computes the worst-case MSE.
@@
 In each plot, black lines show $\mu_{w}(\cdot)$ functions with maximum (solid) or minimum (dotted) bias in the respective function class.
-This figure illustrates how the RDHonest function class can be too modest for many regression discontinuity designs in practice.
-A.4 Ablation analysis
+This figure illustrates how the RDHonest function class may accommodate in-hindsight-unreasonable functions for several practical settings.
+B.4 Ablation Analysis
 Here we present an empirical comparison of our PLRD method with other alternative methods.
@@
 Table 4 compares our PLRD method with four other alternative approaches:
+Dataset Alt. (a) Alt. (b) Alt. (c) Alt. (d) plrd
+Lee (2008) 0.041 0.051 0.045 0.043 0.045
+(96.6%) (95.8%) (98.2%) (99.7%) (98.1%)
+Cattaneo et al. (2015) 5.567 6.880 5.652 9.709 3.870
+(96.6%) (96.0%) (98.3%) (99.8%) (96.1%)
+Ludwig and Miller (2007) 6.626 8.104 7.049 3.905 4.272
+(96.8%) (96.3%) (97.8%) (87.4%) (96.8%)
+Meyersson (2014) 7.011 8.736 6.432 7.599 4.618
+(96.3%) (95.9%) (96.7%) (97.9%) (96.3%)
+Matsudaira (2008) (read.) 0.056 0.070 0.059 0.085 0.059
+(90.5%) (95.4%) (97.7%) (94.8%) (97.7%)
+Matsudaira (2008) (math) 0.051 0.061 0.052 0.075 0.052
+(95.0%) (95.1%) (97.6%) (92.4%) (97.6%)
+Jacob and Lefgren (2004) (read.) 0.076 0.099 0.082 0.070 0.076
+(96.4%) (95.1%) (95.4%) (97.8%) (97.5%)
+Jacob and Lefgren (2004) (math) 0.066 0.085 0.070 0.070 0.061
+(96.9%) (95.0%) (95.5%) (97.1%) (97.9%)
+Oreopoulos (2006) 0.210 0.393 0.244 0.183 0.124
+(96.2%) (96.6%) (98.3%) (98.3%) (96.2%)
+Pure noise simulation 1.611 2.053 1.678 1.049 1.011
+(96.2%) (98.4%) (97.1%) (98.3%) (97.2%)
+Simulation Setting 1 0.238 0.301 0.250 0.169 0.237
+(93.5%) (97.5%) (95.2%) (88.9%) (94.8%)
+Simulation Setting 2 0.330 0.313 0.389 0.468 0.389
+(87.9%) (86.6%) (97.3%) (83.9%) (97.3%)
+Simulation Setting 3 0.252 0.300 0.290 0.421 0.290
+(94.9%) (97.5%) (99.6%) (99.9%) (99.6%)
+Simulation Setting 4 0.235 0.300 0.276 0.368 0.178
+(96.2%) (98.2%) (99.3%) (100.0%) (96.1%)
+Table 4: Comparison of PLRD with other alternatives: first two columns correspond to inflating conventional confidence intervals: instead of $1.96$, we multiply the standard error by a factor of (a) $2.181$ or (b) $2.113$, as suggested by Armstrong and Kolesár (2020); the next two columns correspond to (c) PLRD minus Assumption 2, and (d) PLRD minus Assumption 1; the last column corresponds to our PLRD method.
+We write the empirical coverages in parentheses and the average width in the rows.
+Other details of this table are same as for Tables 1 and 3 in the main text.
   1. (a)
-    Bias-aware intervals for local linear regression following Armstrong and Kolesár (2020):
+    Bias-aware intervals for local linear regression with inflated critical values as proposed by Armstrong and Kolesár (2020): [^17] Our end-to-end implementation differs from that in Armstrong and Kolesár (2020), as we choose bandwidths for local linear regression using the bandwidth selector of Imbens and Kalyanaraman (2012) as implemented in rdrobust; whereas Armstrong and Kolesár (2020) discuss adaptive bandwidth choice using an alternative rule of thumb.
+Here, we chose to pick bandwidths as in rdrobust to enable a cleaner comparison to the conventional estimator.
     1. (i)
@@
     3. (iii)
-      Build the confidence interval $\widehat{\tau}_{\texttt{LLR}}\pm 2.181\cdot\widehat{\text{SE}}_{\texttt{LLR}}$ (Armstrong and Kolesár, 2020).
+      Build the confidence interval $\widehat{\tau}_{\texttt{LLR}}\pm 2.181\cdot\widehat{\text{SE}}_{\texttt{LLR}}$.
   2. (b)
-    Bias-aware intervals for local quadratic regression following Armstrong and Kolesár (2020):
+    Bias-aware intervals for local quadratic regression with inflated critical values as proposed by Armstrong and Kolesár (2020):
     1. (i)
-      Fit a local quadratic regression using the rdrobust package (i.e., with $p=2$ in Section A.1), with default bandwidth selection rule.
+      Fit a local quadratic regression using the rdrobust package (i.e., with $p=2$ in Section B.1), with default bandwidth selection rule.
     2. (ii)
-      Obtain the LQR estimator $\widehat{\tau}_{\texttt{LQR}}$ (defined as $\widehat{\tau}_{2}(h_{n})$ in Section A.1) with its conventional standard error estimate $\widehat{\text{SE}}_{\texttt{LQR}}$
+      Obtain the LQR estimator $\widehat{\tau}_{\texttt{LQR}}$ (defined as $\widehat{\tau}_{2}(h_{n})$ in Section B.1) with its conventional standard error estimate $\widehat{\text{SE}}_{\texttt{LQR}}$
     3. (iii)
-      Build the confidence interval $\widehat{\tau}_{\texttt{LQR}}\pm 2.113\cdot\widehat{\text{SE}}_{\texttt{LQR}}$ (Armstrong and Kolesár, 2020).
+      Build the confidence interval $\widehat{\tau}_{\texttt{LQR}}\pm 2.113\cdot\widehat{\text{SE}}_{\texttt{LQR}}$.
   3. (c)
@@
       As a by-product of the optimization problem, obtain a bound on the worst-case bias of our estimator within this class, and construct “bias aware” confidence intervals for $\tau_{c}$ using the Imbens and Manski (2004) construction.
-    Algorithmically, this is similar to our Algorithm 1; we have also implemented this as a part of our plrd package.
+    Algorithmically, this is similar to our Algorithm 1; we also implemented this as a part of our plrd package.
   4. (d)
@@
       Obtain an upper bound on the worst-case bias of our estimator within this class, and construct “bias aware” confidence intervals for $\tau_{c}$ using the Imbens and Manski (2004) construction.
-    Algorithmically, this is similar to the optrdd package by Imbens and Wager (2019). We also provide an implementation of this method as a part of our replication files.
-Dataset Alt. (a) Alt. (b) Alt. (c) Alt. (d) plrd
-Lee (2008) 0.040 0.047 0.041 0.038 0.035
-(96.5%) (96.3%) (97.8%) (99.8%) (98.3%)
-Cattaneo et al. (2015) 5.956 7.267 5.259 9.376 4.525
-(96.6%) (96.4%) (97.3%) (99.7%) (97.3%)
-Matsudaira (2008) (read.) 0.068 0.070 0.071 0.084 0.060
-(96.1%) (96.7%) (97.1%) (96.8%) (97.1%)
-Matsudaira (2008) (math) 0.050 0.056 0.059 0.073 0.051
-(94.4%) (95.7%) (97.3%) (95.2%) (97.0%)
-Jacob and Lefgren (2004) (read.) 0.075 0.092 0.078 0.075 0.065
-(96.8%) (96.5%) (96.0%) (99.3%) (94.6%)
-Jacob and Lefgren (2004) (math) 0.087 0.102 0.082 0.077 0.069
-(95.6%) (94.9%) (97.2%) (96.2%) (95.0%)
-Oreopoulos (2006) 0.218 0.401 0.225 0.185 0.194
-(95.2%) (96.9%) (98.6%) (94.5%) (96.1%)
-Ludwig and Miller (2007) 4.516 5.564 4.075 3.846 3.542
-(97.2%) (96.5%) (97.6%) (98.2%) (97.4%)
-Meyersson (2014) 7.079 8.796 5.905 7.837 5.095
-(95.7%) (94.7%) (96.7%) (98.5%) (97.3%)
-Pure noise simulation 1.098 1.375 1.612 1.179 0.847
-(96.6%) (98.6%) (97.3%) (98.2%) (96.3%)
-Simulation Setting 1 0.223 0.278 0.276 0.202 0.235
-(92.7%) (97.5%) (97.2%) (92.5%) (94.6%)
-Simulation Setting 2 0.297 0.289 0.375 0.493 0.309
-(86.4%) (86.1%) (97.7%) (98.2%) (51.5%)
-Simulation Setting 3 0.236 0.279 0.280 0.300 0.239
-(93.3%) (96.2%) (97.1%) (98.2%) (95.9%)
-Simulation Setting 4 0.223 0.276 0.279 0.255 0.236
-(93.6%) (96.5%) (96.7%) (98.1%) (97.5%)
-Table 5: Comparison of PLRD with other alternatives: first two columns correspond to inflating conventional confidence intervals: instead of $1.96$, we multiply the standard error by a factor of (a) $2.181$ or (b) $2.113$, as suggested by Armstrong and Kolesár (2020); the next two columns correspond to (c) PLRD minus Assumption 2, and (d) PLRD minus Assumption 1; the last column corresponds to our PLRD method.
-We write the empirical coverages in parentheses and the average width in the rows.
-Other details of this table are same as for Tables 1, 2 and 4 in the main text.
-The empirical results demonstrate that the rule-of-thumb methods (AK20 (a) and AK20 (b)) achieve valid coverage but at the cost of wider intervals, mirroring our earlier findings in Tables 1, 2 and 4.
+    Algorithmically, this is similar to the optrdd package by Imbens and Wager (2019).
+[^18] In fact, over the weaker smoothness class $\mathcal{M}_{\text{H2},\,M}$, Assumption 2 on partial linearity becomes much less useful, because the second derivative of the baseline effect can have (bounded) sudden jumps anyways.
+As such, we also expect numerical performance of this ablation baseline to be close to optrdd.
+We also provide an implementation of this method as a part of our replication files.
+The empirical results demonstrate that the rule-of-thumb methods (AK20 (a) and AK20 (b)) achieve valid coverage but at the cost of wider intervals, mirroring our earlier findings in Tables 1 and 3.
 The relaxed variants of PLRD offer insights into the role of each assumption: Dropping either Assumption 1 or Assumption 2 generally leads to more conservative intervals, suggesting that both assumptions contribute to the efficiency of PLRD.
-Appendix B Additional details on our real data experiments
-Here we provide detailed descriptions of the real datasets analyzed in the main paper. We reference each dataset by the paper that introduced it, same as in Tables 1, 3 and 5.
+Appendix C Additional Details on Real Data Experiments
+Here we provide detailed descriptions of the real datasets analyzed in the main paper. We reference each dataset by the paper that introduced it, same as in Tables 1, 2 and 4.
   1. 1.
@@
 The outcome of interest is the probability of winning subsequent elections.
-The dataset consists of $6,558$ observations from U.S.
+The dataset consists of 6,5586{,}558 observations from U.S.
 House elections between 1946 and 1998.
@@
 Senate elections instead of House elections.
-The running variable is the margin of victory, defined as the difference in vote percentages (ranging from $-100$ to $100$) between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$.
+The running variable is the margin of victory, defined as the difference in vote percentages (ranging from 100-100 to $100$) between the Democratic candidate and their strongest opponent, with the threshold set at $c=0$.
 The outcome of interest is the percentage of winning subsequent elections.
-The dataset consists of $1,390$ observations ($1,297$ excluding missing outcomes) from U.S.
+The dataset consists of 1,3901{,}390 observations (1,2971{,}297 excluding missing outcomes) from U.S.
 House elections between 1914 and 2010.
@@
 Imbens in 2008.
-This dataset has a slightly larger sample size $n=68,798$ than the one used in the original paper ($n=66,839$).
+This dataset has a slightly larger sample size n=68,798n=68{,}798 than the one used in the original paper (n=66,839n=66{,}839).
 Our findings mirror the results in Matsudaira (2008, Table 4) for fifth-grade students.
@@
 Jacob and Lefgren (2004) investigated the effectiveness of remedial education programs on student achievement, using a regression discontinuity design based on test score.
-We use the same dataset as Imbens and Zajonc (2009), which contains $70,831$ observations after dropping entries with missing information.
+We use the same dataset as Imbens and Zajonc (2009), which contains 70,83170{,}831 observations after dropping entries with missing information.
 The running variable is the minimum of the math and reading score for an exam in the spring and the outcome is the post-remedial program reading (resp.
@@
 The outcome here is years of completed education.
-Our dataset comprises $73,954$ observations, differing slightly from the original study due to variations in preprocessing; again, we use the same dataset also used in Kolesár and Rothe (2018) and focus on the reduced form analysis rather than a fuzzy RD approach (see Table 1, Panel C of the corrigendum to Oreopoulos (2006)).
+Our dataset comprises 73,95473{,}954 observations, differing slightly from the original study due to variations in preprocessing; again, we use the same dataset also used in Kolesár and Rothe (2018) and focus on the reduced form analysis rather than a fuzzy RD approach (see Table 1, Panel C of the corrigendum to Oreopoulos (2006)).
 Our findings are in line with Table 2 of Kolesár and Rothe (2018).
@@
 (2014), we consider as outcome variable the mortality rates per 100,000 for children between 5 and 9 years old, with Head Start-related causes, for 1973–1983.
-After cleaning missing values, this dataset includes $2,783$ observations.
+After cleaning missing values, this dataset includes 2,7832{,}783 observations.
 Table III and Figure IV (Panel A) in Ludwig and Miller (2007) illustrate that the Head Start program had a negative impact on the outcomes.
@@
 The study employs an RD design where the running variable is the Islamic margin of victory—the vote share gap between the leading Islamic and secular parties, thus with a natural cutoff of zero.
-The analysis is conducted at the municipality level and has $2,629$ observations.
-We replicate part of the study where the primary outcome is the school attainment for women who were (potentially) in high school during the period 1994-2000, measured with variables extracted from the 2000 census.
+The analysis is conducted at the municipality level and has 2,6292{,}629 observations.
+We replicate part of the study where the primary outcome is the school attainment for women who were (potentially) in high school during the period 1994–2000, measured with variables extracted from the 2000 census.
 This data have been used as a running example in Cattaneo et al.
@@
 (2024) and the original paper Meyersson (2014) — in municipalities where the Islamic party barely won, the educational attainment of women is roughly 3% higher than in municipalities where the party barely lost.
-Appendix C Implementation via Convex Optimization
+Appendix D Implementation via Convex Optimization
 In this section, we discuss how one can solve the optimization problem (9) using readily available software described in, e.g., Boyd and Vandenberghe (2004).
@@
 Since $\gamma_{i}$ are measurable functions of $\{X_{i}\}_{i=1}^{n}$, the bias of the estimator $\widehat{\tau}(\gamma)=\sum_{i=1}^{n}\gamma_{i}Y_{i}$ conditional on the running variable is given by
-$\mathbb{E}\left[\sum_{i=1}^{n}\gamma_{i}Y_{i}\mid\{X_{i}\}_{i=1}^{n}\right]-\tau_{c}$ $=\sum_{i=1}^{n}\gamma_{i}\mu_{W_{i}}(X_{i})-\tau_{c}$
+$\mathbb{E}\left[\sum_{i=1}^{n}\gamma_{i}Y_{i}\,\bigg|\,\{X_{i}\}_{i=1}^{n}\right]-\tau_{c}$ $=\sum_{i=1}^{n}\gamma_{i}\mu_{W_{i}}(X_{i})-\tau_{c}$
 $=\sum_{i=1}^{n}\gamma_{i}(W_{i}(\tau_{c}+\beta(X_{i}-c))+\mu_{0}(X_{i}))-\tau_{c}$
@@
 $\rho(x):=\frac{1}{B}\left(\mu_{0}(x+c)-\mu_{0}(c)-\mu_{0}^{\prime}(c)x-\frac{1}{2}\mu_{0}^{\prime\prime}(c)x^{2}\right).$
-[…]
+Note that for any $\mu_{w}(\cdot)\in\mathcal{M}_{B}$, the function $\rho(\cdot)$ belongs to the function class $\mathcal{R}:=\{f:f(0)=f^{\prime}(0)=f^{\prime\prime}(0)=0,f^{\prime\prime}\text{ is }1\text{-Lipschitz}\}$.
+We can therefore rewrite the optimization problem (25) as:
 $\widehat{\gamma}:=\mathop{\rm{minimize}}_{\gamma}\left[B^{2}\left(\sup_{\rho\in\mathcal{R}}\sum_{i=1}^{n}\gamma_{i}\rho(X_{i}-c)\right)^{2}+\sum_{i=1}^{n}\sigma_{i}^{2}\gamma_{i}^{2}\right].$
 Finally, we make the implicit constraints explicit and use an auxiliary variable $t$ and rewrite the above optimization problem as:
-$(\widehat{\gamma},\widehat{t}\,):=$ $\mathop{\rm{minimize}}_{\gamma,t}\left\{B^{2}t^{2}+\sum_{i=1}^{n}\gamma_{i}^{2}\sigma_{i}^{2}\right\}$
+[…]
 $\text{subject to }\ \sup_{\rho\in\mathcal{R}}\sum_{i=1}^{n}\gamma_{i}\rho(X_{i}-c)\leq t,\ \sum_{i=1}^{n}\gamma_{i}W_{i}=1,\ \sum_{i=1}^{n}\gamma_{i}\left(1-W_{i}\right)=-1,$ (26)
@@
 $G(\widetilde{\rho},\lambda;x,w):=\widetilde{\rho}(x)+\lambda_{2}w+\lambda_{3}(1-w)+\lambda_{4}wx+\lambda_{5}(1-w)x+\lambda_{6}x^{2}.$
-The function $\widetilde{\rho}(\cdot)$ in the above formulation corresponds to the function $\lambda_{1}\rho(\cdot)$ in the original problem.
-Thus, we can recover the optimal variables $(\widehat{\gamma},\widehat{t}\,)$ for the primal problem (24) from the dual optimal variable $\widehat{\theta}$ (i.e., solution of (25)), as follows.
+[…]
 $\widehat{\gamma}_{i}=-\frac{1}{2\sigma_{i}^{2}}G(\widehat{\theta};X_{i}-c,W_{i}),\quad\widehat{t}=\frac{1}{2B^{2}}\widehat{\lambda}_{1}.$
@@
 In our software implementation, we adopt this finite-dimensional approximation approach to compute the optimal weights using quadprog, a function for quadratic programming in R (R Core Team, 2023).
-Appendix D Main proofs
+Appendix E Main Proofs
 Proof of Lemma 4.
 For simplicity in notation, we assume without loss of generality that the regression discontinuity threshold is $c=0$.
-We split the proof into some auxiliary results, namely, Lemmas 10, 8, 7, 9 and 11 in Appendix E.
+We split the proof into some auxiliary results, namely, Lemmas 11, 9, 8, 10 and 12 in Appendix F.
 First, using tools from empirical process theory (see van der Vaart and Wellner (1996) for a book-level treatment), we prove a uniform concentration bound for the loss functions of the empirical and population dual problems.
@@
 which is $o_{p}(1)$ since $\ell=o(n^{-1/12})$. $\square$
-Appendix E Auxiliary results
-[…]
+Appendix F Auxiliary Results
+To set the stage, we recall the key concepts of entropy with and without bracketing for function classes, which will be central to the results presented in this section.
+Consider a class of real-valued functions $\mathcal{F}$ defined on a set $\mathcal{Z}$, equipped with a norm $\|\cdot\|$.
+The covering number $N(\varepsilon,\mathcal{F},\|\cdot\|)$ is the minimum number of balls of radius $\varepsilon$ needed to cover $\mathcal{F}$.
+The centers of the covering balls need not belong to $\mathcal{F}$, but they are required to have finite norms.
+The entropy (without bracketing) is defined as the logarithm of the covering number.
+An $\varepsilon$-bracket is a pair of functions $(l,u)$ such that $\|u-l\|<\varepsilon$ and $l\leq f\leq u$ for all $f\in\mathcal{F}$.
+The bracketing number $N_{[\,]}(\varepsilon,\mathcal{F},\|\cdot\|)$ is the minimum number of $\varepsilon$-brackets needed to cover $\mathcal{F}$.
+The upper and lower bounds $u$ and $l$ of the brackets need not belong to $\mathcal{F}$, but they are required to have finite norms.
+The entropy with bracketing is defined as the logarithm of the bracketing number.
 Following van der Vaart and Wellner (1996, Section 2.7), we consider the class of functions $f:\mathcal{Z}\to\mathbb{R}$ defined on a bounded set $\mathcal{Z}\subset\mathbb{R}$ such that the derivatives of $f$ up to order $k$ are uniformly bounded, and $f^{(k)}$ is Lipschitz continuous.
@@
 Consider the following class of functions:
-[…]
+$\mathcal{F}_{B,\,\ell,\,k}=\{\rho\mid\rho:[-\ell,\ell]\to\mathbb{R},\ \rho(0)=\rho^{\prime}(0)=\cdots=\rho^{(k)}(0)=0,\ \rho^{(k)}\text{ is $B$-Lipschitz}\},$
 where $k\geq 0$ is an integer and $0<\ell\leq 1$. It holds that
@@
 $\theta_{n}^{*}(\sigma^{2},B)=\mathop{\rm argmin}_{\theta\in\Theta(\ell)}L_{n}(\theta;\sigma^{2},B),\quad\text{where}\quad L_{n}(\theta;\sigma^{2},B):=\frac{1}{4\sigma^{2}}\mathbb{E}\left[G^{2}(\theta;X,W)\right]+\frac{\lambda_{1}^{2}}{4nB^{2}}+\frac{\lambda_{2}-\lambda_{3}}{n}.$
-In Algorithm 1, we sample splitting to estimate the unknown Lipschitz constant $B$ and the homoscedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$.
+In Algorithm 1, we sample splitting to estimate the unknown Lipschitz constant $B$ and the homoskedastic variance $\sigma^{2}={\rm Var}(Y_{i}\mid X_{i})$.
 However, for ease of exposition, we assume throughout this proof that an estimate $\widehat{B}$ of $B$ and an estimate $\widehat{\sigma}^{2}$ of $\sigma^{2}$ is available independently of the data using which we solve the empirical dual problem.
@@
 $\mathbb{E}\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})\rho^{2}(X)|$ $=\mathbb{E}\sup_{r\in\mathcal{A}}|(\mathbb{P}_{n}-\mathbb{P})r|$
-[…]
-[…]
+$\lesssim\frac{1}{\sqrt{n}}\int_{0}^{\|F\|_{P,2}}\sqrt{\log N_{[\,]}(\varepsilon^{\prime},\mathcal{A},L_{2}(P))}\,d\varepsilon^{\prime}$
+$\lesssim\frac{1}{\sqrt{n}}\int_{0}^{F}(\varepsilon^{\prime})^{-1/4}\left(\lambda_{n,1}^{*}\ell^{3}\right)^{1/2}\,d\varepsilon^{\prime}$
 $\lesssim\frac{1}{\sqrt{n}}F^{3/4}F^{1/4}=(\lambda_{n,1}^{*})^{2}\frac{\ell^{6}}{\sqrt{n}}.$
@@
 We can now apply (van der Vaart, 1998, Corollary 19.35) with envelop $(4\|\lambda_{n,-1}^{*}\|_{2})^{2}$ to conclude that
-[…]
+$\mathbb{E}\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})(\lambda_{-1}^{\top}v(X,W))^{2}|$ $\lesssim\frac{1}{\sqrt{n}}\int_{0}^{16\|\lambda_{n,-1}^{*}\|^{2}}\sqrt{\log N_{[\,]}(\varepsilon^{\prime},\Lambda_{n},L_{2}(P))}\,d\varepsilon^{\prime}$
 $\lesssim\|\lambda_{n,-1}^{*}\|_{2}^{2}\frac{\sqrt{\varepsilon}}{\sqrt{n}}.$
@@
 We now apply (van der Vaart, 1998, Corollary 19.35) once more to conclude that
-[…]
+$\mathbb{E}\sup_{(\rho,\lambda)\in\Theta_{n}}|(\mathbb{P}_{n}-\mathbb{P})\rho(X)\lambda_{-1}^{\top}v(X,W)|$ $\lesssim\frac{1}{\sqrt{n}}\int_{0}^{F_{ab}}\left((\varepsilon^{\prime})^{-1/6}F_{ab}^{1/6}+(\varepsilon^{\prime})^{-1/2}\varepsilon^{1/2}F_{ab}^{1/2}\right)\,d\varepsilon^{\prime}$
 $\lesssim\frac{1}{\sqrt{n}}\left((F_{ab})^{5/6}(F_{ab})^{1/6}+(F_{ab})^{1/2}\varepsilon^{1/2}(F_{ab})^{1/2}\right)$
@@
     $\frac{2\lambda_{n,1}^{*}}{4\sigma^{2}}\mathbb{E}\left[\left(\widetilde{\rho}(X)-\widetilde{\rho}_{n}^{*}(X)\right)\left(\lambda_{n,1}^{*}\widetilde{\rho}_{n}^{*}(X)+\langle\lambda_{n,-1}^{*},v(X,W)\rangle\right)\right]\geq 0.$ (38)
-Equipped with Eqs.
-34, 35 and 36, we are ready to simplify $L_{n}(\theta)-L_{n}(\theta_{n}^{*})=\widetilde{L}_{n}(\widetilde{\rho},\lambda)-\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})$, as follows.
+Equipped with Equations 36, 37 and 38, we are ready to simplify $L_{n}(\theta)-L_{n}(\theta_{n}^{*})=\widetilde{L}_{n}(\widetilde{\rho},\lambda)-\widetilde{L}_{n}(\widetilde{\rho}_{n}^{*},\lambda_{n}^{*})$, as follows.
 $L_{n}(\theta)-L_{n}(\theta_{n}^{*})-\frac{1}{4\sigma^{2}}\mathbb{E}\left(\rho(X)-\rho_{n}^{*}(X)+\langle\lambda_{-1}-\lambda_{n,-1}^{*},v(X,W)\rangle\right)^{2}$
@@
 Lemma 11.
-Consider the neighborhood $\Theta_{n}(\ell,\delta,\varepsilon)$ of the optimizer $\theta_{n}^{*}=\theta_{n}^{*}(\widehat{\sigma}^{2},B)$ of the population dual problem, as defined in Lemma 7.
-Denote by $\widehat{\theta}_{n}=\widehat{\theta}_{n}(\widehat{\sigma}^{2},\widehat{B})$ the optimizer of the empirical dual problem.
-Assume that $\widehat{B}\rightarrow_{p}B$, and that $\ell=o(n^{-1/12})$.
-For any $\delta,\eta\in(0,1)$, and $\varepsilon=O(n^{-1/4})$, it holds that
+[…]
 $\mathbb{P}(\widehat{\theta}_{n}\in\Theta_{n}(\ell,\delta,\varepsilon))\geq 1-\eta,$
@@
 $\mathbb{P}\left(\widehat{L}_{n}(\theta_{n}^{*};\widehat{\sigma}^{2},\widehat{B})\leq\inf_{\theta\in\partial\widetilde{\Theta}_{n}}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})\right)\geq 1-\eta,$
-for all sufficiently large $n$.
-Finally, note that since the empirical loss function $\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})$ is convex in $\theta$, the above event implies that $\widehat{\theta}_{n}=\mathop{\rm argmin}_{\theta}\widehat{L}_{n}(\theta;\widehat{\sigma}^{2},\widehat{B})$ must be inside $\widetilde{\Theta}_{n}$.
-Thus,
+[…]
 $\mathbb{P}\left(\widehat{\theta}_{n}\in\widetilde{\Theta}_{n}\right)\geq 1-\eta,$
@@
 Proof of Lemma 12.
-[…]
+Recall that $\widehat{\lambda}_{n,1}(\widehat{\sigma}^{2},\widehat{B})=2\widehat{B}^{2}\widehat{t}_{n}$, where $\widehat{B}^{2}\widehat{t}_{n}^{\,2}$ is bounded above by the worst-case MSE of any linear estimator of the form $\sum_{i=1}^{n}\gamma_{i}Y_{i}$.
+In particular, there exists a kernel regression estimator that has worst-case MSE (supremum taken over our function class $\mathcal{M}_{\widehat{B}}$) of the order of $\mathcal{O}_{p}(n^{-6/7})$.
+To make it precise, consider the kernel regression estimator
 $\left(\widehat{\tau}(h_{n}),\widehat{\alpha},\widehat{\beta}\right)=\mathop{\rm argmin}_{\tau,\alpha,\beta_{1},\beta_{0}\in\mathbb{R}}\ \sum_{i=1}^{n}K\left(\frac{X_{i}}{h_{n}}\right)\left(Y_{i}-\alpha-\tau W_{i}-\beta_{0}X_{i}^{-}-\beta_{1}X_{i}^{+}-\beta_{2}X_{i}^{2}\right)^{2}.$
@@
 Recall that, ignoring cross-fitting notation as mentioned in the beginning of this section, we define
-$(\widehat{\gamma}_{n},\widehat{t}_{n}):=\mathop{\rm argmin}_{\gamma,\,t}\left(\widehat{B}^{2}t^{2}+\widehat{\sigma}^{2}\sum_{i=1}^{n}\widehat{\gamma}_{i}^{2}\right)\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\sum_{i=1}^{n}\gamma_{i}\rho(X_{i})\leq t,$
 […]
+where $\widehat{B}$ and $\widehat{\sigma}^{2}$ are fit externally, independent of the data $\{X_{i},Y_{i}\}_{i=1}^{n}$.
+The worst-case bias is given by $\widehat{b}=\widehat{B}\widehat{t}_{n}$, and thus the goal here is to show that $\widehat{t}_{n}^{\,2}/\widehat{V}_{n}=\mathcal{O}_{p}(1)$.
+Recall the population-level problem with externally fit nuisance parameters:
 $(\gamma^{*}_{n},t_{n}^{*}):=\mathop{\rm argmin}_{\gamma,t}\left\{\widehat{B}^{2}t^{2}+\frac{1}{n}\widehat{\sigma}^{2}\mathbb{E}[\gamma^{2}(X)]\right\}\quad\text{subject to}\quad\sup_{\rho\in\mathcal{R}}\mathbb{E}[\gamma(X)\rho(X)]\leq t.$
-[…]
+It suffices to prove that $t_{n}^{*}/\sqrt{V_{n}^{*}}=\mathcal{O}_{p}(1)$, where $V_{n}^{*}=\widehat{\sigma}^{2}\mathbb{E}[\gamma_{n}^{*,2}(X)]/n$, since we have consistency (i.e., that $\widehat{t}_{n}/t^{*}_{n}\rightarrow_{p}1$ and $\widehat{V}_{n}/V_{n}^{*}\rightarrow_{p}1$).
+It follows from Lemma 12 that $\widehat{B}^{2}t_{n}^{\,*,\,2}+V_{n}^{*}\leq_{p}a_{+}n^{-6/7}$ for some $a_{+}\in(0,\infty)$, for all sufficiently large $n$.
+On the other hand, it follows from the minimax lower bound (see, e.g.Cheng et al.
+(1997)) that $\widehat{B}^{2}t_{n}^{\,*,\,2}+V_{n}^{*}\geq_{p}a_{-}n^{-6/7}$ for some $0<a_{-}\leq a_{+}$ for all sufficiently large $n$.
+Therefore, to claim that $t_{n}^{*}/\sqrt{V_{n}^{*}}=\mathcal{O}_{p}(1)$, it suffices to show that there exists some $c>0$ such that $V_{n}^{*}\geq_{p}cn^{-6/7}$ for all sufficiently large $n$.
 To prove this by contradiction, assume that for any $c>0$, we have $V_{n}^{*}\leq_{p}cn^{-6/7}$ infinitely often. We examine the MSE of $\gamma_{n}^{*}$ for the objective
@@
 $=\widehat{B}^{2}(t_{n}^{*})^{2}+kV_{n}^{*}\leq_{p}(a_{+}+kc)k^{-6/7}m^{-6/7}.$
-[…]
+Choose $k$ large enough so that $a_{+}k^{-6/7}\leq a_{-}/2$, and then choose $c>0$ small enough such that $ck^{1/7}<a_{-}/2$.
+This, along with the above display, implies that $R_{m}(\gamma_{m}^{*})\leq R_{m}(\gamma_{n}^{*})\leq_{p}a_{-}m^{-6/7}$, which contradicts the minimax lower bound (we choose $n$ is large enough so that $m=n/k$ satisfies the minimax lower bound $\widehat{B}^{2}t_{m}^{\,*,\,2}+V_{m}^{*}\geq_{p}a_{-}m^{-6/7}$).
+This contradiction completes the proof.
+$\square$
 Lemma 14.